### v1

# Numerical simulation of type II primordial black hole formation

Abstract
This study investigates the formation of primordial black holes (PBHs) resulting from extremely large amplitudes of initial fluctuations in a radiation-dominated universe. We find that, for a sufficiently large initial amplitude, the configuration of trapping horizons shows characteristic structure due to the existence of bifurcating trapping horizons. We call this structure of the trapping horizons “Type II PBH”, while the structure without a bifurcating trapping horizon “Type I PBH”, which is typically generated from a relatively small amplitude of the initial fluctuation. In Ref. [1], in the dust-dominated universe, the Type II PBH can be realized by the Type II initial fluctuation, which is characterized by a non-monotonic areal radius as a function of the radial coordinate (throat structure) in contrast with the standard case with a monotonic areal radius (Type I fluctuation). Our research reveals that a type II fluctuation does not necessarily result in a type II PBH in the case of the radiation fluid. We also find that for the initial amplitude well above the threshold value, the resulting PBH mass may either increase or decrease with the initial amplitude depending on its specific profile rather than its fluctuation type.
1 Introduction
Primordial black holes (PBHs) play a crucial role in exploring the early universe, particularly in investigating inhomogeneities whose scales are much smaller than those of the fluctuations probed by the cosmological microwave background. Since PBHs are remnants of non-linear inhomogeneity in the primordial universe following the inflationary era, they are unique probes of statistical properties of the small-scale primordial inhomogeneity. They are considered as one of the candidates for dark matter and can be detected as black hole binaries by gravitational wave interferometers. According to the standard scenario, PBHs were formed through the gravitational collapse of significantly high-density regions during the radiation-dominated era, which had been generated in the inflationary period [2, 3, 4].
The mechanism of PBH formation can be briefly understood by considering a model of the gravitational collapse in the flat Friedmann–Lemaître–Robertson–Walker (FLRW) universe (see, e.g., Refs. [5, 6]). A criterion for the collapse has been firstly proposed by Carr [7] as $\delta\gtrsim w=1/3$, where $\delta$ is the amplitude of the density perturbation, and $w$ is the coefficient of the assumed linear equation of state $p=w\rho$ with $p$ and $\rho$ being the pressure and the energy density, respectively. A more accurate threshold value has been explored by numerically analyzing general relativistic non-linear dynamics [8, 9, 10, 5, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21].
In cosmological perturbation theory, the growth of the perturbation is often described by the evolution of the curvature perturbation $\zeta$ which represents the growing adiabatic modes. To perform numerical simulations of PBH formation, one must establish the initial conditions for the spacetime geometry and the fluid in the system. This involves setting the initial profile of curvature fluctuations at the super-horizon scale before the horizon entry to be consistent with the assumption of the growing adiabatic modes. These analytical constructions are typically achieved using the gradient expansion techniques [9, 14, 21].
In a spherically symmetric system, we introduce a type II primordial fluctuation that is characterized by a non-monotonic behavior of the areal radius as a function of the radial coordinate $r$. The areal radius $R$ is often expressed as $R\propto re^{\zeta}$ with $\zeta$ and $r$ being the curvature perturbation and radial coordinate, respectively. Then, if $\partial_{r}\zeta<-1/r$, which can be realized by a sufficiently large amplitude of $\zeta$ with the scale fixed, the Type II fluctuation is generally recognized. These fluctuations exhibit a distinct neck-like structure, with a minimal areal radius at $\partial_{r}R=0$.
Kopp, Hofmann, and Weller (KHW) [1] investigated PBH formation associated with the type II fluctuations by employing the Lemaître–Tolman–Bondi (LTB) solution, which is the analytic solution for the dust fluid ($w=0$). They investigated the causal structure of the spacetime generated from a specific profile of the type II fluctuation. They explicitly showed that it describes a spacetime of BH formation, which they call ‘type II PBH.’ In addition, they revealed distinct properties of the type II PBH from ‘type I PBH,’ which arises from typical type I fluctuations with the amplitude over the threshold. We aim to investigate the dynamics and the spacetime structure generated from type II fluctuations in the radiation-dominated universe, namely, to clarify the impact of the existence of the radiation pressure.
When we consider the abundance of PBHs, the contribution from type II fluctuations is often neglected because they are highly unlikely to be generated statistically, and the abundance of PBHs originating from them is typically suppressed. However, as mentioned in Ref. [22], they can be significant when the statistics of primordial fluctuations are highly non-Gaussian. Moreover, in the context of a single-field inflationary model, it has been recently discovered in Ref. [23] that type II fluctuations surround the bubbles formed when the inflaton overshoots the barrier. As non-Gaussianities increase, the production of vacuum bubbles becomes dominant, leading to a mass function and abundance primarily driven by these fluctuations. In addition, a similar black hole formation mechanism due to the birth of a baby universe with a tunneling process during inflation has been reported in Refs. [24, 25]. The baby universe is connected to our universe through a throat inside the black hole, like the neck structure in the type II fluctuation.
In the Misner–Sharp formalism [26] (see also Refs. [1, 27]), the evolution equations suffer from the trouble associated with $/0$ 0 term at the neck radius satisfying $\partial_{r}R=0$ when we consider type II fluctuations. To overcome this problem, it is necessary to conveniently use and modify the equations, as done in Ref. [24] for another scenario of PBH formation from the collapse of vacuum bubbles. On the other hand, in this paper, we use the simulation code developed from a 3+1 dimensional simulation code [28] with BSSN [29, 30] formalism with CARTOON method [31], so that we can avoid the same problem.
This paper is organized as follows. Section 2 describes the setup and numerical code of spherically symmetric numerical simulation of PBH formation. In Section 3.1, we present the simulation results, highlighting the time evolution of fluctuations. Then, we explore the formation of trapping horizons, including the BH horizon in Section 3.2, and give an interpretation of the resultant trapping horizon configurations in Section 3.3. In Section 4, we examine the evolution of PBH mass. In Section 4.2, to check the profile dependence of the PBH mass, we explore PBH formation from another initial profile that models the overdense region, as outlined by KHW. Finally, we offer our conclusions in the last section.
Throughout the paper, we use the geometrical unit in which the speed of light and the gravitational Newton constant are unity, i.e., $c=G=1$.
2 Initial data setting
We start our simulation from the time when the length scale of the fluctuation is much larger than the Hubble length $H_{\text{b}}^{-1}$, that is, $k^{-1}\gg(aH_{\text{b}})^{-1}$ with $k$, $a=a(t)$ and $H_{\text{b}}:=(da/dt)/a$ being the characteristic comoving wave number, the scale factor and the Hubble parameter of the background flat FLRW universe, respectively. We obtain the initial condition by applying the cosmological long-wavelength approximation, in which the small parameter $\epsilon:=k/(aH_{\rm b})$ is introduced, and the leading order spatial metric is given by [32]
$\gamma_{ij}=a^{2}(t)e^{2\zeta(x^{k})}f_{ij},\,$ (2.1)
with $f_{ij}$ being the reference 3-metric of the background flat geometry, and $\partial_{i}\zeta/(aH_{\rm b})=\mathcal{O}(\epsilon)$. We note that the function $\zeta$ is a time-independent arbitrary function of the spatial coordinates $x^{k}$. Once we initially set $\zeta$ as a function of the spatial coordinates $x^{i}$ and fix the gauge conditions, we can successively obtain the growing mode solution up to the following leading terms in this expansion [9, 21]. These long-wavelength solutions are used as the initial data for the time evolution. In this paper, we initially fix the gauge by imposing the constant mean curvature slice (uniform Hubble slice) and the normal threading, for which the trace of the extrinsic curvature of the time slice is uniformly given by $-3H_{\rm b}$ and the shift vector is equal to zero.
In a spherically symmetric system, the line element at the lowest order is denoted as
$ds^{2}=-dt^{2}+a^{2}(t)e^{2\zeta(r)}\quantity(dr^{2}+r^{2}d\Omega^{2}),$ (2.2)
where $d\Omega^{2}$ represents the line element of a 2-sphere. The areal radius is given by $R=are^{\zeta}$. Throughout this paper, we use the scale-up coordinate $z$, which is related to $r$ as
$r=z-\frac{\eta}{1+\eta}\frac{L}{\pi}\sin\left(\frac{\pi}{L}z\right)$ (2.3)
with a parameter $\eta=20$ (see Ref. [28] for details). In this coordinate, the resolution in the central region is $1+2\eta$ times finer than that around the boundary $z=r=L$.
Up through section 4, we use the following Gaussian-shape profile in the numerical simulation,
$\zeta(r)=\mu e^{-\frac{1}{2}k^{2}r^{2}}W(r),$ (2.4)
where $\mu$ is the amplitude of the fluctuation and $W(r)$ is a window function introduced to eliminate the tail of the Gaussian function and impose $\zeta=0$ at the outer boundary of the numerical domain (see Ref. [28] for the explicit functional form). Figure 1 shows the functional form of this profile (left) and the areal radius $R$ (right). By taking a sufficiently large amplitude of the fluctuation $\mu$, the areal radius $R(r)$ of the perturbed region becomes non-monotonic along the radial coordinate $r$, i.e., a pair of the local minimum and maximum points appears with $\partial_{r}R=0$ for $\mu\gtrsim 1.4$ as in Fig. 1. We call this structure neck or throat structure, and KHW defined such fluctuations as of type II and distinguished them from type I fluctuations, which have a monotonic areal radius [1].
Figure 1: Initial profile of the curvature fluctuation $\zeta$ (left) and areal radius $R$ as functions of the scale-up coordinate radius $z/L$. By taking the amplitude parameter $\mu$ larger, a type II fluctuation appears for $\mu=1.4$ and above with the neck $\partial_{r}R=0$.
The feature of type II fluctuations can also be understood as two peaks [23] of the Shibata–Sasaki compaction function $\mathcal{C}_{\text{SS}}$ defined by
$\mathcal{C}_{\text{SS}}(r)=\frac{1}{2}\left[1-\quantity(1+r\frac{d\zeta}{dr})^{2}\right]=\frac{1}{2}\left[1-\left(a^{-1}e^{-\zeta}\partial_{r}R\right)^{2}\right].$ (2.5)
This function has been firstly introduced in Ref. [9] as an estimator for the criterion of PBH formation, and its usefulness has been recognized through subsequent related works (see Ref. [33] and references therein for details including its relations to legitimate compaction functions in different gauges and physical interpretations). The outer peak of $\mathcal{C}_{\text{SS}}$ corresponds to the neck in the type II fluctuation, the radius at which the areal radius takes the minimum value considered relevant for the PBH formation. We may also consider a two-peak configuration with type I fluctuations [34] while we only focus on two peak configurations associated with type II fluctuations. The form of the compaction function for the specific function of $\zeta$ (2.4) is shown in Fig. 2. By using position of the peak of $\mathcal{C}_{\text{SS}}(r)$, $r_{\text{m}}=\sqrt{2}/k$, we analytically obtain the threathold value $\mu=e/2=1.359\cdots$.
Figure 2: Compaction function $\mathcal{C}_{\text{SS}}$. The horizontal axes are the scale-up coordinate radius $z/L$ (left) and the areal radius $R$ (right). By taking the amplitude $\mu$ larger, the number of peaks becomes two for $\mu\simeq 1.4$ and above, corresponding to type II fluctuations.
To follow the time evolution of gravitational collapse from the initial profile of fluctuation, we use the numerical code COSMOS-S [35, 36, 37] based on a 3+1 dimensional code of fourth order Runge–Kutta method with BSSN formalism [29, 30]. This code is specialized for spherically symmetric systems using the CARTOON method [31]. We consider the time evolution of the initial fluctuation given by the functional form of $\zeta$ (2.4) with the equation of state parameter $w=1/3$, namely, radiation fluid. Here, the asymptotically spatially flat FLRW boundary condition is implemented as in Ref. [9]. However, we could not avoid the constraint violation at the outer boundary, and this propagates into the bulk region. Therefore, in the following analyses, we terminate the time evolution before the violation reaches the central region relevant for the analyses of PBH formation.
We also find the constraint violation near the center for cases where an apparent horizon forms. Nevertheless, since we find both independent future-directed null directions are inward there, the constraint violation does not propagate outward. In practice, to avoid a crash of numerical calculation, we excise several grids inside the apparent horizon if needed.
We take the coordinate size $L$ of the numerical region ($0\leq z\leq L$ or, equivalently, $0\leq r\leq L$) as the unit length scale and set the initial value of the scale factor to unity. We fix the ratio between the background initial Hubble length $H_{\rm b}^{-1}$ and the typical scale $k^{-1}$ of the initial fluctuation as $k^{-1}=10H_{\text{b}}^{-1}$ which ensure the long-wavelength approximation. For a longer duration of the time evolution, we need to take a larger region than the initial Hubble length to avoid the propagation of the constraint violation. In the following, we take $H_{\text{b}}^{-1}=0.01L$ or $H_{\text{b}}^{-1}=0.005L$ depending on the duration of the required time evolution. The horizon entry time $t_{H}$ is defined as $k^{-1}=(aH_{\text{b}})^{-1}|_{t=t_{H}}$ throughout this paper.
3 Spacetime structure and a new definition of type I/II PBHs
3.1 Time evolution of type I/II fluctuations
First, we show the time evolutions of the fluid energy density and the lapse function in Fig. 3 for $\mu=0.5$, $1.2$, and $1.8$. For the cases of type I fluctuations $\mu=0.5$ (left) and $\mu=1.2$ (middle), the former disperses, but the latter contracts and forms a peak at the center in the density plot in Fig. 3. The latter case (middle) indicates a typical gravitational collapse, for which the lapse is decreasing to a very small value as plotted in Fig. 3 as Ref. [37]. For a type II fluctuation $\mu=1.8$ case (right), the collapsing behavior happens similarly to the $\mu=1.2$ (middle) case in the plots of density and lapse.
(a) $\mu=0.5$
(b) $\mu=1.2$
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
(a) $\mu=0.5$
(b) $\mu=1.2$
(c) $\mu=1.8$
Figure 3: Snapshots of the fluid energy density contrast $\rho/\rho_{\text{bg}}$ and lapse function $\alpha$ of type I fluctuations (left and middle) and a type II fluctuation (right). The initial configurations ($t=0.01t_{H}$) are shown in black on a different scale.
To check the convergence of the simulation, we plot the spatial average value of the Hamiltonian constraint (Fig. 4). One can find roughly quadratic convergence in the different spatial resolutions, and it is consistent with our numerical code with a second-order scheme implemented in evaluating the fluid flux.
(a) $\mu=0.5$
(b) $\mu=1.2$
(a) $\mu=0.5$
(b) $\mu=1.2$
(c) $\mu=1.8$
Figure 4: The average value of the Hamiltonian constraint violation with different resolutions of the number of grid points for three typical cases for no collapse (left), collapse from a type I fluctuation (middle), and collapse from a type II fluctuation (right), respectively. The horizontal axis is a time coordinate normalized by the time of the horizon entry. We use the simulation result to $t=50t_{H}$ to evaluate spacetime structures around horizon formation before the violation propagation reaches the bulk region from the outer boundary. For evaluating a longer time evolution of the PBH mass, the average is taken outside the BH horizon (or the future outer trapping horizon, described in Sec. 3.2) after the horizon formation at around $t=40t_{H}$ for $\mu=1.2$ and $t=24t_{H}$ for $\mu=1.8$.
3.2 Identification of BH horizon
First, let us introduce two independent future-directed radial null vector fields $l$ and $l^{\prime}$. We also define the null expansion for $l$ and $l^{\prime}$ as $\theta$ and $\theta^{\prime}$. Trapping horizons are defined by the surfaces that satisfy $\theta\theta^{\prime}=0$ or equivalently the compactness $12M/R=$ 1 with $M$ being the Misner–Sharp mass defined by $M(t,r)\coloneqq R(1-\nabla_{\mu}R\,\nabla^{\mu}R)/2$ [26, 38, 39, 6].
In Fig. 5, the snapshots of the compactness $\quantity(2M/R)(z)$ are plotted. For $\mu=0.5$, the intersection between $\quantity(2M/R)(r)$ and the horizontal line of $12M/R=$ 1 is always only one. On the other hand, for the cases $\mu=1.2$ and $1.8$, the line intersects the $12M/R=$ 1 more than three times at a sufficiently late time. The worldlines of the intersections are shown in Fig. 6. In order to identify the future-trapped ($\theta<0$ and $\theta^{\prime}<0$) and past-trapped ($\theta>0$ and $\theta^{\prime}>0$) regions we also show directions of two independent null vector fields $l$ and $l^{\prime}$ together with the contour curves of the areal radius in Figs. 7 and 8. Comparing the directions of the null vector fields and contour curves of the areal radius, we can realize that the red and blue shaded regions in Fig. 6 are future-trapped and past-trapped regions, respectively. More specifically, in the future(past)-trapped region, both arrows direct to directions along which the areal radius decreases (increases), and the curve described by red points indicates the boundary satisfying $\theta\theta^{\prime}=0$. Then, we can identify the outermost boundary surface of the future-trapped region as an apparent horizon. The cosmological horizon would be the outermost boundary between the past-trapped and non-trapped regions. Still, the following section will present a more concrete classification of trapping horizons.
(a) $\mu=0.5$
(b) $\mu=1.2$
(a) $\mu=0.5$
(b) $\mu=1.2$
(c) $\mu=1.8$
Figure 5: Snapshots of the compactness $2M(t,r)/R(t,r)$. For the $\mu=1.2$ case, the trapping horizon at the center is formed at around $t=40t_{H}$ (middle). For the $\mu=1.8$ case, the horizon is formed at around $t=24t_{H}$ (right).
(a) $\mu=0.5$
(b) $\mu=1.2$
(a) $\mu=0.5$
(b) $\mu=1.2$
(c) $\mu=1.8$
Figure 6: Configurations of the trapping horizons $R(t,r)=2M(t,r)$. It corresponds to the points $(t,r)$ of the compactness $12M/R=$ 1 in Fig. 5.
Figure 7: Directions of two independent null vector fields $l$ and $l^{\prime}$ are depicted as orange and blue arrows for $\mu=1.2$ case (type I) in the middle panel of Fig. 6. $R$-contour with 1.0 increments starting from $R=1.0$ and increasing from left bottom to right top is plotted as black curves.
Figure 8: Directions of two independent null vector fields $l$ and $l^{\prime}$ are depicted as orange and blue arrows for $\mu=1.8$ case (type II) in the right panel of Fig. 6. $R$-contour with 1.0 increments starting from $R=1.0$ and increasing from left bottom to right top is plotted as black curves. In addition to the original plot of $R$-levels with 1.0 increments, it is also plotted with 0.1 increments from $R=$ 8.1 to 8.9.
3.3 Spacetime structures of type I/II PBH
Let us introduce the classification of trapping horizons following Ref. [39, 40]. For this purpose, to avoid confusion, we introduce another notation for the two radial null vector fields and associated expansions as $l_{\pm}$ and $\theta_{\pm}$ for a specific trapping horizon. Then we always set $\theta_{+}=0$ when we consider a trapping horizon satisfying $\theta_{+}\theta_{-}=0$. The classification of trapping horizons can be summarized as Table 1.
$\theta_{+}=0$ + 0 -
$\theta_{-}$ Past Bifurcating Future
$\mathcal{L}_{-}\theta_{+}$ Inner Degenerate Outer
Table 1: Classification of trapping horizons satisfying $\theta_{+}\theta_{-}=0$ [39]. A black hole, white hole, and cosmological horizons with $w=1/3$ $(w=0)$ are identified with a future outer, past outer, and past degenerate (past inner) trapping horizons, respectively.
(a) Type I PBH
(a) Type I PBH
(b) Type II PBH
Figure 9: Sketch of type I and II PBHs in radiation domination. The left figure defines a type I PBH as a causal structure of the typical case of PBH. The right one includes bifurcating trapping horizons and defines a type II PBH.
Unlike the dust case discussed in Appendix A, we can only obtain numerical solutions such as those shown in Figs. 7 and 8 for the radiation case. From these figures and carefully comparing the similarities and differences from the exact conformal diagrams of the dust solution discussed in Appendix A, we can deduce possible conformal diagrams and proper identifications of the trapping horizons for the radiation case as shown in the left and right panels in Fig. 9 as possible conformal diagrams of type I and type II PBHs, respectively. The critical difference between the dust and the radiation cases is that trapping horizons in the FLRW regions are timelike and null for the former and the latter, respectively. See Figs. 16, 17 and 19 and Ref. [40]. Then, we confirm that the outermost boundary surface of the future-trapped region is the future outer trapping horizon, which describes an apparent horizon. In addition, one can find that the outermost boundary surface between the past-trapped and non-trapped regions describes the cosmological horizon, which asymptotically approaches the past degenerate trapping horizon in the background FLRW universe.
The crucial difference between the right and left panels in Fig. 9 is the existence of the bifurcating trapping horizons $(\theta_{-};\mathcal{L}_{-}\theta_{+})|_{\theta_{+}=0}=(0;*)$, which is described by the intersection points of the two trapping horizon trajectories in Fig. 8. In this paper, we propose to use this difference to classify PBH formation processes into types I and II; that is, we identify the PBH as of type I if there is no bifurcating trapping horizon and of type II if there appears at least one bifurcating trapping horizon during the time evolution associated with the intersection point. This crucial difference has already been pointed out by KHW in Ref. [1] for the case of the dust fluid. Unlike the current proposal, the authors in Ref. [1] did not clearly distinguish the definitions of type I/II PBHs and type I/II fluctuations. However, in our simulation with $w=1/3$, we find that the horizon configuration is of type I for the $\mu=1.6$ case, but the initial fluctuation is of type II nevertheless. Therefore, there is a gap between the definitions of type I/II PBHs and type I/II fluctuations. This fact is our motivation for the new definition of the PBH.
It is worth noting that the schematic diagram given by Fig. 9 implies that all light rays emitted from the center reach not the future null infinity but the future singularity in the type II PBH. Although we cannot make any general statement for the moment, the null trajectories indicated by the orange arrows in Fig. 8 seem consistent with this observation.
Precisely, one could also consider the critical case between type I and type II PBHs when the two bifurcating trapping horizons degenerate, then we additionally classify this case as a type IId PBH, which has just one bifurcating trapping horizon.
4 PBH mass
4.1 PBH mass with the Gaussian initial fluctuation
Let us consider how the mass of the PBH depends on the initial amplitude $\mu$ of the fluctuation in the type II regime by using the three-zone model with dust as seen in Fig. 18. The PBH mass is then given by the mass parameter for the Schwarzschild spacetime in the middle patch. If we fix the scale of the spatial curvature of the closed FLRW region, the horizon area of the Schwarzschild region is fixed by the horizon area on the matching surface between the closed FLRW and the Schwarzschild regions. Increasing the initial amplitude, we expect that the closed FLRW region will be extended and the boundary will be shifted more to the right. In the closed FLRW side, the shift of the boundary to the right results in the decreasing area of the horizons. Therefore, the mass parameter for the Schwarzschild region is also expected to decrease. From the above consideration of the model with dust, we may naively guess that the mass of the resultant PBH would decrease with the initial amplitude value for the type II regime. See also Ref. [41]. However, as will be shown soon, this guess is not necessarily true for the radiation case, and the behavior also strongly depends on the profile of the initial fluctuation.
Here, we evaluate the mass of the PBH in two different ways at each characteristic time. The one is to calculate the Misner-Sharp mass inside the apparent horizon at its first appearance with the future-trapped region (see Fig. 6). We refer to the mass at this time as the initial black hole mass. The apparent horizon appears as a future outer trapping horizon in a type I PBH, while it appears as a bifurcating outer trapping horizon in type II. We also calculate the final PBH mass by estimating the effect of the accretion. It is known that the accretion process is well described by the mass accretion model, often called Novikov-Zeldovich or Bondi accretion [2, 42]. In this accretion model, the time evolution of the mass is given by
[…]
where $\rho_{b}=\rho_{b,0}(t_{0}/t)^{2}$ is the background energy density, and $F$ is a constant that corresponds to the efficiency of accretion, which is commonly numerically found to be of order $\mathcal{O}(1)$. It is important to note that Eq. (4.1) is not valid at the time of formation of the trapping horizon since it neglects the cosmological expansion. But it will correctly describe the accretion into the PBHs for sufficiently large times once we have a quasi-stationary flow into the PBH. The integrated solution is,
$M(t)=\frac{1}{\frac{1}{M_{f}}+\frac{3F}{2t}},$ (4.2)
where $M_{f}$ is the final BH mass. We fix the values of $M_{f}$ and $F$ by fitting Eq. (4.2) to the numerical data in a sufficiently late finite time interval denoted in Fig. 10 together with the time evolution of the mass for each value of $\mu$. The values of $F$ obtained by the fitting are listed in Table 2. The value of $F$ decreases with increasing $\mu$, which means the accretion is getting smaller. This behavior is consistent with the results in Ref. [43].
Figure 10: The time evolution of the BH mass for each value of $\mu$. The black dashed curves are fitting curves with Eq. (4.2) including the two parameters $M_{f}$ and $F$. The fitting region is taken for the interval between $t=450t_{H}$ and $t=500t_{H}$.
$\mu$ $M_{f}[M_{\text{H}}]$ $F$
1.2 11.9 5.0
1.4 14.7 5.0
1.6 16.8 4.9
1.8 18.6 4.7
2.0 20.1 4.5
2.2 21.4 4.4
Table 2: The values of $F$ and $M_{f}$ for each amplitude of $\mu$ result from the fitting in Fig. 10.
Let us check the $\mu$ dependence of the PBH mass. In Fig. 11, we plot the initial and final PBH mass values for each value of $\mu$.
Figure 11: PBH mass in terms of the horizon mass at horizon entry $M_{H}$ to the amplitude of fluctuation $\mu$. In this case, in terms of fluctuation types, $\mu=1.2$ is classified as type I, while $\mu=1.4\>\text{to}\>2.2$ as type II, as seen in Fig. 2. In terms of formation types, on the other hand, $\mu=1.2$ to $1.6$ are classified as type I, while $\mu=1.8$ to $2.2$ as type II, as seen in Figs. 6 and 9.
The initial and final PBH masses get larger with increasing amplitude $\mu$. We will check the profile dependence in the next subsection.
4.2 PBH mass with the initial fluctuation motivated by the three-zone model
To realize a similar situation to the schematic conformal diagram, let us consider the closed FLRW region for the initial overdense region. Then, as is shown in the Appendix B, under the boundary condition $\zeta=0$ at the surface of the overdense region, we obtain the following form of $\zeta(r)$:
$\zeta_{\text{cFLRW}}(r)=2\ln\frac{\cos k\chi_{a}/2}{\sqrt{\cos^{4}k\chi_{a}/2+\frac{1}{4}k^{2}r^{2}}}\quad\text{for}\quad r\leq r_{a}:=\frac{1}{k}\sin(k\chi_{a}),$ (4.3)
where $1/k$ is the comoving radius of $S^{3}$ associated with the closed FLRW and $0<\chi_{a}\leq\pi/k$ describes the radial coordinate of the edge of the closed FLRW region. Although this expression is relevant for the region $r\leq r_{a}$, we simply extend this expression in the region $r>r_{a}$, and to realize the condition $\zeta(L)=0$, we use the following specific functional form of $\zeta(r)$:
$\zeta(r)$ $=2\tilde{\mu}\ln\quantity(\Psi_{3}(r)-\Psi_{3}(L))W(r),$ (4.4)
$\Psi_{3}(r)$ $\coloneqq e^{\zeta_{\text{cFLRW}}(r)/2}+e^{\zeta_{\text{ref}}(r)/2},$ (4.5)
where $\zeta_{\text{ref}}\coloneqq 0.001e^{-\frac{1}{2}k^{2}r^{2}}$ and $\tilde{\mu}\coloneqq k\chi_{a}/\pi$. The overall factor $\tilde{\mu}$ is multiplied for convenience to realize the background flat FLRW universe in the limit $\tilde{\mu}=0$. The value of $l$ is fixed as $k^{-1}=10H_{\rm b}^{-1}$ as before, and the value of $\chi_{a}$ is set as $\pi\tilde{\mu}/k$ for a given value of $\tilde{\mu}$.
Figure 12: Functional form of $\zeta_{\text{cFLRW}}$ for overdense region $\chi\leq\chi_{a}$. This profile is obtained by imposing a boundary condition that $\zeta=0$ at the edge of the overdense region $\chi=\chi_{a}$ [1].
The initial profiles of $\zeta_{\rm cFLRW}$, $\zeta$, $R$ and the compaction function $\mathcal{C}_{\rm SS}$ are shown in Figs. 12, 13 and 14. This profile also exhibits the type II feature with the neck structure for a large amplitude $(\tilde{\mu}\gtrsim 0.85)$.
Figure 13: Initial three-zone profile of $\zeta(r)$ (left) and areal radius $R(r)$ (right) with respect to the amplitude parameter $\tilde{\mu}$. By taking the amplitude parameter $\tilde{\mu}$ larger, a type II fluctuation appeared for $\tilde{\mu}$ larger than $0.80$ with the neck $\partial_{r}R=0$.
Figure 14: Compaction function $\mathcal{C}_{\text{SS}}(r)$ for thnon-linearprofile. The horizontal axis is the coordinate radius with scale-up $z/L$ (left)/ the areal radius $R$ (right).By taking the amplitude $\tilde{\mu}$ larger, the number of peaks becomes two for $\tilde{\mu}$ = 0.89, corresponding to type II fluctuation.
The behavior of the horizon configuration is similar to the case of the Gaussian profile, and we do not explicitly show it. In contrast to the Gaussian profile, the initial PBH mass formed by this profile decreases as the amplitude of fluctuation $\tilde{\mu}$ increases for $\tilde{\mu}>0.65$ as shown in Fig. 15, even though they are classified as type I fluctuations and formations for $\tilde{\mu}\leq 0.88$. This result, as well as the behavior of the Gaussian profile, indicates that whether the initial PBH mass increases or decreases with the amplitude parameter is not only up to its type but also strongly dependent on its profile, together with the ambiguity in quantifying the amplitude of nonlinear perturbation. We also find the gap between the parameter region of type I/II fluctuations and type I/II PBHs, as indicated in Fig. 15.
Figure 15: Initial PBH mass with respect to the amplitude of fluctuation for the three-zone profile.
The decreasing behavior of the initial PBH mass with the amplitude can be easily predicted by focusing on the profile of the compaction function as a function of the areal radius. The initial PBH mass would be roughly estimated by the areal radius of the peak of the compaction function at the horizon entry. Therefore, if the areal radius around the peak of the compaction function increases and decreases with the amplitude of the perturbation profile, the initial PBH mass is expected to increase and decrease, respectively. In the case of the Gaussian profile, one can find that the areal radius around the peak of the compaction function increases with the value of the amplitude $\mu$ as the right panel of Fig. 2, while it decreases with $\tilde{\mu}$ in the case of the specific profile (4.4) as the right one of Fig. 14.
5 Conclusion
We have numerically investigated the PBH formation from a type II initial fluctuation $\zeta(r)$, which is defined by the non-monotonicity of the areal radius $R(r)\propto re^{\zeta}$ with respect to the coordinate radius $r$ on the spatial geometry, i.e., the existence of $\partial_{r}R(r)=0$ (called neck or throat structure), in the radiation-dominated situation. As in the dust case reported in KHW, the PBH formation process can be classified into types I and II according to the configuration of the trapped regions $R(t,r)\leq 2M(t,r)$ (or trapping horizons $R(t,r)=2M(t,r)$). Specifically, if there is a bifurcating trapping horizon, the PBH is classified into type II, and type I otherwise. In the dust case, KHW reported that a type I fluctuation results in a type I PBH and a type II fluctuation does in a type II PBH. However, in the radiation fluid case, this is not true. We found a gap between the classifications in terms of the shape of the initial spatial geometry (fluctuation type) and the structure of trapping horizons (formation type). That is, we found that a type II fluctuation may result in a type I PBH. Therefore, we propose classifying PBHs in terms of the existence of the bifurcating trapping horizon independently of the type of the initial fluctuation.
In the leading order of the long-wavelength limit, the overdense region can be roughly approximated by the closed FLRW universe, and the amplitude of the density perturbation can be translated into the value of the radial coordinate at the edge of the closed FLRW region. Then, the type II fluctuation indicates that the edge is located in the other hemisphere of $S^{3}$ initial hypersurface beyond the great circle. Therefore, in the case of type II PBHs, one may expect that the PBH mass would decrease with increasing the amplitude of initial fluctuation since the initial time slice of $S^{3}$ in the overdensity region closes in the limit of large-density perturbation, which indicates the area of the edge is vanishing. However, this is not the case for the Gaussian-type initial fluctuation profile, while the other profile exhibits decreasing mass with increasing initial amplitude. Therefore, we conclude that the behavior of the initial PBH mass depends on not only the type of fluctuation but also the fluctuation profile. Thus, systematic investigation of the profile dependence would be needed, e.g., by using the shape parameter defined as the curvature peak of the compaction function $\mathcal{C}_{\text{SS}}$ [19, 43]. In addition to the profile dependence, the dependence on the equation of state $p=w\rho$ is also an interesting issue, which fills the gap between KHW and this work [20].
Acknowledgments
The authors are grateful to K.-I. Nakao, B.J. Carr, T. Igata, and J.M.M. Senovilla for useful comments and fruitful discussion. A.E. acknowledges support from the JSPS Postdoctoral Fellowships for Research in Japan (Graduate School of Sciences, Nagoya University). T.H. is supported in part by JSPS KAKENHI Grant Nos. 19K03876 and 20H05853. D.S. would like to take this opportunity to thank the “Nagoya University Interdisciplinary Frontier Fellowship” supported by Nagoya University and JST, the establishment of university fellowships towards the creation of science technology innovation, Grant Number JPMJFS2120. C.Y. is supported in part by JSPS KAKENHI Grant Nos. 20H05850 and 20H05853. C.Y. and T.H. acknowledge the hospitality at APCTP during the focus research program ”Black Hole and Gravitational Waves: from modified theories of gravity to data analysis”, where part of this work was discussed.
Appendix A Conformal diagrams of the three-zone model for the dust case
In the three-zone model [5], the spacetime of PBH is described by the patchwork of the three different regions : the innermost overdense region given by a spatially closed FLRW universe, the intermediate underdense region, and the outermost flat FLRW universe as the background. We also note that the intermediate underdense region must be introduced to compensate for the mass excess of the innermost region and connect the spacetime to the background spacetime without a mass gap. Without the intermediate underdense region, this model cannot be the exact solution, even for the dust case.
For the underdense region with a perfect fluid with $w=1/3$, it would not be tractable to obtain any analytical and exact solutions of such dynamical inhomogeneous spacetimes with nonlinear sound waves. Therefore, we stick to the three-zone model with dust. Furthermore, for simplicity, we focus on the model where the underdense region is described by the Schwarzschild solution, where the matching is possible without any singular hypersurface for the dust case. In spite of this great simplification, the model provides us with a very powerful tool for capturing the overall dynamics of PBH and the structure of trapping horizons. Note that one can also approach the same problem by using the Lemaître-Tolman-Bondi exact dust solution (c.f. [1]).
The conformal diagrams for the three spacetimes of which the three-zone model consists are shown in Fig. 16, where future $(\theta_{-};\mathcal{L}_{-}\theta_{+})|_{\theta_{+}=0}=(-;*)$, past (+;*)(+;*)( + ; * ), and bifurcating $(0;*)$ trapping horizons are denoted by red lines, blue lines, and the intersection point between them in the Schwarzschild region, respectively. For the dust case $w=0$, the BH horizon and cosmological horizon become future outer (;)(-;-)( - ; - ) and past inner (+;+)(+;+)( + ; + ), denoted by red null and blue timelike lines, respectively.
First, let us construct the spacetime diagram of the case of the type I fluctuation. The extracted regions for the patchwork in the type I case are described by the yellow shaded regions in Fig. 16, and the resultant spacetime diagram is shown in the left panel of Fig. 18. In these conformal diagrams, we define the right and left directed future null vectors as $l_{\rm out}$ and $l_{\rm in}$, respectively. For the closed FLRW region, on the boundary of the extracted region, we can find $\theta_{\rm out}>0$ and $\theta_{\rm in}<0$ at the maximum expansion. Therefore, in the outside Schwarzschild region, we also require the same sign of $\theta_{\rm out}$ and $\theta_{\rm in}$. Similarly, one can understand the matching of the Schwarzsdhild region and the background spatially flat FLRW spacetime.
For the type II fluctuation, the extracted regions are yellow-shaded in Fig. 17. Since the type II fluctuation should have a maximal value of the areal radius, the boundary of the closed FLRW region must be outside the central point at which the areal radius takes the maximum value corresponding to the area of the big sphere at the maximum expansion. Then we find $\theta_{\rm out}<0$ and $\theta_{\rm in}>0$ on the boundary, which can be matched with the Schwarzschild region only in the left part of the Schwarzschild conformal diagram. Thus, the resultant conformal diagram for the type II is given by the right panel of Fig. 18.
(a) closed FLRW for dust
(b) Schwarzschild
(a) closed FLRW for dust
(b) Schwarzschild
(c) flat FLRW for dust
Figure 16: Conformal diagrams to use for a patchwork for the spacetime of the type I PBH schematically in Fig. 18 for dust case $w=0$. Dashed gray curves represent $R$-levels. Red(blue)-shaded regions represent future(past) trapped regions. Future outer $(\theta_{-};\mathcal{L}_{-}\theta_{+})|_{\theta_{+}=0}=(-;-)$, past outer (+;)(+;-)( + ; - ), future inner (;+)(-;+)( - ; + ), and past inner (+;+)(+;+)( + ; + ) trapping horizons are denoted by red null, blue null, red timelike, and blue timelike lines, respectively.
(a) closed FLRW for dust
(b) Schwarzschild
(a) closed FLRW for dust
(b) Schwarzschild
(c) flat FLRW for dust
Figure 17: Same as Fig. 16 but for the type II PBH. We use the yellow-shaded regions for a patchwork.
(a) Type I PBH
(a) Type I PBH
(b) Type II PBH
Figure 18: Conformal diagrams of PBH for dust case $w=0$ as a patchwork of the three diagrams and trapping horizons. The left figure defines a type I PBH as a causal structure of typical case PBH. The right one is constructed by changing the cutting diagram range and defining a type II PBH.
(a) closed FLRW for w=1/3
(a) closed FLRW for w=1/3
(b) flat FLRW for w=1/3
Figure 19: Conformal diagrams of the closed and flat FLRW solutions for the radiation fluid case $w=1/3$. In this case, all trapping horizons are null.
Appendix B Derivation of $\zeta_{\text{cFLRW}}$ in radial coordinate $r$
The leading order of the long-wavelength solution in terms of spatial 3-metric is expressed as
$dl^{2}=a^{2}e^{2\zeta(r)}\quantity(dr^{2}+r^{2}d\Omega^{2}),$ (B.1)
where $r$ is a radial coordinate in the conformally flat geometry. We consider this perturbed metric of an overdense region as closed FLRW geometry with a constant positive curvature $k^{2}$,
$dl^{2}$ $=a^{2}\quantity(d\chi^{2}+\frac{1}{k^{2}}\sin^{2}k\chi\,d\Omega^{2}),$ (B.2)
where $\chi$ means the comoving radial coordinate. By comparing the correspondence of the metric, we obtain
$d\chi$ =±$=\pm dre^{\zeta},$ (B.3)
$\frac{1}{k}\sin k\chi$ $=re^{\zeta},$ (B.4)
then we get the relation,
$R$ $=\frac{a}{k}\sin k\chi=are^{\zeta},$ (B.5)
$r$ $=C\exp\int\frac{d(k\chi)}{\sin k\chi}=C\tan\frac{k\chi}{2},$ (B.6)
where C is an integral constant. By using these relations, $\zeta$ is obtained as
$e^{\zeta}$ $=\frac{\sin k\chi}{kr},$ (B.7)
$=\frac{2}{Ck}\cos^{2}\frac{k\chi}{2},$ (B.8)
$=\frac{2}{k}\frac{1/C}{1+\quantity(r/C)^{2}}.$ (B.9)
If we write $C$ as $2C^{\prime}/k$ for simplicity,
$e^{\zeta}=\frac{1/C^{\prime}}{1+\frac{k^{2}}{4}\quantity(r/C^{\prime})^{2}}\quad\therefore\quad\zeta=\ln\frac{1/C^{\prime}}{1+\frac{k^{2}}{4}\quantity(r/C^{\prime})^{2}}.$ (B.10)
Here we impose the boundary condition that $\zeta=0$ for $\chi=\chi_{a}$, we get
$C^{\prime}=\cos^{2}\frac{k\chi_{a}}{2}.$ (B.11)
Then
$\zeta=2\ln\frac{\cos k\chi_{a}/2}{\sqrt{\cos^{4}k\chi_{a}/2+\frac{1}{4}k^{2}r^{2}}}.$ (B.12)

### v2

@@
 They explicitly showed that it describes a spacetime of BH formation, which they call ‘type II PBH.’ In addition, they revealed distinct properties of the type II PBH from ‘type I PBH,’ which arises from typical type I fluctuations with the amplitude over the threshold.
+Carr and Harada (2015) [22] analytically discussed type II PBHs with more general equations of state using simplified models.
+Type I/II fluctuations are also important because the threshold of these two cases gives a possible maximum value of the averaged density perturbation at horizon entry in the simplified models [1, 22].
 We aim to investigate the dynamics and the spacetime structure generated from type II fluctuations in the radiation-dominated universe, namely, to clarify the impact of the existence of the radiation pressure.
@@
 In this paper, we propose to use this difference to classify PBH formation processes into types I and II; that is, we identify the PBH as of type I if there is no bifurcating trapping horizon and of type II if there appears at least one bifurcating trapping horizon during the time evolution associated with the intersection point.
-This crucial difference has already been pointed out by KHW in Ref.
+This difference has already been pointed out by KHW in Ref.
 [1] for the case of the dust fluid.
@@
 4.1 PBH mass with the Gaussian initial fluctuation
-Let us consider how the mass of the PBH depends on the initial amplitude $\mu$ of the fluctuation in the type II regime by using the three-zone model with dust as seen in Fig.
-18.
-The PBH mass is then given by the mass parameter for the Schwarzschild spacetime in the middle patch.
-If we fix the scale of the spatial curvature of the closed FLRW region, the horizon area of the Schwarzschild region is fixed by the horizon area on the matching surface between the closed FLRW and the Schwarzschild regions.
-Increasing the initial amplitude, we expect that the closed FLRW region will be extended and the boundary will be shifted more to the right.
-In the closed FLRW side, the shift of the boundary to the right results in the decreasing area of the horizons.
-Therefore, the mass parameter for the Schwarzschild region is also expected to decrease.
-From the above consideration of the model with dust, we may naively guess that the mass of the resultant PBH would decrease with the initial amplitude value for the type II regime.
-See also Ref.
-[41].
-However, as will be shown soon, this guess is not necessarily true for the radiation case, and the behavior also strongly depends on the profile of the initial fluctuation.
+[…]
 Here, we evaluate the mass of the PBH in two different ways at each characteristic time.
@@
 It is important to note that Eq.
-(4.1) is not valid at the time of formation of the trapping horizon since it neglects the cosmological expansion.
+(4.1) is not valid at the time of formation of the trapping horizon since it neglects the cosmological expansion as Carr and Hawking (1974) [4] already pointed out.
 But it will correctly describe the accretion into the PBHs for sufficiently large times once we have a quasi-stationary flow into the PBH.
@@
 $\zeta_{\text{cFLRW}}(r)=2\ln\frac{\cos k\chi_{a}/2}{\sqrt{\cos^{4}k\chi_{a}/2+\frac{1}{4}k^{2}r^{2}}}\quad\text{for}\quad r\leq r_{a}:=\frac{1}{k}\sin(k\chi_{a}),$ (4.3)
-where $1/k$ is the comoving radius of $S^{3}$ associated with the closed FLRW and $0<\chi_{a}\leq\pi/k$ describes the radial coordinate of the edge of the closed FLRW region.
+where $k^{-1}$ is the comoving radius of $S^{3}$ associated with the closed FLRW and $0<\chi_{a}\leq\pi/k$ describes the radial coordinate of the edge of the closed FLRW region.
 Although this expression is relevant for the region $r\leq r_{a}$, we simply extend this expression in the region $r>r_{a}$, and to realize the condition $\zeta(L)=0$, we use the following specific functional form of $\zeta(r)$:
@@
 15, even though they are classified as type I fluctuations and formations for $\tilde{\mu}\leq 0.88$.
-This result, as well as the behavior of the Gaussian profile, indicates that whether the initial PBH mass increases or decreases with the amplitude parameter is not only up to its type but also strongly dependent on its profile, together with the ambiguity in quantifying the amplitude of nonlinear perturbation.
+This result, as well as the behavior of the Gaussian profile, indicates that whether the initial PBH mass increases or decreases with the amplitude parameter is not only up to its type but also strongly dependent on its profile, together with the ambiguity in quantifying the amplitude of non-linear perturbation.
 We also find the gap between the parameter region of type I/II fluctuations and type I/II PBHs, as indicated in Fig.
@@
 Without the intermediate underdense region, this model cannot be the exact solution, even for the dust case.
-For the underdense region with a perfect fluid with $w=1/3$, it would not be tractable to obtain any analytical and exact solutions of such dynamical inhomogeneous spacetimes with nonlinear sound waves.
+For the underdense region with a perfect fluid with $w=1/3$, it would not be tractable to obtain any analytical and exact solutions of such dynamical inhomogeneous spacetimes with non-linear sound waves.
 Therefore, we stick to the three-zone model with dust.

### v3

@@
 Since PBHs are remnants of non-linear inhomogeneity in the primordial universe following the inflationary era, they are unique probes of statistical properties of the small-scale primordial inhomogeneity.
-They are considered as one of the candidates for dark matter and can be detected as black hole binaries by gravitational wave interferometers.
-According to the standard scenario, PBHs were formed through the gravitational collapse of significantly high-density regions during the radiation-dominated era, which had been generated in the inflationary period [2, 3, 4].
+They are considered as one of the candidates for dark matter and can be detected as black hole (BH) binaries by gravitational wave interferometers.
+According to the standard scenario, PBHs were formed through the gravitational collapse of significantly high-density regions during the radiation-dominated era, which had been generated in the inflationary period [2, 3, 4, 5, 6].
 The mechanism of PBH formation can be briefly understood by considering a model of the gravitational collapse in the flat Friedmann–Lemaître–Robertson–Walker (FLRW) universe (see, e.g., Refs.
@@
 These analytical constructions are typically achieved using the gradient expansion techniques [11, 16, 21].
-In a spherically symmetric system, we introduce a type II primordial fluctuation that is characterized by a non-monotonic behavior of the areal radius as a function of the radial coordinate $r$.
-The areal radius $R$ is often expressed as $R\propto re^{\zeta}$ with $\zeta$ and $r$ being the curvature perturbation and radial coordinate, respectively.
-Then, if $\partial_{r}\zeta<-1/r$, which can be realized by a sufficiently large amplitude of $\zeta$ with the scale fixed, the Type II fluctuation is generally recognized.
-These fluctuations exhibit a distinct neck-like structure, with a minimal areal radius at $\partial_{r}R=0$.
+In a spherically symmetric system, we introduce a type II primordial fluctuation characterized by a non-monotonic behavior of the areal radius $R=\sqrt{A/4\pi}$, where $A$ is the area of the 2-sphere as a function of the radial coordinate $r$ with constant $t$.
+These fluctuations exhibit distinct neck-like structure, with a minimal areal radius at $\partial_{r}R=0$.
+In cosmological perturbation theory, the areal radius $R$ is often expressed via the curvature fluctuation $\zeta$, $R\propto re^{\zeta}$.
+The type II fluctuation is only realized by a sufficiently large amplitude of $\zeta$ with the scale fixed.
 Kopp, Hofmann, and Weller (KHW) [1] investigated PBH formation associated with the type II fluctuations by employing the Lemaître–Tolman–Bondi (LTB) solution, which is the analytic solution for the dust fluid ($w=0$).
@@
 As non-Gaussianities increase, the production of vacuum bubbles becomes dominant, leading to a mass function and abundance primarily driven by these fluctuations.
-In addition, a similar black hole formation mechanism due to the birth of a baby universe with a tunneling process during inflation has been reported in Refs.
+In addition, a similar BH formation mechanism due to the birth of a baby universe with a tunneling process during inflation has been reported in Refs.
 [27, 28].
-The baby universe is connected to our universe through a throat inside the black hole, like the neck structure in the type II fluctuation.
+The baby universe is connected to our universe through a throat inside the BH, like the neck structure in the type II fluctuation.
+If one would like to follow the time evolution of the type II fluctuations via a usual areal radial coordinate system such as in Ref.
+[21], the neck throat structure would be seriously difficult to deal with.
 In the Misner–Sharp formalism [29] (see also Refs.
@@
 Finally, we offer our conclusions in the last section.
+Appendix A provides the conformal diagrams of the exact three-zone solutions for PBH formation.
+Appendix B provides the derivation of the curvature perturbation for the three-zone model.
+Appendix C proves the appearance of two peaks in the compaction function and the averaged density perturbation for the type II fluctuation.
 Throughout the paper, we use the geometrical unit in which the speed of light and the gravitational Newton constant are unity, i.e., $c=G=1$.
@@
 We start our simulation from the time when the length scale of the fluctuation is much larger than the Hubble length $H_{\text{b}}^{-1}$, that is, $k^{-1}\gg(aH_{\text{b}})^{-1}$ with $k$, $a=a(t)$ and $H_{\text{b}}:=(da/dt)/a$ being the characteristic comoving wave number, the scale factor and the Hubble parameter of the background flat FLRW universe, respectively.
-We obtain the initial condition by applying the cosmological long-wavelength approximation, in which the small parameter $\epsilon:=k/(aH_{\rm b})$ is introduced, and the leading order spatial metric is given by [33]
+We obtain the initial condition by applying the cosmological long-wavelength approximation, in which the small parameter $\epsilon:=k/(aH_{\rm b})$ is introduced111 One would use the other expression, $\epsilon:=1/(H_{\rm b}R_{\rm m})$ with $R_{\rm m}\propto a/k$ being the areal radius, where $\mathcal{C}_{\rm SS}(r)$ takes a maximum.
+, and the zero-th order spatial metric is given by [35]
 $\gamma_{ij}=a^{2}(t)e^{2\zeta(x^{k})}f_{ij},\,$ (2.1)
@@
 In this paper, we initially fix the gauge by imposing the constant mean curvature slice (uniform Hubble slice) and the normal threading, for which the trace of the extrinsic curvature of the time slice is uniformly given by $-3H_{\rm b}$ and the shift vector is equal to zero.
-In a spherically symmetric system, the line element at the lowest order is denoted as
+In a spherically symmetric system, the line element at the zero-th order in powers of $\epsilon$ is denoted as
 $ds^{2}=-dt^{2}+a^{2}(t)e^{2\zeta(r)}\quantity(dr^{2}+r^{2}d\Omega^{2}),$ (2.2)
-where $d\Omega^{2}$ represents the line element of a 2-sphere. The areal radius is given by $R=are^{\zeta}$. Throughout this paper, we use the scale-up coordinate $z$, which is related to $r$ as
+where $d\Omega^{2}$ represents the line element of a 2-sphere.
+We can see that the areal radius is given by $R=are^{\zeta}$ in this line element.
+Throughout this paper, to resolve the fine structure, such as the trapping horizon formation near the center $r=0$, we use the scale-up coordinate $z$, which is defined as
 $r=z-\frac{\eta}{1+\eta}\frac{L}{\pi}\sin\left(\frac{\pi}{L}z\right)$ (2.3)
-with a parameter $\eta=20$ (see Ref. [29] for details). In this coordinate, the resolution in the central region is $1+2\eta$ times finer than that around the boundary $z=r=L$.
-Up through section 4, we use the following Gaussian-shape profile in the numerical simulation,
+with a parameter $\eta=20$ [31]. In this coordinate, the resolution in the central region is $1+2\eta$ times finer than that around the boundary $z=r=L$;
+[…]
+Up through section 4, we use the following Gaussian-shaped profile in the numerical simulation [31],
 $\zeta(r)=\mu e^{-\frac{1}{2}k^{2}r^{2}}W(r),$ (2.5)
-where $\mu$ is the amplitude of the fluctuation and $W(r)$ is a window function introduced to eliminate the tail of the Gaussian function and impose $\zeta=0$ at the outer boundary of the numerical domain (see Ref.
-[29] for the explicit functional form).
+where $\mu$ is the amplitude of the fluctuation and $W(r)$ is a window function [31] introduced to eliminate the tail of the Gaussian function and impose $\zeta=0$ at the outer boundary of the numerical domain.222 The explicit functional form [31] is 11 1$\quad$1 $\text{for}\quad 0\leq r\leq r_{\text{w}},$ $1-\frac{\quantity((r_{\text{w}}-L)^{6}-(L-r)^{6})^{6}}{(r_{\text{w}}-L)^{36}}\quad$ $\text{for}\quad r_{\text{w}}\leq r\leq L,$ (2.6) 00 0$\quad$ $\text{for}\quad L\leq r,$ where $r_{\rm w}$ is set to $0.8L$.
 Figure 1 shows the functional form of this profile (left) and the areal radius $R$ (right).
@@
 By taking the amplitude parameter $\mu$ larger, a type II fluctuation appears for $\mu=1.4$ and above with the neck $\partial_{r}R=0$.
-The feature of type II fluctuations can also be understood as two peaks [24] of the Shibata–Sasaki compaction function $\mathcal{C}_{\text{SS}}$ defined by
-$\mathcal{C}_{\text{SS}}(r)=\frac{1}{2}\left[1-\quantity(1+r\frac{d\zeta}{dr})^{2}\right]=\frac{1}{2}\left[1-\left(a^{-1}e^{-\zeta}\partial_{r}R\right)^{2}\right].$ (2.5)
+The feature of type II fluctuations can also be understood as two peaks333 It is shown in Ref.
+[26] that the appearance of fluctuations of type II naturally appears via a specific single-field inflation model with a bump, but in this case, one of the peaks of $\mathcal{C}_{\rm SS}$ is hidden inside the bubble.
+of the Shibata–Sasaki compaction function $\mathcal{C}_{\text{SS}}$ defined by
+[…]
 This function has been firstly introduced in Ref.
@@
 The outer peak of $\mathcal{C}_{\text{SS}}$ corresponds to the neck in the type II fluctuation, the radius at which the areal radius takes the minimum value considered relevant for the PBH formation.
-We may also consider a two-peak configuration with type I fluctuations [35] while we only focus on two peak configurations associated with type II fluctuations.
+We will give a general proof for the existence of two peaks and the minimum between them for the compaction function in the type II fluctuation in Appendix C.
+We focus on the two peak configurations associated with type II fluctuations (see also Ref.
+[37] for two peak configurations with fluctuations of type I ).
 The form of the compaction function for the specific function of $\zeta$ (2.5) is shown in Fig.
 2.
-By using position of the peak of $\mathcal{C}_{\text{SS}}(r)$, $r_{\text{m}}=\sqrt{2}/k$, we analytically obtain the threathold value $\mu=e/2=1.359\cdots$.
+By using the position of the peak of $\mathcal{C}_{\text{SS}}(r)$, $r_{\text{m}}=\sqrt{2}/k$, we analytically obtain the threshold value $\mu_{\rm c,II}=e/2=1.359\cdots$ for fluctuations of type II.444 Here, we ignore the window function $W(r)$ to estimate the location $r_{\text{m}}=\sqrt{2}/k$ of the peak of compaction function $\mathcal{C}_{\rm SS}(r)$ since this acts on $\zeta(r)$ only near the outer boundary, $r=r_{\rm w}=0.8L$ For $\mu>\mu_{\rm c,II}$ the compaction function takes the maximum value of $0.5$ at least one location with the radial coordinate, and the number of peak positions becomes two, $r_{m1}$ and $r_{m2}$ with $r_{m1}<r_{m2}$.
+This implies that the type II fluctuations always reach the threshold for collapse and form PBHs irrespectively on the initial profile of $\zeta(r)$.
+The isotropic coordinate $r$ does not always give a geometrical size of the enclosed region and may often lead to misunderstanding.
+Therefore, we also explicitly show the compaction function $\mathcal{C}_{\rm SS}$ as a function of the areal radius $R$ in the right panel of Fig.
+2, which will be contrasted with the right panel of Fig.
+13.
 Figure 2: Compaction function $\mathcal{C}_{\text{SS}}$.
@@
 For a type II fluctuation $\mu=1.8$ case (right), the collapsing behavior happens similarly to the $\mu=1.2$ (middle) case in the plots of density and lapse.
+Refer to captionRefer to caption
 (a) $\mu=0.5$
+Refer to captionRefer to caption
 (b) $\mu=1.2$
 Refer to captionRefer to caption
-Refer to captionRefer to caption
-Refer to captionRefer to caption
-(a) $\mu=0.5$
-(b) $\mu=1.2$
 (c) $\mu=1.8$
@@
 The initial configurations ($t=0.01t_{H}$) are shown in black on a different scale.
-To check the convergence of the simulation, we plot the spatial average value of the Hamiltonian constraint (Fig.
+To check the convergence of the simulation, we plot the spatial average value of the Hamiltonian constraint555 First, we evaluate the appropriately normalized value of the constraint violation $\mathcal{H}_{p}$ at each grid point labeled by $p$.
+Then the average is taken as $\sum_{p=1}^{N}\mathcal{H}_{p}/N$, where $N$ is the total number of the grid points.
+[21](Fig.
 4).
 One can find roughly quadratic convergence in the different spatial resolutions, and it is consistent with our numerical code with a second-order scheme implemented in evaluating the fluid flux.
-(a) $\mu=0.5$
-(b) $\mu=1.2$
 (a) $\mu=0.5$
@@
 We use the simulation result to $t=50t_{H}$ to evaluate spacetime structures around horizon formation before the violation propagation reaches the bulk region from the outer boundary.
-For evaluating a longer time evolution of the PBH mass, the average is taken outside the BH horizon (or the future outer trapping horizon, described in Sec.
+For evaluating a longer time evolution of the PBH mass, the average is taken outside the BH apparent horizon (or the future-outer trapping horizon, described in Sec.
 3.2) after the horizon formation at around $t=40t_{H}$ for $\mu=1.2$ and $t=24t_{H}$ for $\mu=1.8$.
-3.2 Identification of BH horizon
-First, let us introduce two independent future-directed radial null vector fields $l$ and $l^{\prime}$.
+3.2 Identification of BH apparent horizon
+First, let us introduce two independent future-directed radial null vector fields, $l$ and $l^{\prime}$.
 We also define the null expansion for $l$ and $l^{\prime}$ as $\theta$ and $\theta^{\prime}$.
+A null expansion is expressed as $\theta\coloneqq 2\mathcal{L}_{l}R/R$, where $\mathcal{L}_{l}$ denotes the Lie derivative along $l$.
 Trapping horizons are defined by the surfaces that satisfy $\theta\theta^{\prime}=0$ or equivalently the compactness $12M/R=$ 1 with $M$ being the Misner–Sharp mass defined by $M(t,r)\coloneqq R(1-\nabla_{\mu}R\,\nabla^{\mu}R)/2$ [29, 41, 42, 8].
@@
 6.
-In order to identify the future-trapped ($\theta<0$ and $\theta^{\prime}<0$) and past-trapped ($\theta>0$ and $\theta^{\prime}>0$) regions we also show directions of two independent null vector fields $l$ and $l^{\prime}$ together with the contour curves of the areal radius in Figs.
+To identify the future-trapped ($\theta<0$ and $\theta^{\prime}<0$) and past-trapped ($\theta>0$ and $\theta^{\prime}>0$) regions we also show directions of two independent null vector fields $l$ and $l^{\prime}$ together with the contour curves of the areal radius in Figs.
 7 and 8.
 Comparing the directions of the null vector fields and contour curves of the areal radius, we can realize that the red and blue shaded regions in Fig.
-6 are future-trapped and past-trapped regions, respectively.
+6 are future- and past-trapped regions, respectively.
 More specifically, in the future(past)-trapped region, both arrows direct to directions along which the areal radius decreases (increases), and the curve described by red points indicates the boundary satisfying $\theta\theta^{\prime}=0$.
-Then, we can identify the outermost boundary surface of the future-trapped region as an apparent horizon.
-The cosmological horizon would be the outermost boundary between the past-trapped and non-trapped regions.
-Still, the following section will present a more concrete classification of trapping horizons.
-(a) $\mu=0.5$
-(b) $\mu=1.2$
+Then, we refer to a marginally future trapped surface that is the outer boundary of the future-trapped region on the $t=$const hypersurface as a BH apparent horizon.
+Similarly, we refer to a marginally past-trapped surface that is the inner boundary of the past-trapped region on the $t=$ const hypersurface as a cosmological apparent horizon.
+It is not easy to understand the horizon configuration.
+Therefore, we carefully investigate it and present a more concrete classification of trapping horizons in the following sections.
+Readers who would like to get a quick intuition about the horizon configuration may refer to the schematic conformal diagram in Figures 16 and 17.
 (a) $\mu=0.5$
@@
 For the $\mu=1.8$ case, the horizon is formed at around $t=24t_{H}$ (right).
-(a) $\mu=0.5$
-(b) $\mu=1.2$
 (a) $\mu=0.5$
@@
 6.
-$R$-contour with 1.0 increments starting from $R=1.0$ and increasing from left bottom to right top is plotted as black curves.
+The $R$ contour is plotted as black curves with 1.0 increments starting from $R=1.0$ and increasing from left bottom to right top.
 In addition to the original plot of $R$-levels with 1.0 increments, it is also plotted with 0.1 increments from $R=8.1$ to $8.9$.
 3.3 Spacetime structures of type I/II PBH
-Let us introduce the classification of trapping horizons following Ref.
+Let us introduce the classification of trapping horizons following Refs.
 [42, 43].
 For this purpose, to avoid confusion, we introduce another notation for the two radial null vector fields and associated expansions as $l_{\pm}$ and $\theta_{\pm}$ for a specific trapping horizon.
-Then we always set $\theta_{+}=0$ when we consider a trapping horizon satisfying $\theta_{+}\theta_{-}=0$.
+Then we always set $\theta_{+}=0$ when considering a trapping horizon satisfying $\theta_{+}\theta_{-}=0$.
 The classification of trapping horizons can be summarized as Table 1.
@@
 Table 1: Classification of trapping horizons satisfying $\theta_{+}\theta_{-}=0$ [42].
-A black hole, white hole, and cosmological horizons with $w=1/3$ $(w=0)$ are identified with a future outer, past outer, and past degenerate (past inner) trapping horizons, respectively.
-(a) Type I PBH
+A black hole, white hole, and cosmological trapping horizons with $w=1/3$ $(w=0)$ are identified with a future-outer, past-outer, and past-degenerate (past-inner) trapping horizons, respectively.
 (a) Type I PBH
 (b) Type II PBH
-Figure 9: Sketch of type I and II PBHs in radiation domination.
+Figure 9: The conformal diagrams of type I and II PBHs in radiation domination inferred from the present numerical results.
 The left figure defines a type I PBH as a causal structure of the typical case of PBH.
 The right one includes bifurcating trapping horizons and defines a type II PBH.
+The red and blue curves depict the future and past trapping horizons, respectively.
+The bifurcating trapping horizons ($\theta_{+}=\theta_{-}=0$) are depicted as circles in the type II case.
 Unlike the dust case discussed in Appendix A, we can only obtain numerical solutions such as those shown in Figs.
@@
 9 as possible conformal diagrams of type I and type II PBHs, respectively.
-The critical difference between the dust and the radiation cases is that trapping horizons in the FLRW regions are timelike and null for the former and the latter, respectively.
-See Figs.
-16, 17 and 19 and Ref.
-[41].
-Then, we confirm that the outermost boundary surface of the future-trapped region is the future outer trapping horizon, which describes an apparent horizon.
-In addition, one can find that the outermost boundary surface between the past-trapped and non-trapped regions describes the cosmological horizon, which asymptotically approaches the past degenerate trapping horizon in the background FLRW universe.
 The crucial difference between the right and left panels in Fig.
@@
 [1] did not clearly distinguish the definitions of type I/II PBHs and type I/II fluctuations.
-However, in our simulation with $w=1/3$, we find that the horizon configuration is of type I for the $\mu=1.6$ case, but the initial fluctuation is of type II nevertheless.
+However, in our simulation with $w=1/3$, we find that the horizon configuration is of type I for the $\mu=1.6$ case, but the initial fluctuation is of type II.
 Therefore, there is a gap between the definitions of type I/II PBHs and type I/II fluctuations.
 This fact is our motivation for the new definition of the PBH.
+The critical difference between the dust and the radiation cases is that trapping horizons in the FLRW regions are timelike and null for the former and the latter, respectively.
+See Figs.
+15, 16 and 18 and Ref.
+[43].
+Then, we confirm that the outermost boundary surface of the future-trapped region is the future-outer trapping horizon, which describes a BH apparent horizon.
+In addition, the outermost boundary surface between the past-trapped and non-trapped regions describes a cosmological apparent horizon, which asymptotically approaches the past degenerate trapping horizon in the background FLRW universe.
 It is worth noting that the schematic diagram given by Fig.
@@
 […]
-Here, we evaluate the mass of the PBH in two different ways at each characteristic time.
-The one is to calculate the Misner-Sharp mass inside the apparent horizon at its first appearance with the future-trapped region (see Fig.
-6).
-We refer to the mass at this time as the initial black hole mass.
-The apparent horizon appears as a future outer trapping horizon in a type I PBH, while it appears as a bifurcating outer trapping horizon in type II.
-We also calculate the final PBH mass by estimating the effect of the accretion.
-It is known that the accretion process is well described by the mass accretion model, often called Novikov-Zeldovich or Bondi accretion [2, 42].
-In this accretion model, the time evolution of the mass is given by
 […]
@@
 (4.1) is not valid at the time of formation of the trapping horizon since it neglects the cosmological expansion as Carr and Hawking (1974) [4] already pointed out.
-But it will correctly describe the accretion into the PBHs for sufficiently large times once we have a quasi-stationary flow into the PBH.
+However, it will correctly describe the accretion into the PBHs for sufficiently large times once we have a quasi-stationary flow into the PBH.
 The integrated solution is,
@@
 4.2 PBH mass with the initial fluctuation motivated by the three-zone model
+In order to construct a specific initial profile that can be contrasted with the former Gaussian profile, let us try to utilize the three-zone profile as a reference, which is an extension of the model used in Ref.
+[1].
 To realize a similar situation to the schematic conformal diagram, let us consider the closed FLRW region for the initial overdense region.
@@
 The value of $l$ is fixed as $k^{-1}=10H_{\rm b}^{-1}$ as before, and the value of $\chi_{a}$ is set as $\pi\tilde{\mu}/k$ for a given value of $\tilde{\mu}$.
-Figure 12: Functional form of $\zeta_{\text{cFLRW}}$ for overdense region $\chi\leq\chi_{a}$.
-This profile is obtained by imposing a boundary condition that $\zeta=0$ at the edge of the overdense region $\chi=\chi_{a}$ [1].
-The initial profiles of $\zeta_{\rm cFLRW}$, $\zeta$, $R$ and the compaction function $\mathcal{C}_{\rm SS}$ are shown in Figs.
-12, 13 and 14.
+The initial profiles of $\zeta$, $R$ and the compaction function $\mathcal{C}_{\rm SS}$ are shown in Figs.
+12 and 13.
+The tail profile is different from the original three-zone model, but it is similar to the three-zone profile around the peak of $\zeta$.
 This profile also exhibits the type II feature with the neck structure for a large amplitude $(\tilde{\mu}\gtrsim 0.85)$.
 Figure 12: Initial three-zone profile of $\zeta(r)$ (left) and areal radius $R(r)$ (right) with respect to the amplitude parameter $\tilde{\mu}$.
-By taking the amplitude parameter $\tilde{\mu}$ larger, a type II fluctuation appeared for $\tilde{\mu}$ larger than $0.80$ with the neck $\partial_{r}R=0$.
+Taking the amplitude parameter $\tilde{\mu}$ larger, a type II fluctuation appeared for $\tilde{\mu}$ larger than $0.80$ with the neck $\partial_{r}R=0$.
+The original profile near the peak $\zeta_{\rm cFLRW}$ [1] is plotted in Fig.
+19 in Appendix B.
 Figure 13: Compaction function $\mathcal{C}_{\text{SS}}(r)$ for thnon-linearprofile.
-The horizontal axis is the coordinate radius with scale-up $z/L$ (left)/ the areal radius $R$ (right).By taking the amplitude $\tilde{\mu}$ larger, the number of peaks becomes two for $\tilde{\mu}$ = 0.89, corresponding to type II fluctuation.
+The horizontal axis is the coordinate radius with scale-up $z/L$ (left)/ the areal radius $R$ (right).
+By taking the amplitude $\tilde{\mu}$ larger, the number of peaks becomes two for $\tilde{\mu}$ = 0.89, corresponding to type II fluctuation.
 The behavior of the horizon configuration is similar to the case of the Gaussian profile, and we do not explicitly show it.
 In contrast to the Gaussian profile, the initial PBH mass formed by this profile decreases as the amplitude of fluctuation $\tilde{\mu}$ increases for $\tilde{\mu}>0.65$ as shown in Fig.
-15, even though they are classified as type I fluctuations and formations for $\tilde{\mu}\leq 0.88$.
+14, even though they are classified as type I fluctuations and formations for $\tilde{\mu}\leq 0.88$.
 This result, as well as the behavior of the Gaussian profile, indicates that whether the initial PBH mass increases or decreases with the amplitude parameter is not only up to its type but also strongly dependent on its profile, together with the ambiguity in quantifying the amplitude of non-linear perturbation.
 We also find the gap between the parameter region of type I/II fluctuations and type I/II PBHs, as indicated in Fig.
-15.
+14.
 Figure 14: Initial PBH mass with respect to the amplitude of fluctuation for the three-zone profile.
@@
 2, while it decreases with $\tilde{\mu}$ in the case of the specific profile (4.4) as the right one of Fig.
-14.
+13.
 5 Conclusion
@@
 Specifically, if there is a bifurcating trapping horizon, the PBH is classified into type II, and type I otherwise.
-In the dust case, KHW reported that a type I fluctuation results in a type I PBH and a type II fluctuation does in a type II PBH.
-However, in the radiation fluid case, this is not true.
-We found a gap between the classifications in terms of the shape of the initial spatial geometry (fluctuation type) and the structure of trapping horizons (formation type).
+In the dust case, KHW reported that a type I fluctuation results in a type I PBH, and a type II fluctuation does in a type II PBH.
+However, in the case of radiation fluids, this is not true.
+We found a gap between the classifications regarding the shape of the initial spatial geometry (fluctuation type) and the structure of trapping horizons (formation type).
 That is, we found that a type II fluctuation may result in a type I PBH.
-Therefore, we propose classifying PBHs in terms of the existence of the bifurcating trapping horizon independently of the type of the initial fluctuation.
+Therefore, we propose classifying PBHs based on the existence of the bifurcating trapping horizon independently of the type of the initial fluctuation.
 In the leading order of the long-wavelength limit, the overdense region can be roughly approximated by the closed FLRW universe, and the amplitude of the density perturbation can be translated into the value of the radial coordinate at the edge of the closed FLRW region.
@@
 Appendix A Conformal diagrams of the three-zone model for the dust case
-In the three-zone model [5], the spacetime of PBH is described by the patchwork of the three different regions : the innermost overdense region given by a spatially closed FLRW universe, the intermediate underdense region, and the outermost flat FLRW universe as the background.
+In the three-zone model [7], the spacetime of PBH is described by the patchwork of the three different regions: the innermost overdense region is given by a spatially closed FLRW universe, the intermediate underdense region, and the outermost flat FLRW universe as the background.
 We also note that the intermediate underdense region must be introduced to compensate for the mass excess of the innermost region and connect the spacetime to the background spacetime without a mass gap.
@@
 Furthermore, for simplicity, we focus on the model where the underdense region is described by the Schwarzschild solution, where the matching is possible without any singular hypersurface for the dust case.
-In spite of this great simplification, the model provides us with a very powerful tool for capturing the overall dynamics of PBH and the structure of trapping horizons.
+Despite this great simplification, the model provides a very powerful tool for capturing the overall dynamics of PBH and the structure of trapping horizons.
 Note that one can also approach the same problem by using the Lemaître-Tolman-Bondi exact dust solution (c.f.
@@
 The conformal diagrams for the three spacetimes of which the three-zone model consists are shown in Fig.
-16, where future $(\theta_{-};\mathcal{L}_{-}\theta_{+})|_{\theta_{+}=0}=(-;*)$, past (+;*)(+;*)( + ; * ), and bifurcating $(0;*)$ trapping horizons are denoted by red lines, blue lines, and the intersection point between them in the Schwarzschild region, respectively.
-For the dust case $w=0$, the BH horizon and cosmological horizon become future outer (;)(-;-)( - ; - ) and past inner (+;+)(+;+)( + ; + ), denoted by red null and blue timelike lines, respectively.
+15, where future $(\theta_{-};\mathcal{L}_{-}\theta_{+})|_{\theta_{+}=0}=(-;*)$, past (+;)(+;*)( + ; ∗ ), and bifurcating $(0;*)$ trapping horizons are denoted by red lines, blue lines, and the intersection point between them in the Schwarzschild region, respectively.
+For the dust case $w=0$, the BH apparent horizon and cosmological apparent horizon become future-outer (;)(-;-)( - ; - ) and past-inner (+;+)(+;+)( + ; + ), denoted by red -solid and blue-dashed lines, respectively.
 First, let us construct the spacetime diagram of the case of the type I fluctuation.
 The extracted regions for the patchwork in the type I case are described by the yellow shaded regions in Fig.
-16, and the resultant spacetime diagram is shown in the left panel of Fig.
-18.
-In these conformal diagrams, we define the right and left directed future null vectors as $l_{\rm out}$ and $l_{\rm in}$, respectively.
+15, and the resultant spacetime diagram is shown in the left panel of Fig.
+17.
+These conformal diagrams define the right and left-directed future null vectors as $l_{\rm out}$ and $l_{\rm in}$, respectively.
 For the closed FLRW region, on the boundary of the extracted region, we can find $\theta_{\rm out}>0$ and $\theta_{\rm in}<0$ at the maximum expansion.
@@
 Similarly, one can understand the matching of the Schwarzsdhild region and the background spatially flat FLRW spacetime.
-For the type II fluctuation, the extracted regions are yellow-shaded in Fig.
-17.
+The extracted regions are yellow-shaded for the type II fluctuation in Fig.
+16.
 Since the type II fluctuation should have a maximal value of the areal radius, the boundary of the closed FLRW region must be outside the central point at which the areal radius takes the maximum value corresponding to the area of the big sphere at the maximum expansion.
 Then we find $\theta_{\rm out}<0$ and $\theta_{\rm in}>0$ on the boundary, which can be matched with the Schwarzschild region only in the left part of the Schwarzschild conformal diagram.
-Thus, the resultant conformal diagram for the type II is given by the right panel of Fig.
-18.
-(a) closed FLRW for dust
-(b) Schwarzschild
+Thus, the resultant conformal diagram for type II is given by the right panel of Fig.
+17.
 (a) closed FLRW for dust
@@
 Red(blue)-shaded regions represent future(past) trapped regions.
-Future outer $(\theta_{-};\mathcal{L}_{-}\theta_{+})|_{\theta_{+}=0}=(-;-)$, past outer (+;)(+;-)( + ; - ), future inner (;+)(-;+)( - ; + ), and past inner (+;+)(+;+)( + ; + ) trapping horizons are denoted by red null, blue null, red timelike, and blue timelike lines, respectively.
-(a) closed FLRW for dust
-(b) Schwarzschild
+Future-outer $(\theta_{-};\mathcal{L}_{-}\theta_{+})|_{\theta_{+}=0}=(-;-)$, past-outer (+;)(+;-)( + ; - ), future-inner (;+)(-;+)( - ; + ), and past-inner (+;+)(+;+)( + ; + ) trapping horizons are denoted by red-solid, blue -solid, red -dashed, and blue -dashed lines, respectively.
 (a) closed FLRW for dust
@@
 (a) Type I PBH
-(a) Type I PBH
 (b) Type II PBH
 Figure 17: Conformal diagrams of PBH for dust case $w=0$ as a patchwork of the three diagrams and trapping horizons.
-The left figure defines a type I PBH as a causal structure of typical case PBH.
+The left figure defines a type I PBH as a typical case PBH causal structure.
 The right one is constructed by changing the cutting diagram range and defining a type II PBH.
 (a) closed FLRW for w=1/3
-(a) closed FLRW for w=1/3
 (b) flat FLRW for w=1/3
-Figure 19: Conformal diagrams of the closed and flat FLRW solutions for the radiation fluid case $w=1/3$. In this case, all trapping horizons are null.
+Figure 18: Conformal diagrams of the closed and flat FLRW solutions for the radiation fluid case $w=1/3$.
+In this case, all trapping horizons are null.
+Future-degenerate $(\theta_{-};\mathcal{L}_{-}\theta_{+})|_{\theta_{+}=0}=(-;0)$ and past-degenerate $(+;0)$ trapping horizons are denoted by red-dotted and blue-dotted lines, respectively.
 Appendix B Derivation of $\zeta_{\text{cFLRW}}$ in radial coordinate $r$
@@
 $\zeta=2\ln\frac{\cos k\chi_{a}/2}{\sqrt{\cos^{4}k\chi_{a}/2+\frac{1}{4}k^{2}r^{2}}}.$ (B.12)
+This functional form is shown as:
+Figure 19: Functional form of $\zeta_{\text{cFLRW}}$ for overdense region $\chi\leq\chi_{a}$.
+This profile is obtained by imposing a boundary condition that $\zeta=0$ at the edge of the overdense region $\chi=\chi_{a}$ [1].
+Appendix C Double-peaked compaction function and averaged density perturbation in the type II fluctuation
+For type II fluctuations, let us assume that $R^{\prime}=\partial_{r}R$, where $R\coloneqq are^{\zeta}$ is the areal radius, satisfies the following condition
+$R^{\prime}\begin{cases}>0&(0<r<r_{m1},r_{m2}<r)\\ =0&(r=r_{m1},r_{m2})\\ <0&(r_{m1}<r<r_{m2})\end{cases}.$ (C.1)
+Since
+$R^{\prime}=ae^{\zeta}(1+r\zeta^{\prime}),$ (C.2)
+we find $r\zeta^{\prime}>-1$ for $0<r<r_{m1}$, $r_{m2}<r$, $r\zeta^{\prime}=-1$ for $r=r_{m1},r_{m2}$, and $r\zeta^{\prime}<-1$ for $r_{m1}<r<r_{m2}$.
+From the above assumption, we can deduce that $r\zeta^{\prime}$ must have at least one minimum below $-1$ with $(r\zeta^{\prime})^{\prime}=0$ between $r_{m1}$ and $r_{m2}$.
+For simplicity, we additionally assume within this section that $r\zeta^{\prime}$ takes only one minimum at $r=r_{m3}$ between $r_{m1}$ and $r_{m2}$ and $r\zeta^{\prime}$ monotonically decreases for $0<r<r_{m3}$ and monotonically increases for $r_{m3}<r$.
+Since we have the compaction function 777 One can rewrite $C_{\rm SS}(r)$ as $C_{\rm SS}(r)=-r\zeta^{\prime}\quantity(1+\frac{1}{2}r\zeta^{\prime})=\frac{3}{4}\delta_{\rm l}\quantity(1-\frac{3}{8}\delta_{\rm l})=-\frac{9}{32}\left(\delta_{l}-\frac{4}{3}\right)^{2}+\frac{1}{2},$ where $\delta_{\rm l}\coloneqq-\frac{4}{3}r\zeta^{\prime}.$ is the ”linear-order density perturbation”.
+Therefore, $C_{\rm SS}$ as a function of $\delta_{\rm l}$ takes two zeroes at $\delta_{\rm l}=0$ and $\delta_{\rm l}=8/3$ and a maximum $1/2$ at $\delta_{\rm l}=4/3$.
+The type II fluctuation is characterized by the existence of the interval in $r$ with $\delta_{l}(r)>4/3$, where $C_{\rm SS}(r)$ as a function of $r$ takes two maxima $1/2$ at its boundaries with $\delta_{l}(r)=4/3$.
+$\mathcal{C}_{\rm SS}(r)=\frac{1}{2}\quantity(1-\quantity(1+r\zeta^{\prime})^{2})$ (C.3)
+and its derivative
+$\mathcal{C}_{\rm SS}^{\prime}(r)=-\quantity(1+r\zeta^{\prime})\quantity(\zeta^{\prime}+r\zeta^{\prime\prime}),$ (C.4)
+we find that the compaction function $\mathcal{C}_{\rm SS}(r)$ has three extrema at $r=r_{m1}$, $r_{m3}$ and $r_{m2}$ with $\mathcal{C}_{\rm SS}^{\prime}(r)=0$, two of which at $r=r_{m1}$ and $r_{m2}$ correspond to two peaks with the value $1/2$, and the rest at $r=r_{m3}$ to its minimum below $1/2$.
+This feature is typically seen for type II fluctuations and actually the case for the Gaussian-shaped profile Eq.
+(2.5) as seen in Fig.
+20.
+Refer to captionRefer to caption
+Figure 20: The situation of type II fluctuation $r\zeta^{\prime}<-1$ from non-monotonicity of the areal radius $R$, $R^{\prime}<0$, and corresponding to the two peaks of the compaction function $\mathcal{C}_{\rm SS}$ (or averaged density perturbation $\bar{\delta}$ at the horizon entry).
+The density perturbation $\delta\coloneqq\rho/\rho_{\rm b}-1$ in a sub-leading order ($\order{\epsilon^{2}}$ ) of the long-wavelength solutions for uniform Hubble (constant mean curvature) slice is written as [21],
+$\delta\simeq-\frac{4}{3}e^{-5\zeta/2}\bar{\Delta}\quantity(e^{\zeta/2})\quantity(\frac{1}{aH_{\rm b}})^{2},$ (C.5)
+where $\bar{\Delta}$ is the Laplacian of the flat background geometry. The compaction function can be identified with the volume averaged density perturbation at the horizon entry $R=1/H_{\rm b}$,
+$\left.\bar{\delta}\right|_{H_{\rm b}=1/R}\simeq 2\mathcal{C}_{\rm SS}.$ (C.6)
+Then, the plot of $\mathcal{C}_{\rm SS}$ in Fig.
+20 can also be understood as $\bar{\delta}$ up to the factor of $2$.
+For the type II fluctuations ($\mu\gtrsim 1.4$), the value of $\bar{\delta}$ takes a minimal value at $r=\sqrt{2}/k\simeq 0.14$ with $k=10$, and the minimal value decreases with increasing $\mu$.
+It is worth noting that the separate universe (pinch-off) limit will not occur for this Gaussian-shaped profile of $\zeta$, while Kopp et al.
+considered the limit based on the other profile via the closed FLRW (or Three-zone) geometry of Eq.
+(4.3) (or Eq.(4.4)).

### v4

@@
 Abstract
-This study investigates the formation of primordial black holes (PBHs) resulting from extremely large amplitudes of initial fluctuations in a radiation-dominated universe.
+This study investigates the formation of type II primordial black holes (PBHs) resulting from extremely large amplitudes of initial fluctuations in a radiation-dominated universe.
 We find that, for a sufficiently large initial amplitude, the configuration of trapping horizons shows characteristic structure due to the existence of bifurcating trapping horizons.
-We call this structure of the trapping horizons “Type II PBH”, while the structure without a bifurcating trapping horizon “Type I PBH”, which is typically generated from a relatively small amplitude of the initial fluctuation.
+We call this type of configuration of the trapping horizons type II-B PBH, while the structure without a bifurcating trapping horizon type II-A PBH.
 In Ref.
-[1], in the dust-dominated universe, the Type II PBH can be realized by the Type II initial fluctuation, which is characterized by a non-monotonic areal radius as a function of the radial coordinate (throat structure) in contrast with the standard case with a monotonic areal radius (Type I fluctuation).
-Our research reveals that a type II fluctuation does not necessarily result in a type II PBH in the case of the radiation fluid.
-We also find that for the initial amplitude well above the threshold value, the resulting PBH mass may either increase or decrease with the initial amplitude depending on its specific profile rather than its fluctuation type.
+[1], in the dust-dominated universe, the type B PBH can be realized by the type II initial fluctuation, which is characterized by a non-monotonic areal radius as a function of the radial coordinate (throat structure) in contrast with the standard case, type A PBH with a monotonic areal radius (type I fluctuation).
+Our research reveals that a type II fluctuation does not necessarily result in a type B PBH in the radiation-dominated case.
+We also find that for an initial amplitude well above the threshold value, the resulting PBH mass may either increase or decrease with the initial amplitude, depending on its specific profile rather than its fluctuation type.
 1 Introduction
 Primordial black holes (PBHs) play a crucial role in exploring the early universe, particularly in investigating inhomogeneities whose scales are much smaller than those of the fluctuations probed by the cosmological microwave background.
-Since PBHs are remnants of non-linear inhomogeneity in the primordial universe following the inflationary era, they are unique probes of statistical properties of the small-scale primordial inhomogeneity.
+Since PBHs are remnants of nonlinear inhomogeneity in the primordial universe following the inflationary era, they are unique probes of statistical properties of the small-scale primordial inhomogeneity.
 They are considered as one of the candidates for dark matter and can be detected as black hole (BH) binaries by gravitational wave interferometers.
@@
 [7, 8]).
-A criterion for the collapse has been firstly proposed by Carr [9] as $\delta\gtrsim w=1/3$, where $\delta$ is the amplitude of the density perturbation, and $w$ is the coefficient of the assumed linear equation of state $p=w\rho$ with $p$ and $\rho$ being the pressure and the energy density, respectively.
-A more accurate threshold value has been explored by numerically analyzing general relativistic non-linear dynamics [10, 11, 12, 7, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23].
-In cosmological perturbation theory, the growth of the perturbation is often described by the evolution of the curvature perturbation $\zeta$ which represents the growing adiabatic modes.
+A criterion for the collapse has been first proposed by Carr [9] as $\delta\gtrsim w=1/3$, where $\delta$ is the amplitude of the density perturbation, and $w$ is the coefficient of the assumed linear equation of state $p=w\rho$ with $p$ and $\rho$ being the pressure and the energy density, respectively.
+A more accurate threshold value has been explored by numerically analyzing general relativistic nonlinear dynamics [10, 11, 12, 7, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23].
+In cosmological perturbation theory, the growth of the perturbation is often described by the evolution of the curvature perturbation $\zeta$, which represents the growing adiabatic modes.
 To perform numerical simulations of PBH formation, one must establish the initial conditions for the spacetime geometry and the fluid in the system.
@@
 In a spherically symmetric system, we introduce a type II primordial fluctuation characterized by a non-monotonic behavior of the areal radius $R=\sqrt{A/4\pi}$, where $A$ is the area of the 2-sphere as a function of the radial coordinate $r$ with constant $t$.
-These fluctuations exhibit distinct neck-like structure, with a minimal areal radius at $\partial_{r}R=0$.
+These fluctuations exhibit distinct neck-like structures, with a minimal areal radius at $\partial_{r}R=0$.
 In cosmological perturbation theory, the areal radius $R$ is often expressed via the curvature fluctuation $\zeta$, $R\propto re^{\zeta}$.
@@
 They investigated the causal structure of the spacetime generated from a specific profile of the type II fluctuation.
-They explicitly showed that it describes a spacetime of BH formation, which they call ‘type II PBH.’ In addition, they revealed distinct properties of the type II PBH from ‘type I PBH,’ which arises from typical type I fluctuations with the amplitude over the threshold.
-Carr and Harada (2015) [24] analytically discussed type II PBHs with more general equations of state using simplified models.
+They explicitly showed that it describes a spacetime of BH formation, which they call type II PBH.
+In addition, they revealed distinct properties of the type II PBH from the type I PBH, which arises from typical type I fluctuations with the amplitude above the threshold.
+Carr and Harada [24] analytically discussed type II PBHs with more general equations of state using simplified models.
 Type I/II fluctuations are also important because the threshold of these two cases gives a possible maximum value of the averaged density perturbation at horizon entry in the simplified models [1, 24].
-We aim to investigate the dynamics and the spacetime structure generated from type II fluctuations in the radiation-dominated universe, namely, to clarify the impact of the existence of the radiation pressure.
-When we consider the abundance of PBHs, the contribution from type II fluctuations is often neglected because they are highly unlikely to be generated statistically, and the abundance of PBHs originating from them is typically suppressed.
-However, as mentioned in Ref.
-[25], they can be significant when the statistics of primordial fluctuations are highly non-Gaussian.
+We aim to investigate the dynamics and the spacetime structure of type II PBH formation in the radiation-dominated universe, namely, to clarify the impact of the existence of the radiation pressure.
+When we consider the abundance of PBHs, the contribution from type II fluctuations is often neglected because they are highly unlikely to be generated statistically, and the abundance of type II PBHs originating from them is typically suppressed.
+However, as mentioned in [25], they can be significant when the statistics of primordial fluctuations are highly non-Gaussian.
 Moreover, in the context of a single-field inflationary model, it has been recently discovered in Ref.
@@
 If one would like to follow the time evolution of the type II fluctuations via a usual areal radial coordinate system such as in Ref.
-[21], the neck throat structure would be seriously difficult to deal with.
+[21], the neck throat structure would be complicated.
 In the Misner–Sharp formalism [29] (see also Refs.
@@
 In Section 3.1, we present the simulation results, highlighting the time evolution of fluctuations.
-Then, we explore the formation of trapping horizons, including the BH horizon in Section 3.2, and give an interpretation of the resultant trapping horizon configurations in Section 3.3.
+Then, we explore the formation of trapping horizons, including the BH horizon in Section 3.2, and give an interpretation of the resulting trapping horizon configurations in Section 3.3.
 In Section 4, we examine the evolution of PBH mass.
-In Section 4.2, to check the profile dependence of the PBH mass, we explore PBH formation from another initial profile that models the overdense region, as outlined by KHW.
+In Section 4.2, to check the profile dependence of the PBH mass, we explore PBH formation from another initial profile that models the overdense region, as described by KHW.
 Finally, we offer our conclusions in the last section.
@@
 Appendix C proves the appearance of two peaks in the compaction function and the averaged density perturbation for the type II fluctuation.
-Throughout the paper, we use the geometrical unit in which the speed of light and the gravitational Newton constant are unity, i.e., $c=G=1$.
+Throughout the paper, we use the geometrical unit in which the speed of light and the gravitational Newton constant are unity, that is, $c=G=1$.
 2 Initial data setting
@@
 We obtain the initial condition by applying the cosmological long-wavelength approximation, in which the small parameter $\epsilon:=k/(aH_{\rm b})$ is introduced111 One would use the other expression, $\epsilon:=1/(H_{\rm b}R_{\rm m})$ with $R_{\rm m}\propto a/k$ being the areal radius, where $\mathcal{C}_{\rm SS}(r)$ takes a maximum.
-, and the zero-th order spatial metric is given by [35]
+, and the zeroth order spatial metric is given by [35]
 $\gamma_{ij}=a^{2}(t)e^{2\zeta(x^{k})}f_{ij},\,$ (2.1)
@@
 In this paper, we initially fix the gauge by imposing the constant mean curvature slice (uniform Hubble slice) and the normal threading, for which the trace of the extrinsic curvature of the time slice is uniformly given by $-3H_{\rm b}$ and the shift vector is equal to zero.
-In a spherically symmetric system, the line element at the zero-th order in powers of $\epsilon$ is denoted as
+In a spherically symmetric system, the line element at the zeroth order in powers of $\epsilon$ is denoted as
 $ds^{2}=-dt^{2}+a^{2}(t)e^{2\zeta(r)}\quantity(dr^{2}+r^{2}d\Omega^{2}),$ (2.2)
@@
 1.
-We call this structure neck or throat structure, and KHW defined such fluctuations as of type II and distinguished them from type I fluctuations, which have a monotonic areal radius [1].
+We call this structure the neck or throat structure, and KHW defined such fluctuations as type II and distinguished them from type I fluctuations, which have a monotonic areal radius [1].
 Figure 1: Initial profile of the curvature fluctuation $\zeta$ (left) and areal radius $R$ as functions of the scale-up coordinate radius $z/L$.
@@
 [36] and references therein for details including its relations to legitimate compaction functions in different gauges and physical interpretations).
-The outer peak of $\mathcal{C}_{\text{SS}}$ corresponds to the neck in the type II fluctuation, the radius at which the areal radius takes the minimum value considered relevant for the PBH formation.
-We will give a general proof for the existence of two peaks and the minimum between them for the compaction function in the type II fluctuation in Appendix C.
+The outer peak of $\mathcal{C}_{\text{SS}}$ corresponds to the neck in the type II fluctuation, the radius at which the areal radius takes the minimum value considered relevant for PBH formation.
+We will give a general proof for two peaks and the minimum between them for the compaction function in the type II fluctuation in the Appendix C.
 We focus on the two peak configurations associated with type II fluctuations (see also Ref.
@@
 By using the position of the peak of $\mathcal{C}_{\text{SS}}(r)$, $r_{\text{m}}=\sqrt{2}/k$, we analytically obtain the threshold value $\mu_{\rm c,II}=e/2=1.359\cdots$ for fluctuations of type II.444 Here, we ignore the window function $W(r)$ to estimate the location $r_{\text{m}}=\sqrt{2}/k$ of the peak of compaction function $\mathcal{C}_{\rm SS}(r)$ since this acts on $\zeta(r)$ only near the outer boundary, $r=r_{\rm w}=0.8L$ For $\mu>\mu_{\rm c,II}$ the compaction function takes the maximum value of $0.5$ at least one location with the radial coordinate, and the number of peak positions becomes two, $r_{m1}$ and $r_{m2}$ with $r_{m1}<r_{m2}$.
-This implies that the type II fluctuations always reach the threshold for collapse and form PBHs irrespectively on the initial profile of $\zeta(r)$.
-The isotropic coordinate $r$ does not always give a geometrical size of the enclosed region and may often lead to misunderstanding.
+This implies that type II fluctuations always reach the collapse threshold and form PBHs irrespectively of the initial $\zeta(r)$ profile.
+The isotropic coordinate $r$ does not always give the geometrical size of the enclosed region, which can often lead to misunderstanding.
 Therefore, we also explicitly show the compaction function $\mathcal{C}_{\rm SS}$ as a function of the areal radius $R$ in the right panel of Fig.
@@
 [11].
-However, we could not avoid the constraint violation at the outer boundary, and this propagates into the bulk region.
+However, we could not avoid violating the constraint at the outer boundary, which propagates into the bulk region.
 Therefore, in the following analyses, we terminate the time evolution before the violation reaches the central region relevant for the analyses of PBH formation.
 We also find the constraint violation near the center for cases where an apparent horizon forms.
-Nevertheless, since we find both independent future-directed null directions are inward there, the constraint violation does not propagate outward.
-In practice, to avoid a crash of numerical calculation, we excise several grids inside the apparent horizon if needed.
-We take the coordinate size $L$ of the numerical region ($0\leq z\leq L$ or, equivalently, $0\leq r\leq L$) as the unit length scale and set the initial value of the scale factor to unity.
-We fix the ratio between the background initial Hubble length $H_{\rm b}^{-1}$ and the typical scale $k^{-1}$ of the initial fluctuation as $k^{-1}=10H_{\text{b}}^{-1}$ which ensure the long-wavelength approximation.
-For a longer duration of the time evolution, we need to take a larger region than the initial Hubble length to avoid the propagation of the constraint violation.
+Nevertheless, since we find both independent future-directed null directions inward, the constraint violation does not propagate outward.
+In practice, we excise several grids inside the apparent horizon to avoid a numerical calculation crash if needed.
+We take the coordinate size $L$ of the numerical region ($0\leq z\leq L$ or equivalently $0\leq r\leq L$) as the unit length scale and set the initial value of the scale factor to unity.
+We fix the ratio between the initial Hubble background length $H_{\rm b}^{-1}$ and the typical scale $k^{-1}$ of the initial fluctuation as $k^{-1}=10H_{\text{b}}^{-1}$ that ensures the long-wavelength approximation.
+To avoid propagating the constraint violation, we need to take a larger region than the initial Hubble length for a longer duration of the time evolution.
 In the following, we take $H_{\text{b}}^{-1}=0.01L$ or $H_{\text{b}}^{-1}=0.005L$ depending on the duration of the required time evolution.
 The horizon entry time $t_{H}$ is defined as $k^{-1}=(aH_{\text{b}})^{-1}|_{t=t_{H}}$ throughout this paper.
-3 Spacetime structure and a new definition of type I/II PBHs
+3 Spacetime structure and definition of type A/B PBHs
 3.1 Time evolution of type I/II fluctuations
@@
 The latter case (middle) indicates a typical gravitational collapse, for which the lapse is decreasing to a very small value as plotted in Fig.
-3 as Ref.
+3 as in Ref.
 [40].
-For a type II fluctuation $\mu=1.8$ case (right), the collapsing behavior happens similarly to the $\mu=1.2$ (middle) case in the plots of density and lapse.
-Refer to captionRefer to caption
+For a type II fluctuation case $\mu=1.8$ (right), the collapsing behavior occurs similarly to the $\mu=1.2$ (middle) case in the density and lapse plots.
 (a) $\mu=0.5$
-Refer to captionRefer to caption
 (b) $\mu=1.2$
-Refer to captionRefer to caption
 (c) $\mu=1.8$
@@
 To check the convergence of the simulation, we plot the spatial average value of the Hamiltonian constraint555 First, we evaluate the appropriately normalized value of the constraint violation $\mathcal{H}_{p}$ at each grid point labeled by $p$.
-Then the average is taken as $\sum_{p=1}^{N}\mathcal{H}_{p}/N$, where $N$ is the total number of the grid points.
+Then the average is taken as $\sum_{p=1}^{N}\mathcal{H}_{p}/N$, where $N$ is the total number of the grid.
 [21](Fig.
@@
 3.2 Identification of BH apparent horizon
-First, let us introduce two independent future-directed radial null vector fields, $l$ and $l^{\prime}$.
+First, we introduce two independent future-directed radial null vector fields, $l$ and $l^{\prime}$.
 We also define the null expansion for $l$ and $l^{\prime}$ as $\theta$ and $\theta^{\prime}$.
@@
 For $\mu=0.5$, the intersection between $\quantity(2M/R)(r)$ and the horizontal line of $12M/R=$ 1 is always only one.
-On the other hand, for the cases $\mu=1.2$ and $1.8$, the line intersects the $12M/R=$ 1 more than three times at a sufficiently late time.
+On the other hand, for the cases $\mu=1.2$ and $1.8$, the line intersects $12M/R=$ 1 more than three times at a sufficiently late time.
 The worldlines of the intersections are shown in Fig.
 6.
-To identify the future-trapped ($\theta<0$ and $\theta^{\prime}<0$) and past-trapped ($\theta>0$ and $\theta^{\prime}>0$) regions we also show directions of two independent null vector fields $l$ and $l^{\prime}$ together with the contour curves of the areal radius in Figs.
+To identify future-trapped ($\theta<0$ and $\theta^{\prime}<0$) and past-trapped ($\theta>0$ and $\theta^{\prime}>0$) regions, we also show directions of two independent null vector fields $l$ and $l^{\prime}$ together with the contour curves of the areal radius in Figs.
 7 and 8.
-Comparing the directions of the null vector fields and contour curves of the areal radius, we can realize that the red and blue shaded regions in Fig.
+Comparing the directions of the null vector fields and the contour curves of the areal radius, we can realize that the red and blue shaded regions in Fig.
 6 are future- and past-trapped regions, respectively.
 More specifically, in the future(past)-trapped region, both arrows direct to directions along which the areal radius decreases (increases), and the curve described by red points indicates the boundary satisfying $\theta\theta^{\prime}=0$.
-Then, we refer to a marginally future trapped surface that is the outer boundary of the future-trapped region on the $t=$const hypersurface as a BH apparent horizon.
+Then, we refer to a marginally future-trapped surface that is the outer boundary of the future-trapped region on the $t=$const hypersurface as a BH apparent horizon.
 Similarly, we refer to a marginally past-trapped surface that is the inner boundary of the past-trapped region on the $t=$ const hypersurface as a cosmological apparent horizon.
@@
 Therefore, we carefully investigate it and present a more concrete classification of trapping horizons in the following sections.
-Readers who would like to get a quick intuition about the horizon configuration may refer to the schematic conformal diagram in Figures 16 and 17.
+Readers who would like to get a quick intuition about the horizon configuration can refer to the schematic conformal diagram in Figures 16 and 17.
 (a) $\mu=0.5$
@@
 6.
-$R$-contour with 1.0 increments starting from $R=1.0$ and increasing from left bottom to right top is plotted as black curves.
+The $R$ contour is plotted as black curves with 1.0 increments starting from $R=1.0$ and increasing from left bottom to right top.
 Figure 8: Directions of two independent null vector fields $l$ and $l^{\prime}$ are depicted as orange and blue arrows for $\mu=1.8$ case (type II) in the right panel of Fig.
@@
 The $R$ contour is plotted as black curves with 1.0 increments starting from $R=1.0$ and increasing from left bottom to right top.
-In addition to the original plot of $R$-levels with 1.0 increments, it is also plotted with 0.1 increments from $R=8.1$ to $8.9$.
-3.3 Spacetime structures of type I/II PBH
+In addition to the original plot of $R$ levels with increments of 1.0, it is also plotted with increments of 0.1 from $R=8.1$ to $8.9$.
+3.3 Spacetime structures of type A/B PBH
 Let us introduce the classification of trapping horizons following Refs.
@@
 Then we always set $\theta_{+}=0$ when considering a trapping horizon satisfying $\theta_{+}\theta_{-}=0$.
-The classification of trapping horizons can be summarized as Table 1.
+The classification of the trapping horizons can be summarized in Table 1.
 $\theta_{+}=0$ + 0 -
@@
 Table 1: Classification of trapping horizons satisfying $\theta_{+}\theta_{-}=0$ [42].
-A black hole, white hole, and cosmological trapping horizons with $w=1/3$ $(w=0)$ are identified with a future-outer, past-outer, and past-degenerate (past-inner) trapping horizons, respectively.
-(a) Type I PBH
-(b) Type II PBH
-Figure 9: The conformal diagrams of type I and II PBHs in radiation domination inferred from the present numerical results.
-The left figure defines a type I PBH as a causal structure of the typical case of PBH.
-The right one includes bifurcating trapping horizons and defines a type II PBH.
+A black hole, a white hole, and cosmological trapping horizons with $w=1/3$ $(w=0)$ are identified with future-outer, past-outer, and past-degenerate (past-inner) trapping horizons, respectively.
+(a) Type IA PBH
+(b) Type IIB PBH
+Figure 9: The present numerical results are inferred from the conformal diagrams of type A and B PBHs in radiation domination.
+The left figure defines a type A PBH as a causal structure of the typical case of PBH.
+The right one includes bifurcating trapping horizons and defines a type B PBH.
 The red and blue curves depict the future and past trapping horizons, respectively.
-The bifurcating trapping horizons ($\theta_{+}=\theta_{-}=0$) are depicted as circles in the type II case.
+The bifurcating trapping horizons ($\theta_{+}=\theta_{-}=0$) are depicted as circles in the type B case.
 Unlike the dust case discussed in Appendix A, we can only obtain numerical solutions such as those shown in Figs.
 7 and 8 for the radiation case.
-From these figures and carefully comparing the similarities and differences from the exact conformal diagrams of the dust solution discussed in Appendix A, we can deduce possible conformal diagrams and proper identifications of the trapping horizons for the radiation case as shown in the left and right panels in Fig.
-9 as possible conformal diagrams of type I and type II PBHs, respectively.
+From these figures and carefully comparing the similarities and differences from the exact conformal diagrams of the dust solution discussed in the Appendix A, we can deduce possible conformal diagrams and proper identifications of the trapping horizons for the radiation case as shown in the left and right panels in Fig.
+9 as possible conformal diagrams of type A and type B PBHs, respectively.
 The crucial difference between the right and left panels in Fig.
@@
 8.
-In this paper, we propose to use this difference to classify PBH formation processes into types I and II; that is, we identify the PBH as of type I if there is no bifurcating trapping horizon and of type II if there appears at least one bifurcating trapping horizon during the time evolution associated with the intersection point.
-This difference has already been pointed out by KHW in Ref.
+In this paper, we propose to use this difference to classify the PBH formation processes into types A and B; that is, we identify the PBH as type A if there is no bifurcating trapping horizon and as type B if at least one bifurcating trapping horizon appears during the time evolution associated with the intersection point.
+KHW has already pointed out this difference in Ref.
 [1] for the case of the dust fluid.
 Unlike the current proposal, the authors in Ref.
-[1] did not clearly distinguish the definitions of type I/II PBHs and type I/II fluctuations.
-However, in our simulation with $w=1/3$, we find that the horizon configuration is of type I for the $\mu=1.6$ case, but the initial fluctuation is of type II.
-Therefore, there is a gap between the definitions of type I/II PBHs and type I/II fluctuations.
+[1] identified the definitions of type I/II and type A/B.
+However, in our simulation with $w=1/3$, we find the case of type IIA, the horizon configuration is of type A from type II fluctuation for the $\mu=1.6$ case.
 This fact is our motivation for the new definition of the PBH.
@@
 It is worth noting that the schematic diagram given by Fig.
-9 implies that all light rays emitted from the center reach not the future null infinity but the future singularity in the type II PBH.
+9 implies that all light rays emitted from the center reach not the future null infinity but the future singularity of the type B PBH.
 Although we cannot make any general statement for the moment, the null trajectories indicated by the orange arrows in Fig.
 8 seem consistent with this observation.
-Precisely, one could also consider the critical case between type I and type II PBHs when the two bifurcating trapping horizons degenerate, then we additionally classify this case as a type IId PBH, which has just one bifurcating trapping horizon.
 4 PBH mass
@@
 (4.1) is not valid at the time of formation of the trapping horizon since it neglects the cosmological expansion as Carr and Hawking (1974) [4] already pointed out.
-However, it will correctly describe the accretion into the PBHs for sufficiently large times once we have a quasi-stationary flow into the PBH.
+However, once we have a quasi-stationary flow into the PBH, it will correctly describe the accretion into the PBHs for sufficiently long times.
 The integrated solution is,
@@
 2.
-In terms of formation types, on the other hand, $\mu=1.2$ to $1.6$ are classified as type I, while $\mu=1.8$ to $2.2$ as type II, as seen in Figs.
+In terms of formation types, on the other hand, $\mu=1.2$ to $1.6$ are classified as type A, while $\mu=1.8$ to $2.2$ as type B, as seen in Figs.
 6 and 9.
-The initial and final PBH masses get larger with increasing amplitude $\mu$. We will check the profile dependence in the next subsection.
+The initial and final PBH masses get larger with increasing amplitude $\mu$. We will check the dependence of the profile in the next subsection.
 4.2 PBH mass with the initial fluctuation motivated by the three-zone model
-In order to construct a specific initial profile that can be contrasted with the former Gaussian profile, let us try to utilize the three-zone profile as a reference, which is an extension of the model used in Ref.
+To construct a specific initial profile that can be contrasted with the former Gaussian profile, let us try to utilize the three-zone profile as a reference, an extension of the model used in Ref.
 [1].
-To realize a similar situation to the schematic conformal diagram, let us consider the closed FLRW region for the initial overdense region.
+Let us consider the closed FLRW region for the initial overdense region to realize a situation similar to the schematic conformal diagram.
 Then, as is shown in the Appendix B, under the boundary condition $\zeta=0$ at the surface of the overdense region, we obtain the following form of $\zeta(r)$:
@@
 This result, as well as the behavior of the Gaussian profile, indicates that whether the initial PBH mass increases or decreases with the amplitude parameter is not only up to its type but also strongly dependent on its profile, together with the ambiguity in quantifying the amplitude of non-linear perturbation.
-We also find the gap between the parameter region of type I/II fluctuations and type I/II PBHs, as indicated in Fig.
+We also find the gap between the parameter region of type I/II and type A/B PBHs, as indicated in Fig.
 14.
@@
 5 Conclusion
-We have numerically investigated the PBH formation from a type II initial fluctuation $\zeta(r)$, which is defined by the non-monotonicity of the areal radius $R(r)\propto re^{\zeta}$ with respect to the coordinate radius $r$ on the spatial geometry, i.e., the existence of $\partial_{r}R(r)=0$ (called neck or throat structure), in the radiation-dominated situation.
+We have numerically investigated the type II PBH formation from a type II initial fluctuation $\zeta(r)$, which is defined by the non-monotonicity of the areal radius $R(r)\propto re^{\zeta}$ with respect to the coordinate radius $r$ on the spatial geometry, i.e., the existence of $\partial_{r}R(r)=0$ (called neck or throat structure), in the radiation-dominated situation.
 As in the dust case reported in KHW, the PBH formation process can be classified into types I and II according to the configuration of the trapped regions $R(t,r)\leq 2M(t,r)$ (or trapping horizons $R(t,r)=2M(t,r)$).
-Specifically, if there is a bifurcating trapping horizon, the PBH is classified into type II, and type I otherwise.
-In the dust case, KHW reported that a type I fluctuation results in a type I PBH, and a type II fluctuation does in a type II PBH.
+Specifically, if there is a bifurcating trapping horizon, the PBH is classified as type B, and otherwise, it is classified as type A.
+In the dust case, KHW reported that a type I fluctuation results in a type A PBH and a type II fluctuation in a type B PBH.
 However, in the case of radiation fluids, this is not true.
-We found a gap between the classifications regarding the shape of the initial spatial geometry (fluctuation type) and the structure of trapping horizons (formation type).
-That is, we found that a type II fluctuation may result in a type I PBH.
-Therefore, we propose classifying PBHs based on the existence of the bifurcating trapping horizon independently of the type of the initial fluctuation.
+We found a gap between the classifications regarding the shape of the initial spatial geometry (fluctuation type) and the structure of the trapping horizons (formation type).
+We found that a type II fluctuation may result in a type I PBH, so-called type II-A PBH.
+Therefore, we propose classifying type A/B PBHs based on the existence of the bifurcating trapping horizon independently of the type of the initial fluctuation.
 In the leading order of the long-wavelength limit, the overdense region can be roughly approximated by the closed FLRW universe, and the amplitude of the density perturbation can be translated into the value of the radial coordinate at the edge of the closed FLRW region.
 Then, the type II fluctuation indicates that the edge is located in the other hemisphere of $S^{3}$ initial hypersurface beyond the great circle.
-Therefore, in the case of type II PBHs, one may expect that the PBH mass would decrease with increasing the amplitude of initial fluctuation since the initial time slice of $S^{3}$ in the overdensity region closes in the limit of large-density perturbation, which indicates the area of the edge is vanishing.
+Therefore, in the case of type II PBHs, one may expect that the PBH mass would decrease with increasing the amplitude of initial fluctuation since the initial time slice of $S^{3}$ in the overdensity region closes in the limit of large-density perturbation, which indicates that the area of the edge is vanishing.
 However, this is not the case for the Gaussian-type initial fluctuation profile, while the other profile exhibits decreasing mass with increasing initial amplitude.
 Therefore, we conclude that the behavior of the initial PBH mass depends on not only the type of fluctuation but also the fluctuation profile.
-Thus, systematic investigation of the profile dependence would be needed, e.g., by using the shape parameter defined as the curvature peak of the compaction function $\mathcal{C}_{\text{SS}}$ [19, 45].
-In addition to the profile dependence, the dependence on the equation of state $p=w\rho$ is also an interesting issue, which fills the gap between KHW and this work [20].
+Thus, a systematic investigation of the profile dependence would be needed, e.g., by using the shape parameter defined as the curvature peak of the compaction function $\mathcal{C}_{\text{SS}}$ [19, 45].
+In addition to profile dependence, dependence on the equation of state $p=w\rho$ is also an interesting issue, filling the gap between KHW and this work [20].
 Acknowledgments
@@
 acknowledges support from the JSPS Postdoctoral Fellowships for Research in Japan (Graduate School of Sciences, Nagoya University).
-T.H.
-is supported in part by JSPS KAKENHI Grant Nos.
-19K03876 and 20H05853.
-D.S.
-would like to take this opportunity to thank the “Nagoya University Interdisciplinary Frontier Fellowship” supported by Nagoya University and JST, the establishment of university fellowships towards the creation of science technology innovation, Grant Number JPMJFS2120.
-C.Y.
-is supported in part by JSPS KAKENHI Grant Nos.
-20H05850 and 20H05853.
-C.Y.
-and T.H.
-acknowledge the hospitality at APCTP during the focus research program ”Black Hole and Gravitational Waves: from modified theories of gravity to data analysis”, where part of this work was discussed.
+The authors are supported in part by the JSPS KAKENHI Grant Nos.
+19K03876(TH), 20H05850(CY), 20H05853(TH, CY), and 24KJ1223(DS).
+KU would like to take this opportunity to thank the “Nagoya University Interdisciplinary Frontier Fellowship” supported by Nagoya University and JST, the establishment of university fellowships towards the creation of science technology innovation, Grant Number JPMJFS2125.
+CY and TH acknowledge the hospitality at APCTP during the focus research program ”Black Hole and Gravitational Waves: from modified theories of gravity to data analysis”, where part of this work was discussed.
 Appendix A Conformal diagrams of the three-zone model for the dust case
 In the three-zone model [7], the spacetime of PBH is described by the patchwork of the three different regions: the innermost overdense region is given by a spatially closed FLRW universe, the intermediate underdense region, and the outermost flat FLRW universe as the background.
-We also note that the intermediate underdense region must be introduced to compensate for the mass excess of the innermost region and connect the spacetime to the background spacetime without a mass gap.
+We also note that the intermediate underdense region must be introduced to compensate for the innermost region’s mass excess and connect the spacetime to the background spacetime without a mass gap.
 Without the intermediate underdense region, this model cannot be the exact solution, even for the dust case.
-For the underdense region with a perfect fluid with $w=1/3$, it would not be tractable to obtain any analytical and exact solutions of such dynamical inhomogeneous spacetimes with non-linear sound waves.
+For the underdense region with a perfect fluid with $w=1/3$, obtaining any analytical and exact solutions of such dynamical inhomogeneous spacetimes with non-linear sound waves would not be tractable.
 Therefore, we stick to the three-zone model with dust.
 Furthermore, for simplicity, we focus on the model where the underdense region is described by the Schwarzschild solution, where the matching is possible without any singular hypersurface for the dust case.
-Despite this great simplification, the model provides a very powerful tool for capturing the overall dynamics of PBH and the structure of trapping horizons.
+Despite this significant simplification, the model provides a powerful tool for capturing the overall dynamics of PBH and the structure of trapping horizons.
 Note that one can also approach the same problem by using the Lemaître-Tolman-Bondi exact dust solution (c.f.
@@
 15, where future $(\theta_{-};\mathcal{L}_{-}\theta_{+})|_{\theta_{+}=0}=(-;*)$, past (+;)(+;*)( + ; ∗ ), and bifurcating $(0;*)$ trapping horizons are denoted by red lines, blue lines, and the intersection point between them in the Schwarzschild region, respectively.
-For the dust case $w=0$, the BH apparent horizon and cosmological apparent horizon become future-outer (;)(-;-)( - ; - ) and past-inner (+;+)(+;+)( + ; + ), denoted by red -solid and blue-dashed lines, respectively.
+For the dust case $w=0$, the apparent horizon and cosmological apparent horizon become future-outer (;)(-;-)( - ; - ) and past-inner (+;+)(+;+)( + ; + ), denoted by red-solid and blue-dashed lines, respectively.
 First, let us construct the spacetime diagram of the case of the type I fluctuation.
-The extracted regions for the patchwork in the type I case are described by the yellow shaded regions in Fig.
+The extracted regions for the patchwork in the type I (type A) case are described by the yellow shaded regions in Fig.
 15, and the resultant spacetime diagram is shown in the left panel of Fig.
@@
 Then we find $\theta_{\rm out}<0$ and $\theta_{\rm in}>0$ on the boundary, which can be matched with the Schwarzschild region only in the left part of the Schwarzschild conformal diagram.
-Thus, the resultant conformal diagram for type II is given by the right panel of Fig.
+Thus, the resultant conformal diagram for type II (type B) is given by the right panel of Fig.
 17.
@@
 (c) flat FLRW for dust
-Figure 15: Conformal diagrams to use for a patchwork for the spacetime of the type I PBH schematically in Fig.
+Figure 15: Conformal diagrams to use for a patchwork for the spacetime of the type I(type A) PBH schematically in Fig.
 17 for dust case $w=0$.
@@
 (c) flat FLRW for dust
-Figure 16: Same as Fig. 15 but for the type II PBH. We use the yellow-shaded regions for a patchwork.
+Figure 16: Same as Fig. 15 but for the type II(type B) PBH. We use the yellow-shaded regions for a patchwork.
 (a) Type I PBH
@@
 Figure 17: Conformal diagrams of PBH for dust case $w=0$ as a patchwork of the three diagrams and trapping horizons.
-The left figure defines a type I PBH as a typical case PBH causal structure.
-The right one is constructed by changing the cutting diagram range and defining a type II PBH.
+The left figure defines a type I(type A) PBH as a typical case PBH causal structure.
+The right one is constructed by changing the cutting diagram range and defining a type II(type B) PBH.
 (a) closed FLRW for w=1/3
@@
 20.
-Refer to captionRefer to caption
 Figure 20: The situation of type II fluctuation $r\zeta^{\prime}<-1$ from non-monotonicity of the areal radius $R$, $R^{\prime}<0$, and corresponding to the two peaks of the compaction function $\mathcal{C}_{\rm SS}$ (or averaged density perturbation $\bar{\delta}$ at the horizon entry).

### v5

@@
 Abstract
-This study investigates the formation of type II primordial black holes (PBHs) resulting from extremely large amplitudes of initial fluctuations in a radiation-dominated universe.
+This study investigates the formation of primordial black holes (PBHs) resulting from extremely large amplitudes of initial fluctuations in a radiation-dominated universe.
 We find that, for a sufficiently large initial amplitude, the configuration of trapping horizons shows characteristic structure due to the existence of bifurcating trapping horizons.
-We call this type of configuration of the trapping horizons type II-B PBH, while the structure without a bifurcating trapping horizon type II-A PBH.
+We call this structure of the trapping horizons “Type II PBH”, while the structure without a bifurcating trapping horizon “Type I PBH”, which is typically generated from a relatively small amplitude of the initial fluctuation.
 In Ref.
-[1], in the dust-dominated universe, the type B PBH can be realized by the type II initial fluctuation, which is characterized by a non-monotonic areal radius as a function of the radial coordinate (throat structure) in contrast with the standard case, type A PBH with a monotonic areal radius (type I fluctuation).
-Our research reveals that a type II fluctuation does not necessarily result in a type B PBH in the radiation-dominated case.
-We also find that for an initial amplitude well above the threshold value, the resulting PBH mass may either increase or decrease with the initial amplitude, depending on its specific profile rather than its fluctuation type.
+[1], in the dust-dominated universe, the Type II PBH can be realized by the Type II initial fluctuation, which is characterized by a non-monotonic areal radius as a function of the radial coordinate (throat structure) in contrast with the standard case with a monotonic areal radius (Type I fluctuation).
+Our research reveals that a type II fluctuation does not necessarily result in a type II PBH in the case of the radiation fluid.
+We also find that for the initial amplitude well above the threshold value, the resulting PBH mass may either increase or decrease with the initial amplitude depending on its specific profile rather than its fluctuation type.
 1 Introduction
 Primordial black holes (PBHs) play a crucial role in exploring the early universe, particularly in investigating inhomogeneities whose scales are much smaller than those of the fluctuations probed by the cosmological microwave background.
-Since PBHs are remnants of nonlinear inhomogeneity in the primordial universe following the inflationary era, they are unique probes of statistical properties of the small-scale primordial inhomogeneity.
+Since PBHs are remnants of non-linear inhomogeneity in the primordial universe following the inflationary era, they are unique probes of statistical properties of the small-scale primordial inhomogeneity.
 They are considered as one of the candidates for dark matter and can be detected as black hole (BH) binaries by gravitational wave interferometers.
@@
 [7, 8]).
-A criterion for the collapse has been first proposed by Carr [9] as $\delta\gtrsim w=1/3$, where $\delta$ is the amplitude of the density perturbation, and $w$ is the coefficient of the assumed linear equation of state $p=w\rho$ with $p$ and $\rho$ being the pressure and the energy density, respectively.
-A more accurate threshold value has been explored by numerically analyzing general relativistic nonlinear dynamics [10, 11, 12, 7, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23].
-In cosmological perturbation theory, the growth of the perturbation is often described by the evolution of the curvature perturbation $\zeta$, which represents the growing adiabatic modes.
+A criterion for the collapse has been firstly proposed by Carr [9] as $\delta\gtrsim w=1/3$, where $\delta$ is the amplitude of the density perturbation, and $w$ is the coefficient of the assumed linear equation of state $p=w\rho$ with $p$ and $\rho$ being the pressure and the energy density, respectively.
+A more accurate threshold value has been explored by numerically analyzing general relativistic non-linear dynamics [10, 11, 12, 7, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23].
+In cosmological perturbation theory, the growth of the perturbation is often described by the evolution of the curvature perturbation $\zeta$ which represents the growing adiabatic modes.
 To perform numerical simulations of PBH formation, one must establish the initial conditions for the spacetime geometry and the fluid in the system.
@@
 In a spherically symmetric system, we introduce a type II primordial fluctuation characterized by a non-monotonic behavior of the areal radius $R=\sqrt{A/4\pi}$, where $A$ is the area of the 2-sphere as a function of the radial coordinate $r$ with constant $t$.
-These fluctuations exhibit distinct neck-like structures, with a minimal areal radius at $\partial_{r}R=0$.
+These fluctuations exhibit distinct neck-like structure, with a minimal areal radius at $\partial_{r}R=0$.
 In cosmological perturbation theory, the areal radius $R$ is often expressed via the curvature fluctuation $\zeta$, $R\propto re^{\zeta}$.
@@
 They investigated the causal structure of the spacetime generated from a specific profile of the type II fluctuation.
-They explicitly showed that it describes a spacetime of BH formation, which they call type II PBH.
-In addition, they revealed distinct properties of the type II PBH from the type I PBH, which arises from typical type I fluctuations with the amplitude above the threshold.
-Carr and Harada [24] analytically discussed type II PBHs with more general equations of state using simplified models.
+They explicitly showed that it describes a spacetime of BH formation, which they call ‘type II PBH.’ In addition, they revealed distinct properties of the type II PBH from ‘type I PBH,’ which arises from typical type I fluctuations with the amplitude over the threshold.
+Carr and Harada (2015) [24] analytically discussed type II PBHs with more general equations of state using simplified models.
 Type I/II fluctuations are also important because the threshold of these two cases gives a possible maximum value of the averaged density perturbation at horizon entry in the simplified models [1, 24].
-We aim to investigate the dynamics and the spacetime structure of type II PBH formation in the radiation-dominated universe, namely, to clarify the impact of the existence of the radiation pressure.
-When we consider the abundance of PBHs, the contribution from type II fluctuations is often neglected because they are highly unlikely to be generated statistically, and the abundance of type II PBHs originating from them is typically suppressed.
-However, as mentioned in [25], they can be significant when the statistics of primordial fluctuations are highly non-Gaussian.
+We aim to investigate the dynamics and the spacetime structure generated from type II fluctuations in the radiation-dominated universe, namely, to clarify the impact of the existence of the radiation pressure.
+When we consider the abundance of PBHs, the contribution from type II fluctuations is often neglected because they are highly unlikely to be generated statistically, and the abundance of PBHs originating from them is typically suppressed.
+However, as mentioned in Ref.
+[25], they can be significant when the statistics of primordial fluctuations are highly non-Gaussian.
 Moreover, in the context of a single-field inflationary model, it has been recently discovered in Ref.
@@
 If one would like to follow the time evolution of the type II fluctuations via a usual areal radial coordinate system such as in Ref.
-[21], the neck throat structure would be complicated.
+[21], the neck throat structure would be seriously difficult to deal with.
 In the Misner–Sharp formalism [29] (see also Refs.
@@
 In Section 3.1, we present the simulation results, highlighting the time evolution of fluctuations.
-Then, we explore the formation of trapping horizons, including the BH horizon in Section 3.2, and give an interpretation of the resulting trapping horizon configurations in Section 3.3.
+Then, we explore the formation of trapping horizons, including the BH horizon in Section 3.2, and give an interpretation of the resultant trapping horizon configurations in Section 3.3.
 In Section 4, we examine the evolution of PBH mass.
-In Section 4.2, to check the profile dependence of the PBH mass, we explore PBH formation from another initial profile that models the overdense region, as described by KHW.
+In Section 4.2, to check the profile dependence of the PBH mass, we explore PBH formation from another initial profile that models the overdense region, as outlined by KHW.
 Finally, we offer our conclusions in the last section.
@@
 Appendix C proves the appearance of two peaks in the compaction function and the averaged density perturbation for the type II fluctuation.
-Throughout the paper, we use the geometrical unit in which the speed of light and the gravitational Newton constant are unity, that is, $c=G=1$.
+Throughout the paper, we use the geometrical unit in which the speed of light and the gravitational Newton constant are unity, i.e., $c=G=1$.
 2 Initial data setting
@@
 We obtain the initial condition by applying the cosmological long-wavelength approximation, in which the small parameter $\epsilon:=k/(aH_{\rm b})$ is introduced111 One would use the other expression, $\epsilon:=1/(H_{\rm b}R_{\rm m})$ with $R_{\rm m}\propto a/k$ being the areal radius, where $\mathcal{C}_{\rm SS}(r)$ takes a maximum.
-, and the zeroth order spatial metric is given by [35]
+, and the zero-th order spatial metric is given by [35]
 $\gamma_{ij}=a^{2}(t)e^{2\zeta(x^{k})}f_{ij},\,$ (2.1)
@@
 In this paper, we initially fix the gauge by imposing the constant mean curvature slice (uniform Hubble slice) and the normal threading, for which the trace of the extrinsic curvature of the time slice is uniformly given by $-3H_{\rm b}$ and the shift vector is equal to zero.
-In a spherically symmetric system, the line element at the zeroth order in powers of $\epsilon$ is denoted as
+In a spherically symmetric system, the line element at the zero-th order in powers of $\epsilon$ is denoted as
 $ds^{2}=-dt^{2}+a^{2}(t)e^{2\zeta(r)}\quantity(dr^{2}+r^{2}d\Omega^{2}),$ (2.2)
@@
 1.
-We call this structure the neck or throat structure, and KHW defined such fluctuations as type II and distinguished them from type I fluctuations, which have a monotonic areal radius [1].
+We call this structure neck or throat structure, and KHW defined such fluctuations as of type II and distinguished them from type I fluctuations, which have a monotonic areal radius [1].
 Figure 1: Initial profile of the curvature fluctuation $\zeta$ (left) and areal radius $R$ as functions of the scale-up coordinate radius $z/L$.
@@
 [36] and references therein for details including its relations to legitimate compaction functions in different gauges and physical interpretations).
-The outer peak of $\mathcal{C}_{\text{SS}}$ corresponds to the neck in the type II fluctuation, the radius at which the areal radius takes the minimum value considered relevant for PBH formation.
-We will give a general proof for two peaks and the minimum between them for the compaction function in the type II fluctuation in the Appendix C.
+The outer peak of $\mathcal{C}_{\text{SS}}$ corresponds to the neck in the type II fluctuation, the radius at which the areal radius takes the minimum value considered relevant for the PBH formation.
+We will give a general proof for the existence of two peaks and the minimum between them for the compaction function in the type II fluctuation in Appendix C.
 We focus on the two peak configurations associated with type II fluctuations (see also Ref.
@@
 By using the position of the peak of $\mathcal{C}_{\text{SS}}(r)$, $r_{\text{m}}=\sqrt{2}/k$, we analytically obtain the threshold value $\mu_{\rm c,II}=e/2=1.359\cdots$ for fluctuations of type II.444 Here, we ignore the window function $W(r)$ to estimate the location $r_{\text{m}}=\sqrt{2}/k$ of the peak of compaction function $\mathcal{C}_{\rm SS}(r)$ since this acts on $\zeta(r)$ only near the outer boundary, $r=r_{\rm w}=0.8L$ For $\mu>\mu_{\rm c,II}$ the compaction function takes the maximum value of $0.5$ at least one location with the radial coordinate, and the number of peak positions becomes two, $r_{m1}$ and $r_{m2}$ with $r_{m1}<r_{m2}$.
-This implies that type II fluctuations always reach the collapse threshold and form PBHs irrespectively of the initial $\zeta(r)$ profile.
-The isotropic coordinate $r$ does not always give the geometrical size of the enclosed region, which can often lead to misunderstanding.
+This implies that the type II fluctuations always reach the threshold for collapse and form PBHs irrespectively on the initial profile of $\zeta(r)$.
+The isotropic coordinate $r$ does not always give a geometrical size of the enclosed region and may often lead to misunderstanding.
 Therefore, we also explicitly show the compaction function $\mathcal{C}_{\rm SS}$ as a function of the areal radius $R$ in the right panel of Fig.
@@
 [11].
-However, we could not avoid violating the constraint at the outer boundary, which propagates into the bulk region.
+However, we could not avoid the constraint violation at the outer boundary, and this propagates into the bulk region.
 Therefore, in the following analyses, we terminate the time evolution before the violation reaches the central region relevant for the analyses of PBH formation.
 We also find the constraint violation near the center for cases where an apparent horizon forms.
-Nevertheless, since we find both independent future-directed null directions inward, the constraint violation does not propagate outward.
-In practice, we excise several grids inside the apparent horizon to avoid a numerical calculation crash if needed.
-We take the coordinate size $L$ of the numerical region ($0\leq z\leq L$ or equivalently $0\leq r\leq L$) as the unit length scale and set the initial value of the scale factor to unity.
-We fix the ratio between the initial Hubble background length $H_{\rm b}^{-1}$ and the typical scale $k^{-1}$ of the initial fluctuation as $k^{-1}=10H_{\text{b}}^{-1}$ that ensures the long-wavelength approximation.
-To avoid propagating the constraint violation, we need to take a larger region than the initial Hubble length for a longer duration of the time evolution.
+Nevertheless, since we find both independent future-directed null directions are inward there, the constraint violation does not propagate outward.
+In practice, to avoid a crash of numerical calculation, we excise several grids inside the apparent horizon if needed.
+We take the coordinate size $L$ of the numerical region ($0\leq z\leq L$ or, equivalently, $0\leq r\leq L$) as the unit length scale and set the initial value of the scale factor to unity.
+We fix the ratio between the background initial Hubble length $H_{\rm b}^{-1}$ and the typical scale $k^{-1}$ of the initial fluctuation as $k^{-1}=10H_{\text{b}}^{-1}$ which ensure the long-wavelength approximation.
+For a longer duration of the time evolution, we need to take a larger region than the initial Hubble length to avoid the propagation of the constraint violation.
 In the following, we take $H_{\text{b}}^{-1}=0.01L$ or $H_{\text{b}}^{-1}=0.005L$ depending on the duration of the required time evolution.
 The horizon entry time $t_{H}$ is defined as $k^{-1}=(aH_{\text{b}})^{-1}|_{t=t_{H}}$ throughout this paper.
-3 Spacetime structure and definition of type A/B PBHs
+3 Spacetime structure and a new definition of type I/II PBHs
 3.1 Time evolution of type I/II fluctuations
@@
 The latter case (middle) indicates a typical gravitational collapse, for which the lapse is decreasing to a very small value as plotted in Fig.
-3 as in Ref.
+3 as Ref.
 [40].
-For a type II fluctuation case $\mu=1.8$ (right), the collapsing behavior occurs similarly to the $\mu=1.2$ (middle) case in the density and lapse plots.
+For a type II fluctuation $\mu=1.8$ case (right), the collapsing behavior happens similarly to the $\mu=1.2$ (middle) case in the plots of density and lapse.
 (a) $\mu=0.5$
@@
 To check the convergence of the simulation, we plot the spatial average value of the Hamiltonian constraint555 First, we evaluate the appropriately normalized value of the constraint violation $\mathcal{H}_{p}$ at each grid point labeled by $p$.
-Then the average is taken as $\sum_{p=1}^{N}\mathcal{H}_{p}/N$, where $N$ is the total number of the grid.
+Then the average is taken as $\sum_{p=1}^{N}\mathcal{H}_{p}/N$, where $N$ is the total number of the grid points.
 [21](Fig.
@@
 3.2 Identification of BH apparent horizon
-First, we introduce two independent future-directed radial null vector fields, $l$ and $l^{\prime}$.
+First, let us introduce two independent future-directed radial null vector fields, $l$ and $l^{\prime}$.
 We also define the null expansion for $l$ and $l^{\prime}$ as $\theta$ and $\theta^{\prime}$.
@@
 For $\mu=0.5$, the intersection between $\quantity(2M/R)(r)$ and the horizontal line of $12M/R=$ 1 is always only one.
-On the other hand, for the cases $\mu=1.2$ and $1.8$, the line intersects $12M/R=$ 1 more than three times at a sufficiently late time.
+On the other hand, for the cases $\mu=1.2$ and $1.8$, the line intersects the $12M/R=$ 1 more than three times at a sufficiently late time.
 The worldlines of the intersections are shown in Fig.
 6.
-To identify future-trapped ($\theta<0$ and $\theta^{\prime}<0$) and past-trapped ($\theta>0$ and $\theta^{\prime}>0$) regions, we also show directions of two independent null vector fields $l$ and $l^{\prime}$ together with the contour curves of the areal radius in Figs.
+To identify the future-trapped ($\theta<0$ and $\theta^{\prime}<0$) and past-trapped ($\theta>0$ and $\theta^{\prime}>0$) regions we also show directions of two independent null vector fields $l$ and $l^{\prime}$ together with the contour curves of the areal radius in Figs.
 7 and 8.
-Comparing the directions of the null vector fields and the contour curves of the areal radius, we can realize that the red and blue shaded regions in Fig.
+Comparing the directions of the null vector fields and contour curves of the areal radius, we can realize that the red and blue shaded regions in Fig.
 6 are future- and past-trapped regions, respectively.
 More specifically, in the future(past)-trapped region, both arrows direct to directions along which the areal radius decreases (increases), and the curve described by red points indicates the boundary satisfying $\theta\theta^{\prime}=0$.
-Then, we refer to a marginally future-trapped surface that is the outer boundary of the future-trapped region on the $t=$const hypersurface as a BH apparent horizon.
+Then, we refer to a marginally future trapped surface that is the outer boundary of the future-trapped region on the $t=$const hypersurface as a BH apparent horizon.
 Similarly, we refer to a marginally past-trapped surface that is the inner boundary of the past-trapped region on the $t=$ const hypersurface as a cosmological apparent horizon.
@@
 Therefore, we carefully investigate it and present a more concrete classification of trapping horizons in the following sections.
-Readers who would like to get a quick intuition about the horizon configuration can refer to the schematic conformal diagram in Figures 16 and 17.
+Readers who would like to get a quick intuition about the horizon configuration may refer to the schematic conformal diagram in Figures 16 and 17.
 (a) $\mu=0.5$
@@
 6.
-The $R$ contour is plotted as black curves with 1.0 increments starting from $R=1.0$ and increasing from left bottom to right top.
+$R$-contour with 1.0 increments starting from $R=1.0$ and increasing from left bottom to right top is plotted as black curves.
 Figure 8: Directions of two independent null vector fields $l$ and $l^{\prime}$ are depicted as orange and blue arrows for $\mu=1.8$ case (type II) in the right panel of Fig.
@@
 The $R$ contour is plotted as black curves with 1.0 increments starting from $R=1.0$ and increasing from left bottom to right top.
-In addition to the original plot of $R$ levels with increments of 1.0, it is also plotted with increments of 0.1 from $R=8.1$ to $8.9$.
-3.3 Spacetime structures of type A/B PBH
+In addition to the original plot of $R$-levels with 1.0 increments, it is also plotted with 0.1 increments from $R=8.1$ to $8.9$.
+3.3 Spacetime structures of type I/II PBH
 Let us introduce the classification of trapping horizons following Refs.
@@
 Then we always set $\theta_{+}=0$ when considering a trapping horizon satisfying $\theta_{+}\theta_{-}=0$.
-The classification of the trapping horizons can be summarized in Table 1.
+The classification of trapping horizons can be summarized as Table 1.
 $\theta_{+}=0$ + 0 -
@@
 Table 1: Classification of trapping horizons satisfying $\theta_{+}\theta_{-}=0$ [42].
-A black hole, a white hole, and cosmological trapping horizons with $w=1/3$ $(w=0)$ are identified with future-outer, past-outer, and past-degenerate (past-inner) trapping horizons, respectively.
-(a) Type IA PBH
-(b) Type IIB PBH
-Figure 9: The present numerical results are inferred from the conformal diagrams of type A and B PBHs in radiation domination.
+A black hole, white hole, and cosmological trapping horizons with $w=1/3$ $(w=0)$ are identified with a future-outer, past-outer, and past-degenerate (past-inner) trapping horizons, respectively.
+(a) Type A PBH
+(b) Type B PBH
+Figure 9: The conformal diagrams of type A and B PBHs in radiation domination inferred from the present numerical results.
 The left figure defines a type A PBH as a causal structure of the typical case of PBH.
@@
 7 and 8 for the radiation case.
-From these figures and carefully comparing the similarities and differences from the exact conformal diagrams of the dust solution discussed in the Appendix A, we can deduce possible conformal diagrams and proper identifications of the trapping horizons for the radiation case as shown in the left and right panels in Fig.
-9 as possible conformal diagrams of type A and type B PBHs, respectively.
+From these figures and carefully comparing the similarities and differences from the exact conformal diagrams of the dust solution discussed in Appendix A, we can deduce possible conformal diagrams and proper identifications of the trapping horizons for the radiation case as shown in the left and right panels in Fig.
+9 as possible conformal diagrams of type I and type II PBHs, respectively.
 The crucial difference between the right and left panels in Fig.
@@
 8.
-In this paper, we propose to use this difference to classify the PBH formation processes into types A and B; that is, we identify the PBH as type A if there is no bifurcating trapping horizon and as type B if at least one bifurcating trapping horizon appears during the time evolution associated with the intersection point.
-KHW has already pointed out this difference in Ref.
+In this paper, we propose to use this difference to classify PBH formation processes into types I and II; that is, we identify the PBH as of type I if there is no bifurcating trapping horizon and of type II if there appears at least one bifurcating trapping horizon during the time evolution associated with the intersection point.
+This difference has already been pointed out by KHW in Ref.
 [1] for the case of the dust fluid.
 Unlike the current proposal, the authors in Ref.
-[1] identified the definitions of type I/II and type A/B.
-However, in our simulation with $w=1/3$, we find the case of type IIA, the horizon configuration is of type A from type II fluctuation for the $\mu=1.6$ case.
+[1] did not clearly distinguish the definitions of type I/II PBHs and type I/II fluctuations.
+However, in our simulation with $w=1/3$, we find that the horizon configuration is of type I for the $\mu=1.6$ case, but the initial fluctuation is of type II.
+Therefore, there is a gap between the definitions of type I/II PBHs and type I/II fluctuations.
 This fact is our motivation for the new definition of the PBH.
@@
 It is worth noting that the schematic diagram given by Fig.
-9 implies that all light rays emitted from the center reach not the future null infinity but the future singularity of the type B PBH.
+9 implies that all light rays emitted from the center reach not the future null infinity but the future singularity in the type II PBH.
 Although we cannot make any general statement for the moment, the null trajectories indicated by the orange arrows in Fig.
 8 seem consistent with this observation.
+Precisely, one could also consider the critical case between type I and type II PBHs when the two bifurcating trapping horizons degenerate, then we additionally classify this case as a type IId PBH, which has just one bifurcating trapping horizon.
 4 PBH mass
@@
 (4.1) is not valid at the time of formation of the trapping horizon since it neglects the cosmological expansion as Carr and Hawking (1974) [4] already pointed out.
-However, once we have a quasi-stationary flow into the PBH, it will correctly describe the accretion into the PBHs for sufficiently long times.
+However, it will correctly describe the accretion into the PBHs for sufficiently large times once we have a quasi-stationary flow into the PBH.
 The integrated solution is,
@@
 2.
-In terms of formation types, on the other hand, $\mu=1.2$ to $1.6$ are classified as type A, while $\mu=1.8$ to $2.2$ as type B, as seen in Figs.
+In terms of formation types, on the other hand, $\mu=1.2$ to $1.6$ are classified as type I, while $\mu=1.8$ to $2.2$ as type II, as seen in Figs.
 6 and 9.
-The initial and final PBH masses get larger with increasing amplitude $\mu$. We will check the dependence of the profile in the next subsection.
+The initial and final PBH masses get larger with increasing amplitude $\mu$. We will check the profile dependence in the next subsection.
 4.2 PBH mass with the initial fluctuation motivated by the three-zone model
-To construct a specific initial profile that can be contrasted with the former Gaussian profile, let us try to utilize the three-zone profile as a reference, an extension of the model used in Ref.
+In order to construct a specific initial profile that can be contrasted with the former Gaussian profile, let us try to utilize the three-zone profile as a reference, which is an extension of the model used in Ref.
 [1].
-Let us consider the closed FLRW region for the initial overdense region to realize a situation similar to the schematic conformal diagram.
+To realize a similar situation to the schematic conformal diagram, let us consider the closed FLRW region for the initial overdense region.
 Then, as is shown in the Appendix B, under the boundary condition $\zeta=0$ at the surface of the overdense region, we obtain the following form of $\zeta(r)$:
@@
 This result, as well as the behavior of the Gaussian profile, indicates that whether the initial PBH mass increases or decreases with the amplitude parameter is not only up to its type but also strongly dependent on its profile, together with the ambiguity in quantifying the amplitude of non-linear perturbation.
-We also find the gap between the parameter region of type I/II and type A/B PBHs, as indicated in Fig.
+We also find the gap between the parameter region of type I/II fluctuations and type I/II PBHs, as indicated in Fig.
 14.
@@
 5 Conclusion
-We have numerically investigated the type II PBH formation from a type II initial fluctuation $\zeta(r)$, which is defined by the non-monotonicity of the areal radius $R(r)\propto re^{\zeta}$ with respect to the coordinate radius $r$ on the spatial geometry, i.e., the existence of $\partial_{r}R(r)=0$ (called neck or throat structure), in the radiation-dominated situation.
+We have numerically investigated the PBH formation from a type II initial fluctuation $\zeta(r)$, which is defined by the non-monotonicity of the areal radius $R(r)\propto re^{\zeta}$ with respect to the coordinate radius $r$ on the spatial geometry, i.e., the existence of $\partial_{r}R(r)=0$ (called neck or throat structure), in the radiation-dominated situation.
 As in the dust case reported in KHW, the PBH formation process can be classified into types I and II according to the configuration of the trapped regions $R(t,r)\leq 2M(t,r)$ (or trapping horizons $R(t,r)=2M(t,r)$).
-Specifically, if there is a bifurcating trapping horizon, the PBH is classified as type B, and otherwise, it is classified as type A.
-In the dust case, KHW reported that a type I fluctuation results in a type A PBH and a type II fluctuation in a type B PBH.
+Specifically, if there is a bifurcating trapping horizon, the PBH is classified into type II, and type I otherwise.
+In the dust case, KHW reported that a type I fluctuation results in a type I PBH, and a type II fluctuation does in a type II PBH.
 However, in the case of radiation fluids, this is not true.
-We found a gap between the classifications regarding the shape of the initial spatial geometry (fluctuation type) and the structure of the trapping horizons (formation type).
-We found that a type II fluctuation may result in a type I PBH, so-called type II-A PBH.
-Therefore, we propose classifying type A/B PBHs based on the existence of the bifurcating trapping horizon independently of the type of the initial fluctuation.
+We found a gap between the classifications regarding the shape of the initial spatial geometry (fluctuation type) and the structure of trapping horizons (formation type).
+That is, we found that a type II fluctuation may result in a type I PBH.
+Therefore, we propose classifying PBHs based on the existence of the bifurcating trapping horizon independently of the type of the initial fluctuation.
 In the leading order of the long-wavelength limit, the overdense region can be roughly approximated by the closed FLRW universe, and the amplitude of the density perturbation can be translated into the value of the radial coordinate at the edge of the closed FLRW region.
 Then, the type II fluctuation indicates that the edge is located in the other hemisphere of $S^{3}$ initial hypersurface beyond the great circle.
-Therefore, in the case of type II PBHs, one may expect that the PBH mass would decrease with increasing the amplitude of initial fluctuation since the initial time slice of $S^{3}$ in the overdensity region closes in the limit of large-density perturbation, which indicates that the area of the edge is vanishing.
+Therefore, in the case of type II PBHs, one may expect that the PBH mass would decrease with increasing the amplitude of initial fluctuation since the initial time slice of $S^{3}$ in the overdensity region closes in the limit of large-density perturbation, which indicates the area of the edge is vanishing.
 However, this is not the case for the Gaussian-type initial fluctuation profile, while the other profile exhibits decreasing mass with increasing initial amplitude.
 Therefore, we conclude that the behavior of the initial PBH mass depends on not only the type of fluctuation but also the fluctuation profile.
-Thus, a systematic investigation of the profile dependence would be needed, e.g., by using the shape parameter defined as the curvature peak of the compaction function $\mathcal{C}_{\text{SS}}$ [19, 45].
-In addition to profile dependence, dependence on the equation of state $p=w\rho$ is also an interesting issue, filling the gap between KHW and this work [20].
+Thus, systematic investigation of the profile dependence would be needed, e.g., by using the shape parameter defined as the curvature peak of the compaction function $\mathcal{C}_{\text{SS}}$ [19, 45].
+In addition to the profile dependence, the dependence on the equation of state $p=w\rho$ is also an interesting issue, which fills the gap between KHW and this work [20].
 Acknowledgments
@@
 acknowledges support from the JSPS Postdoctoral Fellowships for Research in Japan (Graduate School of Sciences, Nagoya University).
-The authors are supported in part by the JSPS KAKENHI Grant Nos.
-19K03876(TH), 20H05850(CY), 20H05853(TH, CY), and 24KJ1223(DS).
-KU would like to take this opportunity to thank the “Nagoya University Interdisciplinary Frontier Fellowship” supported by Nagoya University and JST, the establishment of university fellowships towards the creation of science technology innovation, Grant Number JPMJFS2125.
-CY and TH acknowledge the hospitality at APCTP during the focus research program ”Black Hole and Gravitational Waves: from modified theories of gravity to data analysis”, where part of this work was discussed.
+T.H.
+is supported in part by JSPS KAKENHI Grant Nos.
+19K03876 and 20H05853.
+D.S.
+would like to take this opportunity to thank the “Nagoya University Interdisciplinary Frontier Fellowship” supported by Nagoya University and JST, the establishment of university fellowships towards the creation of science technology innovation, Grant Number JPMJFS2120.
+C.Y.
+is supported in part by JSPS KAKENHI Grant Nos.
+20H05850 and 20H05853.
+C.Y.
+and T.H.
+acknowledge the hospitality at APCTP during the focus research program ”Black Hole and Gravitational Waves: from modified theories of gravity to data analysis”, where part of this work was discussed.
 Appendix A Conformal diagrams of the three-zone model for the dust case
 In the three-zone model [7], the spacetime of PBH is described by the patchwork of the three different regions: the innermost overdense region is given by a spatially closed FLRW universe, the intermediate underdense region, and the outermost flat FLRW universe as the background.
-We also note that the intermediate underdense region must be introduced to compensate for the innermost region’s mass excess and connect the spacetime to the background spacetime without a mass gap.
+We also note that the intermediate underdense region must be introduced to compensate for the mass excess of the innermost region and connect the spacetime to the background spacetime without a mass gap.
 Without the intermediate underdense region, this model cannot be the exact solution, even for the dust case.
-For the underdense region with a perfect fluid with $w=1/3$, obtaining any analytical and exact solutions of such dynamical inhomogeneous spacetimes with non-linear sound waves would not be tractable.
+For the underdense region with a perfect fluid with $w=1/3$, it would not be tractable to obtain any analytical and exact solutions of such dynamical inhomogeneous spacetimes with non-linear sound waves.
 Therefore, we stick to the three-zone model with dust.
 Furthermore, for simplicity, we focus on the model where the underdense region is described by the Schwarzschild solution, where the matching is possible without any singular hypersurface for the dust case.
-Despite this significant simplification, the model provides a powerful tool for capturing the overall dynamics of PBH and the structure of trapping horizons.
+Despite this great simplification, the model provides a very powerful tool for capturing the overall dynamics of PBH and the structure of trapping horizons.
 Note that one can also approach the same problem by using the Lemaître-Tolman-Bondi exact dust solution (c.f.
@@
 15, where future $(\theta_{-};\mathcal{L}_{-}\theta_{+})|_{\theta_{+}=0}=(-;*)$, past (+;)(+;*)( + ; ∗ ), and bifurcating $(0;*)$ trapping horizons are denoted by red lines, blue lines, and the intersection point between them in the Schwarzschild region, respectively.
-For the dust case $w=0$, the apparent horizon and cosmological apparent horizon become future-outer (;)(-;-)( - ; - ) and past-inner (+;+)(+;+)( + ; + ), denoted by red-solid and blue-dashed lines, respectively.
+For the dust case $w=0$, the BH apparent horizon and cosmological apparent horizon become future-outer (;)(-;-)( - ; - ) and past-inner (+;+)(+;+)( + ; + ), denoted by red -solid and blue-dashed lines, respectively.
 First, let us construct the spacetime diagram of the case of the type I fluctuation.
-The extracted regions for the patchwork in the type I (type A) case are described by the yellow shaded regions in Fig.
+The extracted regions for the patchwork in the type I case are described by the yellow shaded regions in Fig.
 15, and the resultant spacetime diagram is shown in the left panel of Fig.
@@
 Then we find $\theta_{\rm out}<0$ and $\theta_{\rm in}>0$ on the boundary, which can be matched with the Schwarzschild region only in the left part of the Schwarzschild conformal diagram.
-Thus, the resultant conformal diagram for type II (type B) is given by the right panel of Fig.
+Thus, the resultant conformal diagram for type II is given by the right panel of Fig.
 17.
@@
 (c) flat FLRW for dust
-Figure 15: Conformal diagrams to use for a patchwork for the spacetime of the type I(type A) PBH schematically in Fig.
+Figure 15: Conformal diagrams to use for a patchwork for the spacetime of the type I PBH schematically in Fig.
 17 for dust case $w=0$.
@@
 (c) flat FLRW for dust
-Figure 16: Same as Fig. 15 but for the type II(type B) PBH. We use the yellow-shaded regions for a patchwork.
+Figure 16: Same as Fig. 15 but for the type II PBH. We use the yellow-shaded regions for a patchwork.
 (a) Type I PBH

### v6

@@
 We find that, for a sufficiently large initial amplitude, the configuration of trapping horizons shows characteristic structure due to the existence of bifurcating trapping horizons.
-We call this structure of the trapping horizons “Type II PBH”, while the structure without a bifurcating trapping horizon “Type I PBH”, which is typically generated from a relatively small amplitude of the initial fluctuation.
-In Ref.
-[1], in the dust-dominated universe, the Type II PBH can be realized by the Type II initial fluctuation, which is characterized by a non-monotonic areal radius as a function of the radial coordinate (throat structure) in contrast with the standard case with a monotonic areal radius (Type I fluctuation).
-Our research reveals that a type II fluctuation does not necessarily result in a type II PBH in the case of the radiation fluid.
-We also find that for the initial amplitude well above the threshold value, the resulting PBH mass may either increase or decrease with the initial amplitude depending on its specific profile rather than its fluctuation type.
+We call this type of configuration of the trapping horizons type B PBH, while the structure without a bifurcating trapping horizon type A PBH.
+As shown in Ref.
+[PhysRevD.83.124025], in the matter-dominated universe, the type B PBH can be realized by the type II initial fluctuation, which is characterized by a non-monotonic areal radius as a function of the radial coordinate (throat structure) in contrast with the standard case, type A PBH with a monotonic areal radius (type I fluctuation).
+Our research reveals that a type II fluctuation does not necessarily result in a type B PBH in the radiation-dominated case.
+We also find that for an initial amplitude well above the threshold value, the resulting PBH mass may either increase or decrease with increasing the initial amplitude, depending on its specific profile rather than its fluctuation type.
 1 Introduction
 Primordial black holes (PBHs) play a crucial role in exploring the early universe, particularly in investigating inhomogeneities whose scales are much smaller than those of the fluctuations probed by the cosmological microwave background.
-Since PBHs are remnants of non-linear inhomogeneity in the primordial universe following the inflationary era, they are unique probes of statistical properties of the small-scale primordial inhomogeneity.
+Since PBHs are remnants of nonlinear inhomogeneity in the primordial universe following the inflationary era, they are unique probes of statistical properties of the small-scale primordial inhomogeneity.
 They are considered as one of the candidates for dark matter and can be detected as black hole (BH) binaries by gravitational wave interferometers.
-According to the standard scenario, PBHs were formed through the gravitational collapse of significantly high-density regions during the radiation-dominated era, which had been generated in the inflationary period [2, 3, 4, 5, 6].
+According to the standard scenario, PBHs were formed through the gravitational collapse of significantly high-density regions during the radiation-dominated era, which had been generated in the inflationary period [Zeldovich:1967lct, 10.1093/mnras/152.1.75, 10.1093/mnras/168.2.399, 1978SvA....22..129N, 1980AZh....57..250N].
 The mechanism of PBH formation can be briefly understood by considering a model of the gravitational collapse in the flat Friedmann–Lemaître–Robertson–Walker (FLRW) universe (see, e.g., Refs.
-[7, 8]).
-A criterion for the collapse has been firstly proposed by Carr [9] as $\delta\gtrsim w=1/3$, where $\delta$ is the amplitude of the density perturbation, and $w$ is the coefficient of the assumed linear equation of state $p=w\rho$ with $p$ and $\rho$ being the pressure and the energy density, respectively.
-A more accurate threshold value has been explored by numerically analyzing general relativistic non-linear dynamics [10, 11, 12, 7, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23].
-In cosmological perturbation theory, the growth of the perturbation is often described by the evolution of the curvature perturbation $\zeta$ which represents the growing adiabatic modes.
+[PhysRevD.88.084051, galaxies10060112]).
+A criterion for the collapse has been first proposed by Carr [osti_4109902] as $\delta\gtrsim w=1/3$, where $\delta$ is the amplitude of the density perturbation, and $w$ is the coefficient of the assumed linear equation of state $p=w\rho$ with $p$ and $\rho$ being the pressure and the energy density, respectively.
+A more accurate threshold value has been explored by numerically analyzing general relativistic nonlinear dynamics [Niemeyer:1999ak, PhysRevD.60.084002, Musco_2005, PhysRevD.88.084051, Nakama_2014, Nakama_2014double, osti_4109902, Polnarev_2007, Musco:2008hv, Musco:2012au, ESCRIVA2020100466, Escriva:2019phb, escriva_2021analy, Harada:2015yda, Musco:2018rwt, Musco:2020jjb].
+In cosmological perturbation theory, the growth of the perturbation is often described by the evolution of the curvature perturbation $\zeta$, which represents the growing adiabatic modes.
 To perform numerical simulations of PBH formation, one must establish the initial conditions for the spacetime geometry and the fluid in the system.
 This involves setting the initial profile of curvature fluctuations at the super-horizon scale before the horizon entry to be consistent with the assumption of the growing adiabatic modes.
-These analytical constructions are typically achieved using the gradient expansion techniques [11, 16, 21].
+These analytical constructions are typically achieved using the gradient expansion techniques [PhysRevD.60.084002, Polnarev_2007, Harada:2015yda].
 In a spherically symmetric system, we introduce a type II primordial fluctuation characterized by a non-monotonic behavior of the areal radius $R=\sqrt{A/4\pi}$, where $A$ is the area of the 2-sphere as a function of the radial coordinate $r$ with constant $t$.
-These fluctuations exhibit distinct neck-like structure, with a minimal areal radius at $\partial_{r}R=0$.
+These fluctuations exhibit distinct neck-like structures, with a minimal areal radius at $\partial_{r}R=0$.
 In cosmological perturbation theory, the areal radius $R$ is often expressed via the curvature fluctuation $\zeta$, $R\propto re^{\zeta}$.
 The type II fluctuation is only realized by a sufficiently large amplitude of $\zeta$ with the scale fixed.
-Kopp, Hofmann, and Weller (KHW) [1] investigated PBH formation associated with the type II fluctuations by employing the Lemaître–Tolman–Bondi (LTB) solution, which is the analytic solution for the dust fluid ($w=0$).
+Kopp, Hofmann, and Weller (KHW) [PhysRevD.83.124025] investigated PBH formation associated with the type II fluctuations by employing the Lemaître–Tolman–Bondi (LTB) solution, which is the analytic solution for the dust fluid ($w=0$).
 They investigated the causal structure of the spacetime generated from a specific profile of the type II fluctuation.
-They explicitly showed that it describes a spacetime of BH formation, which they call ‘type II PBH.’ In addition, they revealed distinct properties of the type II PBH from ‘type I PBH,’ which arises from typical type I fluctuations with the amplitude over the threshold.
-Carr and Harada (2015) [24] analytically discussed type II PBHs with more general equations of state using simplified models.
-Type I/II fluctuations are also important because the threshold of these two cases gives a possible maximum value of the averaged density perturbation at horizon entry in the simplified models [1, 24].
-We aim to investigate the dynamics and the spacetime structure generated from type II fluctuations in the radiation-dominated universe, namely, to clarify the impact of the existence of the radiation pressure.
-When we consider the abundance of PBHs, the contribution from type II fluctuations is often neglected because they are highly unlikely to be generated statistically, and the abundance of PBHs originating from them is typically suppressed.
-However, as mentioned in Ref.
-[25], they can be significant when the statistics of primordial fluctuations are highly non-Gaussian.
+They explicitly showed that it describes a spacetime of BH formation, which they call type II PBH.
+In addition, they revealed distinct properties of the type II PBH from the type I PBH, which arises from typical type I fluctuations with the amplitude above the threshold.
+Carr and Harada [PhysRevD.91.084048] analytically discussed type II PBHs with more general equations of state using simplified models.
+Type I/II fluctuations are also important because the threshold of these two cases gives a possible maximum value of the averaged density perturbation at horizon entry in the simplified models [PhysRevD.83.124025, PhysRevD.91.084048].
+We aim to investigate the dynamics and the spacetime structure of type II PBH formation in the radiation-dominated universe, namely, to clarify the impact of the existence of the radiation pressure.
+When we consider the abundance of PBHs, the contribution from type II fluctuations is often neglected because they are statistically unlikely to be generated, and the abundance of type II PBHs originating from them is typically suppressed.
+However, as mentioned in [Gow:2022jfb], they can be significant when the statistics of primordial fluctuations are highly non-Gaussian.
 Moreover, in the context of a single-field inflationary model, it has been recently discovered in Ref.
-[26] that type II fluctuations surround the bubbles formed when the inflaton overshoots the barrier.
+[escriva2023formation] that type II fluctuations surround the bubbles formed when the inflaton overshoots the barrier.
 As non-Gaussianities increase, the production of vacuum bubbles becomes dominant, leading to a mass function and abundance primarily driven by these fluctuations.
 In addition, a similar BH formation mechanism due to the birth of a baby universe with a tunneling process during inflation has been reported in Refs.
-[27, 28].
+[Deng_2017, Deng_2017b].
 The baby universe is connected to our universe through a throat inside the BH, like the neck structure in the type II fluctuation.
-If one would like to follow the time evolution of the type II fluctuations via a usual areal radial coordinate system such as in Ref.
-[21], the neck throat structure would be seriously difficult to deal with.
-In the Misner–Sharp formalism [29] (see also Refs.
-[1, 30]), the evolution equations suffer from the trouble associated with $/0$ 0 term at the neck radius satisfying $\partial_{r}R=0$ when we consider type II fluctuations.
-To overcome this problem, it is necessary to conveniently use and modify the equations, as done in Ref.
-[27] for another scenario of PBH formation from the collapse of vacuum bubbles.
-On the other hand, in this paper, we use the simulation code developed from a 3+1 dimensional simulation code [31] with BSSN [32, 33] formalism with CARTOON method [34], so that we can avoid the same problem.
+For example, in the Misner–Sharp formalism [PhysRev.136.B571] (see also Refs.
+[PhysRevD.83.124025, PhysRevD.74.084013]), the time evolution of the type II fluctuations via a usual areal radial coordinate system such as in Ref.
+[Harada:2015yda], the neck throat structure would be complicated.
+The evolution equations suffer from the trouble associated with $0/0$ term at the neck radius satisfying $\partial_{r}R=0$ when we consider type II fluctuations.
+To overcome this problem, it will be necessary to conveniently modify the equations, as done in Ref.
+[Deng_2017] for another scenario of PBH formation from the collapse of vacuum bubbles.
+In this paper, we use the simulation code developed from a 3+1 dimensional simulation code [Yoo_2019] with BSSN [shibata1995evolution, PhysRevD.59.024007] formalism with CARTOON method [doi:10.1142/S0218271801000834] so that we can avoid the same problem.
 This paper is organized as follows.
 Section 2 describes the setup and numerical code of spherically symmetric numerical simulation of PBH formation.
-In Section 3.1, we present the simulation results, highlighting the time evolution of fluctuations.
-Then, we explore the formation of trapping horizons, including the BH horizon in Section 3.2, and give an interpretation of the resultant trapping horizon configurations in Section 3.3.
-In Section 4, we examine the evolution of PBH mass.
-In Section 4.2, to check the profile dependence of the PBH mass, we explore PBH formation from another initial profile that models the overdense region, as outlined by KHW.
+In Sec.
+3.1, we present the simulation results, highlighting the time evolution of fluctuations.
+Then, we explore the formation of trapping horizons, including the BH horizon in Sec.
+3.2, and give an interpretation of the resulting trapping horizon configurations in Sec.
+3.3.
+In Sec.
+4, we examine the evolution of PBH mass.
+In Sec.
+4.2, to check the profile dependence of the PBH mass, we explore PBH formation from another initial profile that models the overdense region, as described by KHW.
 Finally, we offer our conclusions in the last section.
@@
 Appendix C proves the appearance of two peaks in the compaction function and the averaged density perturbation for the type II fluctuation.
-Throughout the paper, we use the geometrical unit in which the speed of light and the gravitational Newton constant are unity, i.e., $c=G=1$.
+Throughout the paper, we use the geometrical unit in which the speed of light and the gravitational Newton constant are unity, that is, $c=G=1$.
 2 Initial data setting
 We start our simulation from the time when the length scale of the fluctuation is much larger than the Hubble length $H_{\text{b}}^{-1}$, that is, $k^{-1}\gg(aH_{\text{b}})^{-1}$ with $k$, $a=a(t)$ and $H_{\text{b}}:=(da/dt)/a$ being the characteristic comoving wave number, the scale factor and the Hubble parameter of the background flat FLRW universe, respectively.
-We obtain the initial condition by applying the cosmological long-wavelength approximation, in which the small parameter $\epsilon:=k/(aH_{\rm b})$ is introduced111 One would use the other expression, $\epsilon:=1/(H_{\rm b}R_{\rm m})$ with $R_{\rm m}\propto a/k$ being the areal radius, where $\mathcal{C}_{\rm SS}(r)$ takes a maximum.
-, and the zero-th order spatial metric is given by [35]
+We obtain the initial condition by applying the cosmological long-wavelength approximation, in which the small parameter $\epsilon:=k/(aH_{\rm b})$ is introduced [^1] One would use the other expression, $\epsilon:=1/(H_{\rm b}R_{\rm m})$ with $R_{\rm m}\propto a/k$ being the areal radius, where $\mathcal{C}_{\rm SS}(r)$ takes a maximum.
+, and the zeroth order spatial metric is given by [Lyth_2005]
 $\gamma_{ij}=a^{2}(t)e^{2\zeta(x^{k})}f_{ij},\,$ (2.1)
@@
 We note that the function $\zeta$ is a time-independent arbitrary function of the spatial coordinates $x^{k}$.
-Once we initially set $\zeta$ as a function of the spatial coordinates $x^{i}$ and fix the gauge conditions, we can successively obtain the growing mode solution up to the following leading terms in this expansion [11, 21].
+Once we initially set $\zeta$ as a function of the spatial coordinates $x^{i}$ and fix the gauge conditions, we can successively obtain the growing mode solution up to the following leading terms in this expansion [PhysRevD.60.084002, Harada:2015yda].
 These long-wavelength solutions are used as the initial data for the time evolution.
-In this paper, we initially fix the gauge by imposing the constant mean curvature slice (uniform Hubble slice) and the normal threading, for which the trace of the extrinsic curvature of the time slice is uniformly given by $-3H_{\rm b}$ and the shift vector is equal to zero.
-In a spherically symmetric system, the line element at the zero-th order in powers of $\epsilon$ is denoted as
+In this paper, we initially fix the gauge by imposing the constant mean curvature slice (uniform Hubble slice) and the normal threading, for which the trace of the extrinsic curvature of the time slice is uniformly given by $3H_{\rm b}$ and the shift vector is equal to zero.
+In a spherically symmetric system, the line element at the zeroth order in powers of $\epsilon$ is denoted as
 $ds^{2}=-dt^{2}+a^{2}(t)e^{2\zeta(r)}\quantity(dr^{2}+r^{2}d\Omega^{2}),$ (2.2)
@@
 $r=z-\frac{\eta}{1+\eta}\frac{L}{\pi}\sin\left(\frac{\pi}{L}z\right)$ (2.3)
-with a parameter $\eta=20$ [31]. In this coordinate, the resolution in the central region is $1+2\eta$ times finer than that around the boundary $z=r=L$;
+with a parameter $\eta=20$ [Yoo_2019]. In this coordinate, the resolution in the central region is $1+2\eta$ times finer than that near the boundary $z=r=L$;
 […]
-Up through section 4, we use the following Gaussian-shaped profile in the numerical simulation [31],
+Up through Sec. 4, we use the following Gaussian-shaped profile in the numerical simulation [Yoo_2019],
 $\zeta(r)=\mu e^{-\frac{1}{2}k^{2}r^{2}}W(r),$ (2.5)
-where $\mu$ is the amplitude of the fluctuation and $W(r)$ is a window function [31] introduced to eliminate the tail of the Gaussian function and impose $\zeta=0$ at the outer boundary of the numerical domain.222 The explicit functional form [31] is 11 1$\quad$1 $\text{for}\quad 0\leq r\leq r_{\text{w}},$ $1-\frac{\quantity((r_{\text{w}}-L)^{6}-(L-r)^{6})^{6}}{(r_{\text{w}}-L)^{36}}\quad$ $\text{for}\quad r_{\text{w}}\leq r\leq L,$ (2.6) 00 0$\quad$ $\text{for}\quad L\leq r,$ where $r_{\rm w}$ is set to $0.8L$.
+where $\mu$ is the amplitude of the fluctuation and $W(r)$ is a window function [Yoo_2019] introduced to eliminate the tail of the Gaussian function and impose $\zeta=0$ at the outer boundary of the numerical domain.
+[^2] The explicit functional form [Yoo_2019] is $1\quad$ $\text{for}\quad 0\leq r\leq r_{\text{w}},$ $1-\frac{\quantity((r_{\text{w}}-L)^{6}-(L-r)^{6})^{6}}{(r_{\text{w}}-L)^{36}}\quad$ $\text{for}\quad r_{\text{w}}\leq r\leq L,$ (2.6) $0\quad$ $\text{for}\quad L\leq r,$ where $r_{\rm w}$ is set to $0.8L$.
 Figure 1 shows the functional form of this profile (left) and the areal radius $R$ (right).
-By taking a sufficiently large amplitude of the fluctuation $\mu$, the areal radius $R(r)$ of the perturbed region becomes non-monotonic along the radial coordinate $r$, i.e., a pair of the local minimum and maximum points appears with $\partial_{r}R=0$ for $\mu\gtrsim 1.4$ as in Fig.
+By taking a sufficiently large amplitude of the fluctuation $\mu$, the areal radius $R(r)$ in the overdense region becomes non-monotonic along the radial coordinate $r$, i.e., a pair of the local minimum and maximum points appears with $\partial_{r}R=0$ for $\mu\gtrsim 1.4$ as in Fig.
 1.
-We call this structure neck or throat structure, and KHW defined such fluctuations as of type II and distinguished them from type I fluctuations, which have a monotonic areal radius [1].
-Figure 1: Initial profile of the curvature fluctuation $\zeta$ (left) and areal radius $R$ as functions of the scale-up coordinate radius $z/L$.
-By taking the amplitude parameter $\mu$ larger, a type II fluctuation appears for $\mu=1.4$ and above with the neck $\partial_{r}R=0$.
-The feature of type II fluctuations can also be understood as two peaks333 It is shown in Ref.
-[26] that the appearance of fluctuations of type II naturally appears via a specific single-field inflation model with a bump, but in this case, one of the peaks of $\mathcal{C}_{\rm SS}$ is hidden inside the bubble.
-of the Shibata–Sasaki compaction function $\mathcal{C}_{\text{SS}}$ defined by
+We call this structure the neck or throat structure, and KHW defined such fluctuations as type II and distinguished them from type I fluctuations, which have a monotonic areal radius [PhysRevD.83.124025].
+Figure 1: In the left panel, the vertical and horizontal axes are the functional forms for the curvature perturbation $\zeta(r)$ given by Eq.
+(2.5) and the scale-up coordinate $z/L$ defined in Eq.
+(2.3), respectively.
+Note that $z$ is a monotonically increasing function of $r$ and that $L$ is chosen for the computation region’s length.
+We plot $\zeta(r)$ for the set of the amplitude parameter values $\mu=0.5,1.2,1.4,1.6$ and $1.8$.
+In the right panel, the vertical and horizontal axes are the areal radius $R(r)$ and the scale-up coordinate $z/L$, respectively, for the initial data sets of the long-wavelength solutions generated by $\zeta(r)$ plotted in the left panel.
+We can see that $R(r)$ is a monotonically increasing function for $\mu=0.5$ and $1.2$, while it increases, decreases, and increases again as $r$ increases for $\mu=1.4,1.6$, and $1.8$.
+In the latter cases, $R(r)$ has a maximum and a minimum with $\partial_{r}R=0$, where the minimum corresponds to the neck.
+The feature of type II fluctuations can also be understood as two peaks [^3] It is shown in Ref.
+[escriva2023formation] that the appearance of fluctuations of type II naturally appears via a specific single-field inflation model with a bump, but in this case, one of the peaks of $\mathcal{C}_{\rm SS}$ is hidden inside the bubble.
+of the Shibata–Sasaki compaction function $\mathcal{C}_{\text{SS}}$ at super horizon scales, which is calculated as [Harada:2015yda]
 […]
-This function has been firstly introduced in Ref.
-[11] as an estimator for the criterion of PBH formation, and its usefulness has been recognized through subsequent related works (see Ref.
-[36] and references therein for details including its relations to legitimate compaction functions in different gauges and physical interpretations).
-The outer peak of $\mathcal{C}_{\text{SS}}$ corresponds to the neck in the type II fluctuation, the radius at which the areal radius takes the minimum value considered relevant for the PBH formation.
-We will give a general proof for the existence of two peaks and the minimum between them for the compaction function in the type II fluctuation in Appendix C.
+The Shibata-Sasaki compaction function was first introduced in Ref.
+[PhysRevD.60.084002] as an estimator for the criterion of PBH formation.
+See Refs.
+[Harada:2023ffo, Harada:2024trx] and references therein for the initial confusion in its introduction, its usefulness, its relation to legitimate compaction functions in different gauges, and its geometrical origin as a conformal compactness function.
+The outer peak of $\mathcal{C}_{\text{SS}}$ corresponds to the neck in the type II fluctuation, the radius at which the areal radius takes the minimum value considered relevant for PBH formation.
+We will give a general proof for two peaks and the minimum between them for the compaction function in the type II fluctuation in the Appendix C.
 We focus on the two peak configurations associated with type II fluctuations (see also Ref.
-[37] for two peak configurations with fluctuations of type I ).
+[Escriva:2023qnq] for two peak configurations with fluctuations of type I ).
 The form of the compaction function for the specific function of $\zeta$ (2.5) is shown in Fig.
 2.
-By using the position of the peak of $\mathcal{C}_{\text{SS}}(r)$, $r_{\text{m}}=\sqrt{2}/k$, we analytically obtain the threshold value $\mu_{\rm c,II}=e/2=1.359\cdots$ for fluctuations of type II.444 Here, we ignore the window function $W(r)$ to estimate the location $r_{\text{m}}=\sqrt{2}/k$ of the peak of compaction function $\mathcal{C}_{\rm SS}(r)$ since this acts on $\zeta(r)$ only near the outer boundary, $r=r_{\rm w}=0.8L$ For $\mu>\mu_{\rm c,II}$ the compaction function takes the maximum value of $0.5$ at least one location with the radial coordinate, and the number of peak positions becomes two, $r_{m1}$ and $r_{m2}$ with $r_{m1}<r_{m2}$.
-This implies that the type II fluctuations always reach the threshold for collapse and form PBHs irrespectively on the initial profile of $\zeta(r)$.
-The isotropic coordinate $r$ does not always give a geometrical size of the enclosed region and may often lead to misunderstanding.
+By using the position of the peak of $\mathcal{C}_{\text{SS}}(r)$, $r_{\text{m}}=\sqrt{2}/k$, we analytically obtain the threshold value $\mu_{\rm c,II}=e/2=1.359\cdots$ for fluctuations of type II.
+[^4] Here, we ignore the window function $W(r)$ to estimate the location $r_{\text{m}}=\sqrt{2}/k$ of the peak of compaction function $\mathcal{C}_{\rm SS}(r)$ since this acts on $\zeta(r)$ only near the outer boundary, $r=r_{\rm w}=0.8L$ .
+For $\mu>\mu_{\rm c,II}$ the compaction function takes the maximum value of $0.5$ at least one location with the radial coordinate, and the number of peak positions becomes two, $r_{m1}$ and $r_{m2}$ with $r_{m1}<r_{m2}$.
+This implies that type II fluctuations always reach the collapse threshold and form PBHs irrespectively of the initial $\zeta(r)$ profile.
+The isotropic coordinate $r$ does not always give the geometrical size of the enclosed region, which can often lead to misunderstanding.
 Therefore, we also explicitly show the compaction function $\mathcal{C}_{\rm SS}$ as a function of the areal radius $R$ in the right panel of Fig.
-2, which will be contrasted with the right panel of Fig.
-13.
-Figure 2: Compaction function $\mathcal{C}_{\text{SS}}$.
-The horizontal axes are the scale-up coordinate radius $z/L$ (left) and the areal radius $R$ (right).
-By taking the amplitude $\mu$ larger, the number of peaks becomes two for $\mu\simeq 1.4$ and above, corresponding to type II fluctuations.
-To follow the time evolution of gravitational collapse from the initial profile of fluctuation, we use the numerical code COSMOS-S [38, 39, 40] based on a 3+1 dimensional code of fourth order Runge–Kutta method with BSSN formalism [32, 33].
-This code is specialized for spherically symmetric systems using the CARTOON method [34].
+2.
+Figure 2: In the left panel, the vertical and horizontal axes are the compaction function $\mathcal{C}_{\text{SS}}(r)$ and the scale-up coordinate $z/L$, respectively, for the initial data sets of the long-wavelength solutions generated by Eq.
+(2.5) for the same set of the amplitude parameter $\mu$.
+We can see that $\mathcal{C}_{\text{SS}}(r)$ takes only one maximum, which is smaller than $1/2$, for $\mu=0.5$ and $1.2$, while it takes two distinct maxima, which are equal to $1/2$, and a single minimum between the two maxima for $\tilde{\mu}=1.4,1.6$ and $1.8$.
+The latter cases correspond to type II fluctuations as shown in Appendix C.
+In the right panel, the vertical and horizontal axes are the compaction function $\mathcal{C}_{\text{SS}}(r)$ and the areal radius $R(r)$, respectively, for the same initial data sets.
+We can see that $\mathcal{C}_{\text{SS}}(r)$ takes only the maximum or the two maxima at the larger areal radius $R$ for the larger values of $\mu$.
+This implies that the mass of the formed black hole increases as $\mu$ increases, as seen in Fig.
+12.
+To follow the time evolution of gravitational collapse from the initial profile of fluctuation, we use the numerical code COSMOS-S [PhysRevLett.111.161102, PhysRevD.89.104032, PhysRevD.105.103538] based on a 3+1 dimensional code of fourth order Runge–Kutta method with BSSN formalism [shibata1995evolution, PhysRevD.59.024007].
+This code is specialized for spherically symmetric systems using the CARTOON method [doi:10.1142/S0218271801000834].
 We consider the time evolution of the initial fluctuation given by the functional form of $\zeta$ (2.5) with the equation of state parameter $w=1/3$, namely, radiation fluid.
 Here, the asymptotically spatially flat FLRW boundary condition is implemented as in Ref.
-[11].
-However, we could not avoid the constraint violation at the outer boundary, and this propagates into the bulk region.
-Therefore, in the following analyses, we terminate the time evolution before the violation reaches the central region relevant for the analyses of PBH formation.
+[PhysRevD.60.084002].
+However, we could not avoid violating the constraint at the outer boundary, which propagates into the bulk region.
+Therefore, we terminate the time evolution before the violation reaches the central region relevant to the PBH formation in the simulation.
 We also find the constraint violation near the center for cases where an apparent horizon forms.
-Nevertheless, since we find both independent future-directed null directions are inward there, the constraint violation does not propagate outward.
-In practice, to avoid a crash of numerical calculation, we excise several grids inside the apparent horizon if needed.
-We take the coordinate size $L$ of the numerical region ($0\leq z\leq L$ or, equivalently, $0\leq r\leq L$) as the unit length scale and set the initial value of the scale factor to unity.
-We fix the ratio between the background initial Hubble length $H_{\rm b}^{-1}$ and the typical scale $k^{-1}$ of the initial fluctuation as $k^{-1}=10H_{\text{b}}^{-1}$ which ensure the long-wavelength approximation.
-For a longer duration of the time evolution, we need to take a larger region than the initial Hubble length to avoid the propagation of the constraint violation.
+Nevertheless, since we find both independent future-directed null directions inward, the constraint violation does not propagate outward.
+In practice, we excise several grids inside the apparent horizon to avoid the numerical calculation from crashing if needed.
+We take the coordinate size $L$ of the numerical region ($0\leq z\leq L$ or equivalently $0\leq r\leq L$) as the unit length scale and set the initial value of the scale factor to unity.
+We fix the ratio between the initial Hubble background length $H_{\rm b}^{-1}$ and the typical scale $k^{-1}$ of the initial fluctuation as $k^{-1}=10H_{\text{b}}^{-1}$ that ensures the long-wavelength approximation.
+To avoid propagating the constraint violation, we need to take a larger region than the initial Hubble length for a longer duration of the time evolution.
 In the following, we take $H_{\text{b}}^{-1}=0.01L$ or $H_{\text{b}}^{-1}=0.005L$ depending on the duration of the required time evolution.
 The horizon entry time $t_{H}$ is defined as $k^{-1}=(aH_{\text{b}})^{-1}|_{t=t_{H}}$ throughout this paper.
-3 Spacetime structure and a new definition of type I/II PBHs
+3 Spacetime structure and definition of type A/B PBHs
 3.1 Time evolution of type I/II fluctuations
@@
 3 for $\mu=0.5$, $1.2$, and $1.8$.
-For the cases of type I fluctuations $\mu=0.5$ (left) and $\mu=1.2$ (middle), the former disperses, but the latter contracts and forms a peak at the center in the density plot in Fig.
+For the cases of type I fluctuations with $\mu=0.5$ (left) and $\mu=1.2$ (middle), the former disperses, but the latter contracts and forms a peak at the center in the density plot in Fig.
 3.
 The latter case (middle) indicates a typical gravitational collapse, for which the lapse is decreasing to a very small value as plotted in Fig.
-3 as Ref.
-[40].
-For a type II fluctuation $\mu=1.8$ case (right), the collapsing behavior happens similarly to the $\mu=1.2$ (middle) case in the plots of density and lapse.
+3 as in Ref.
+[PhysRevD.105.103538].
+For a type II fluctuation case with $\mu=1.8$ (right), the collapsing behavior occurs similarly to the $\mu=1.2$ (middle) case in the density and lapse plots.
 (a) $\mu=0.5$
@@
 (c) $\mu=1.8$
-Figure 3: Snapshots of the fluid energy density contrast $\rho/\rho_{\text{bg}}$ and lapse function $\alpha$ of type I fluctuations (left and middle) and a type II fluctuation (right).
-The initial configurations ($t=0.01t_{H}$) are shown in black on a different scale.
-To check the convergence of the simulation, we plot the spatial average value of the Hamiltonian constraint555 First, we evaluate the appropriately normalized value of the constraint violation $\mathcal{H}_{p}$ at each grid point labeled by $p$.
-Then the average is taken as $\sum_{p=1}^{N}\mathcal{H}_{p}/N$, where $N$ is the total number of the grid points.
-[21](Fig.
+Figure 3: Snapshots of the fluid energy density contrast $\rho/\rho_{\text{bg}}$ and lapse function $\alpha$ for $\zeta(r)$ given by Eq.
+(2.5) for $\mu=0.5$, $1.2$ and $1.8$ in the left, middle and right panels, respectively.
+The initial configurations ($t=0.01t_{H}$) are shown at the top of each panel in black on a different scale.
+We can see that the evolution does not lead to black hole formation for $\mu=0.5$ but does for $\mu=1.2$ and $1.8$.
+To check the convergence of the simulation, we plot the spatial average value of the Hamiltonian constraint [^5] First, we evaluate the appropriately normalized value of the constraint violation $\mathcal{H}_{p}$ at each grid point labeled by $p$.
+Then the average is taken as $\sum_{p=1}^{N}\mathcal{H}_{p}/N$, where $N$ is the total number of the grid.
+[Harada:2015yda] (Fig.
 4).
@@
 (c) $\mu=1.8$
-Figure 4: The average value of the Hamiltonian constraint violation with different resolutions of the number of grid points for three typical cases for no collapse (left), collapse from a type I fluctuation (middle), and collapse from a type II fluctuation (right), respectively.
+Figure 4: The average value of the Hamiltonian constraint violation with different resolutions of the number of grid points for $\zeta(r)$ given by Eq.
+(2.5) for $\mu=0.5$, $1.2$ and $1.8$ in the left, middle and right panels, respectively.
 The horizontal axis is a time coordinate normalized by the time of the horizon entry.
-We use the simulation result to $t=50t_{H}$ to evaluate spacetime structures around horizon formation before the violation propagation reaches the bulk region from the outer boundary.
-For evaluating a longer time evolution of the PBH mass, the average is taken outside the BH apparent horizon (or the future-outer trapping horizon, described in Sec.
+We use the simulation result upto $t=50t_{H}$ to deduce spacetime structures near the horizon formation before the violation propagation reaches the bulk region from the outer boundary.
+For computing a longer time evolution of the PBH mass, the average is taken outside the BH apparent horizon (or the future-outer trapping horizon, described in Sec.
 3.2) after the horizon formation at around $t=40t_{H}$ for $\mu=1.2$ and $t=24t_{H}$ for $\mu=1.8$.
 3.2 Identification of BH apparent horizon
-First, let us introduce two independent future-directed radial null vector fields, $l$ and $l^{\prime}$.
+First, we introduce two independent future-directed radial null vector fields, $l$ and $l^{\prime}$.
 We also define the null expansion for $l$ and $l^{\prime}$ as $\theta$ and $\theta^{\prime}$.
 A null expansion is expressed as $\theta\coloneqq 2\mathcal{L}_{l}R/R$, where $\mathcal{L}_{l}$ denotes the Lie derivative along $l$.
-Trapping horizons are defined by the surfaces that satisfy $\theta\theta^{\prime}=0$ or equivalently the compactness $12M/R=$ 1 with $M$ being the Misner–Sharp mass defined by $M(t,r)\coloneqq R(1-\nabla_{\mu}R\,\nabla^{\mu}R)/2$ [29, 41, 42, 8].
+A trapping horizon is defined by a hypersurface foliated by marginal surfaces that satisfy $\theta\theta^{\prime}=0$ or equivalently the compactness $2M/R=1$ with $M$ being the Misner–Sharp mass defined by
+$M(t,r)\coloneqq\frac{1}{2}R(1-\nabla_{\mu}R\,\nabla^{\mu}R).$ (3.1)
+A trapping horizon is said to be future, past and bifurcating for $\theta+\theta^{\prime}<0$, $\theta+\theta^{\prime}>0$ and $\theta+\theta^{\prime}=0$, respectively [PhysRev.136.B571, Hayward:1994bu, Hayward:1993wb, galaxies10060112, Maeda:2009tk].
+A trapping horizon can be identified with an apparent horizon in the present setting.
+Black hole and cosmological apparent horizon are future and past trapping horizon, respectively.
 In Fig.
-5, the snapshots of the compactness $\quantity(2M/R)(z)$ are plotted.
-For $\mu=0.5$, the intersection between $\quantity(2M/R)(r)$ and the horizontal line of $12M/R=$ 1 is always only one.
-On the other hand, for the cases $\mu=1.2$ and $1.8$, the line intersects the $12M/R=$ 1 more than three times at a sufficiently late time.
-The worldlines of the intersections are shown in Fig.
+5, the snapshots of the compactness $\quantity(2M/R)(z)$ are plotted for $\mu=0.5,1.2$ and $1.8$.
+For all these three cases, we can see that at the earliest time $t=0.01t_{H}$, the compactness monotonically increases from 0 at $z=0$ and crosses $1$ at some value of $z$, corresponding to a past trapping horizon or a cosmological horizon.
+For $\mu=0.5$, the compaction monotonically increases even in the later times and crosses $1$ only once even in the late times.
+This implies that no black hole forms in this case.
+For $\mu=1.2$, after some evolution, the monotonicity breaks down.
+It obtains a maximum and a minimum both smaller than $1$.
+The maximum is first smaller than $1$ and later gets larger than $1$, while the minimum continues to be below $1$.
+As the maximum exceeds $1$, the number of intersection points changes as $1$, 22, and $3$.
+This implies black hole formation inside the cosmological horizon.
+For $\mu=1.8$, at first sight, the behavior is similar to that of $\mu=1.2$, but we can see that the maximum and minimum are larger than $1$ for $t=6t_{H}$, when the monotonicity breaks down.
+As time proceeds further, the minimum decreases and gets smaller than $1$ for $t=12t_{H}$, while the maximum continues to be above $1$.
+Around $t=18t_{H}$, another set of maximum and minimum points appear between the pre-existing two extrema, and the minimum gets smaller than 1 for $t=24t_{H}$.
+The number of intersection points becomes five.
+Then the outer maximum decreases to 1 and increases again after touching the horizontal line of $2M/R=1$.
+As will be understood from Fig.
+6, the future trapping horizon appears after this moment of the turnaround of the maximum.
+After that, a similar behavior is observed for the inner maximum.
+Then the minimum between the two maxima increases to a value larger than 1, and finally, there are 3 intersections with the line of $2M/R=1$.
+The worldlines of the intersection points, which correspond to trapping horizons, are shown in Fig.
 6.
-To identify the future-trapped ($\theta<0$ and $\theta^{\prime}<0$) and past-trapped ($\theta>0$ and $\theta^{\prime}>0$) regions we also show directions of two independent null vector fields $l$ and $l^{\prime}$ together with the contour curves of the areal radius in Figs.
+We can see a single trapping horizon for $\mu=0.5$, two distinct trapping horizons for $\mu=1.2$ and two intersecting trapping horizons for $\mu=1.8$.
+To identify future-trapped ($\theta<0$ and $\theta^{\prime}<0$) and past-trapped ($\theta>0$ and $\theta^{\prime}>0$) regions, we also show directions of two independent null vector fields $l$ and $l^{\prime}$ together with the contour curves of the areal radius in Figs.
 7 and 8.
-Comparing the directions of the null vector fields and contour curves of the areal radius, we can realize that the red and blue shaded regions in Fig.
+Comparing the directions of the null vector fields and the contour curves of the areal radius, we can realize that the red and blue shaded regions in Fig.
 6 are future- and past-trapped regions, respectively.
-More specifically, in the future(past)-trapped region, both arrows direct to directions along which the areal radius decreases (increases), and the curve described by red points indicates the boundary satisfying $\theta\theta^{\prime}=0$.
-Then, we refer to a marginally future trapped surface that is the outer boundary of the future-trapped region on the $t=$const hypersurface as a BH apparent horizon.
-Similarly, we refer to a marginally past-trapped surface that is the inner boundary of the past-trapped region on the $t=$ const hypersurface as a cosmological apparent horizon.
-It is not easy to understand the horizon configuration.
+More specifically, in the future (past)-trapped region, the areal radius decreases (increases) along both of the arrows, and the red curve indicates the boundary satisfying $\theta\theta^{\prime}=0$.
+Then, the boundary of the future (past)-trapped region is a future (past) trapping horizon.
+We can see in Fig.
+6 that there is only a single past trapping horizon for $\mu=0.5$ and separate future trapping horizon and past trapping horizon for $\mu=1.2$.
+For $\mu=1.8$, two trapping horizons intersect each other twice.
+The intersection points correspond to bifurcating trapping horizons.
+The enlarged views of Fig.
+8 around the intersection points and the region surrounded by the red curves are provided in Fig.
+9.
+It is not easy to understand the horizon configuration, in particular, for the case of $\mu=1.8$.
 Therefore, we carefully investigate it and present a more concrete classification of trapping horizons in the following sections.
-Readers who would like to get a quick intuition about the horizon configuration may refer to the schematic conformal diagram in Figures 16 and 17.
+Readers who would like to get a quick intuition about the horizon configuration can refer to the schematic conformal diagram in Figs.
+17 and 18.
 (a) $\mu=0.5$
@@
 (c) $\mu=1.8$
-Figure 5: Snapshots of the compactness $2M(t,r)/R(t,r)$.
-For the $\mu=1.2$ case, the trapping horizon at the center is formed at around $t=40t_{H}$ (middle).
-For the $\mu=1.8$ case, the horizon is formed at around $t=24t_{H}$ (right).
+Figure 5: Snapshots of the compactness $2M(t,r)/R(t,r)$ for $\zeta(r)$ given by Eq.
+(2.5) for $\mu=0.5$, $1.2$ and $1.8$ in the left, middle and right panels, respectively.
+For all these three cases, we can see that at the earliest time $t=0.01t_{H}$, the compactness monotonically increases from 0 at $z=0$ and crosses $1$ at some value of $z$, corresponding to a past trapping horizon or a cosmological horizon.
+For $\mu=0.5$, the compaction monotonically increases even in the later times and crosses $1$ only once even in the late times.
+This implies that no black hole forms in this case.
+For $\mu=1.2$, after some evolution, the monotonicity breaks down.
+It obtains a maximum and a minimum both smaller than $1$.
+The maximum is first smaller than $1$, becomes equal to 11 at around $t=40t_{H}$, and later gets larger than $1$, while the minimum continues to be below $1$.
+As the maximum exceeds $1$, the number of intersection points changes as $1$, 22, and $3$.
+This implies the formation of a black hole trapping horizon inside the cosmological horizon at around $t=40t_{H}$.
+For $\mu=1.8$, at first sight, the behavior is similar to that of $\mu=1.2$, but we can see that the maximum and minimum are larger than $1$ for $t=6t_{H}$ when the monotonicity breaks down.
+As time proceeds further, the minimum decreases and gets smaller than $1$ for $t=12t_{H}$, while the maximum continues to be above $1$.
+Around $t=18t_{H}$, another set of maximum and minimum points appear between the pre-existing extrema, and the minimum gets smaller than 1 for $t=24t_{H}$.
+As the zoomed subplot shows, the number of intersection points becomes five for $t=24t_{H}$.
+Then, the outer maximum decreases to 1 and increases again after touching the horizontal line of $2M/R=1$.
+As will be understood from Fig.
+6, the future trapping horizon appears after this moment of the turnaround of the maximum.
+After that, a similar behavior is observed for the inner maximum.
+Then, the minimum between the two maxima increases to a value larger than 1, and finally, there are 3 intersections with the line of $2M/R=1$.
 (a) $\mu=0.5$
@@
 (c) $\mu=1.8$
-Figure 6: Configurations of the trapping horizons $R(t,r)=2M(t,r)$. It corresponds to the points $(t,r)$ of the compactness $12M/R=$ 1 in Fig. 5.
-Figure 7: Directions of two independent null vector fields $l$ and $l^{\prime}$ are depicted as orange and blue arrows for $\mu=1.2$ case (type I) in the middle panel of Fig.
-6.
-$R$-contour with 1.0 increments starting from $R=1.0$ and increasing from left bottom to right top is plotted as black curves.
-Figure 8: Directions of two independent null vector fields $l$ and $l^{\prime}$ are depicted as orange and blue arrows for $\mu=1.8$ case (type II) in the right panel of Fig.
-6.
+Figure 6: Configurations of the trapping horizons $R(t,r)=2M(t,r)$ for $\zeta(r)$ given by Eq.
+(2.5) for $\mu=0.5$, $1.2$ and $1.8$ in the left, middle and right panels, respectively.
+It corresponds to the points $(t,r)$ of the compactness $2M/R=1$ in Fig.
+5.
+We can see that there is only a single past trapping horizon for $\mu=0.5$ and separate future trapping horizon and past trapping horizon for $\mu=1.2$.
+For $\mu=1.8$, two trapping horizons intersect each other twice.
+The intersection points correspond to bifurcating trapping horizons.
+The region surrounded by the trapping horizons is untrapped with the null expansions $\theta>0$ and $\theta^{\prime}<0$.
+The future and past boundaries of this region are future and past trapping horizons with the ends being bifurcating trapping horizons, respectively.
+Figure 7: Directions of two independent null vector fields $l$ and $l^{\prime}$ are depicted with blue and orange arrows in the top and bottom panels, respectively, for $\zeta(r)$ given by Eq.
+(2.5) for $\mu=1.2$.
 The $R$ contour is plotted as black curves with 1.0 increments starting from $R=1.0$ and increasing from left bottom to right top.
-In addition to the original plot of $R$-levels with 1.0 increments, it is also plotted with 0.1 increments from $R=8.1$ to $8.9$.
-3.3 Spacetime structures of type I/II PBH
+The red curves indicate trapping horizons satisfying $\theta\theta^{\prime}=0$.
+Comparing the directions of the null vector fields and the contour curves of the areal radius, we can see that the red and blue shaded regions in Fig.
+6 are future- and past-trapped regions, respectively.
+More specifically, in the future (past)-trapped region, the areal radius decreases (increases) along both of the arrows.
+Figure 8: Same as Fig.
+7 but for $\mu=1.8$.
+The $R$ contour is plotted as black curves with 1.0 increments starting from $R=1.0$ and increasing from left bottom to right top.
+In addition to the original plot of $R$ levels with increments of 1.0, it is also plotted with increments of 0.1 from $R=8.1$ to $8.9$.
+We can see that two trapping horizons intersect each other twice.
+The intersection points correspond to bifurcating trapping horizons.
+Looking at these panels very carefully, we can see that the intersection points on the left-hand and right-hand sides are a maximum and a saddle, respectively, in terms of the areal radius.
+The region surrounded by the trapping horizons is untrapped with the null expansions $\theta>0$ and $\theta^{\prime}<0$.
+The future and past boundaries of this region are future and past trapping horizons with the ends being bifurcating trapping horizons, respectively.
+The enlarged views around the intersection points by the two red curves and the region surrounded by the red curves are provided in Fig.
+9.
+Figure 9: Same as Fig.
+8 but the enlarged views of the complicated region near the intersection points.
+Note that the red curves divide the whole region into five, i.e., the bottom left (I), top middle (II), top right (III), bottom right (IV), and the region in the center of the panel surrounded by the red curves (V).
+On each panel, $R$ is small on the bottom left, top middle, and bottom right, while large on the top right.
+So, we can conclude that the left-hand and right-hand intersection points of the red curves correspond to a maximum and a saddle, respectively, in terms of $R$.
+We can see that the blue arrows on the top panel are tangent to the $R$-contour on the red curve convex upward, implying that $\theta=0$ there, while the orange arrows on the bottom panel are tangent to the $R$-contour on the red curve convex downward, implying that $\theta^{\prime}=0$.
+Thus, the signs of the null expansions $(\theta,\theta^{\prime})$ in regions I, II, III, and IV are given by (,+)(-,+), (,)(-,-), (,+)(-,+) and $(+,+)$, respectively.
+We can see that the signs are (+,)(+,-) in region V, implying that this region is untrapped.
+The part of the red curve convex upward that is the boundary of region V has the signs (0,)(0,-), implying that it is a future trapping horizon.
+The same analysis implies that the part of the red curve convex downward that is also the boundary of region V is a past trapping horizon.
+3.3 Spacetime structures of type A/B PBH
 Let us introduce the classification of trapping horizons following Refs.
-[42, 43].
+[Hayward:1993wb, Maeda:2009tk, Harada_2018].
 For this purpose, to avoid confusion, we introduce another notation for the two radial null vector fields and associated expansions as $l_{\pm}$ and $\theta_{\pm}$ for a specific trapping horizon.
 Then we always set $\theta_{+}=0$ when considering a trapping horizon satisfying $\theta_{+}\theta_{-}=0$.
-The classification of trapping horizons can be summarized as Table 1.
+The classification of the trapping horizons can be summarized in Table 1.
 $\theta_{+}=0$ + 0 -
@@
 $\mathcal{L}_{-}\theta_{+}$ Inner Degenerate Outer
-Table 1: Classification of trapping horizons satisfying $\theta_{+}\theta_{-}=0$ [42].
-A black hole, white hole, and cosmological trapping horizons with $w=1/3$ $(w=0)$ are identified with a future-outer, past-outer, and past-degenerate (past-inner) trapping horizons, respectively.
+Table 1: Classification of trapping horizons satisfying $\theta_{+}\theta_{-}=0$ [Hayward:1993wb].
+A black hole, a white hole, and cosmological trapping horizons with $w=1/3$ $(w=0)$ are identified with future-outer, past-outer, and past-degenerate (past-inner) trapping horizons, respectively.
 (a) Type A PBH
 (b) Type B PBH
-Figure 9: The conformal diagrams of type A and B PBHs in radiation domination inferred from the present numerical results.
-The left figure defines a type A PBH as a causal structure of the typical case of PBH.
+Figure 10: The conformal diagrams inferred from the present numerical results in radiation domination.
+The left panel defines a type A PBH as a causal structure of the typical case of PBH.
 The right one includes bifurcating trapping horizons and defines a type B PBH.
@@
 The bifurcating trapping horizons ($\theta_{+}=\theta_{-}=0$) are depicted as circles in the type B case.
-Unlike the dust case discussed in Appendix A, we can only obtain numerical solutions such as those shown in Figs.
-7 and 8 for the radiation case.
-From these figures and carefully comparing the similarities and differences from the exact conformal diagrams of the dust solution discussed in Appendix A, we can deduce possible conformal diagrams and proper identifications of the trapping horizons for the radiation case as shown in the left and right panels in Fig.
-9 as possible conformal diagrams of type I and type II PBHs, respectively.
+Unlike the dust case discussed in Appendix A, we can only obtain numerical solutions for radiation case, such as those shown in Figs.
+7 and 8 for $\mu=1.2$ and $1.8$, respectively.
+The trajectories of $\theta=0$ and $\theta^{\prime}=0$ are independently given by two smooth curves indicated by red lines in the top and bottom panels, respectively.
+In the top panels, on the smooth red line in the bottom right, the blue arrows are tangent to the solid black lines indicating the contour of the areal radius, which implies $\theta=0$ there.
+In the bottom panels, on the other hand, the orange arrows are tangent to the solid black lines on the other smooth red line that is convex downward in the top left, which implies $\theta^{\prime}=0$ there.
+A very careful look at Figs.
+8 and 9 reveals that the intersection points on the left-hand and right-hand sides are a maximum and a saddle, respectively, in terms of the areal radius.
+From these figures, we can find that the region surrounded by the trapping horizons, which we call region V, is untrapped with the null expansions $\theta>0$ and $\theta^{\prime}<0$.
+The part of the red curve convex upward that is the boundary of region V is a future trapping horizon, while the part of the red curve convex downward that is the boundary of region V is a past trapping horizon.
+From these figures and carefully comparing the similarities and differences from the exact conformal diagrams of the dust solution discussed in the Appendix A, we can deduce possible conformal diagrams and proper identifications of the trapping horizons for the radiation case as shown in the left and right panels in Fig.
+10 as possible conformal diagrams of type A and type B PBHs, respectively.
 The crucial difference between the right and left panels in Fig.
@@
 8.
-In this paper, we propose to use this difference to classify PBH formation processes into types I and II; that is, we identify the PBH as of type I if there is no bifurcating trapping horizon and of type II if there appears at least one bifurcating trapping horizon during the time evolution associated with the intersection point.
-This difference has already been pointed out by KHW in Ref.
-[1] for the case of the dust fluid.
+In this paper, we propose to use this difference to classify the PBH formation processes into types A and B; that is, we identify the PBH as type A if there is no bifurcating trapping horizon and as type B if at least one bifurcating trapping horizon appears during the time evolution associated with the intersection point.
+KHW has already pointed out this difference in Ref.
+[PhysRevD.83.124025] for the case of the dust fluid.
 Unlike the current proposal, the authors in Ref.
-[1] did not clearly distinguish the definitions of type I/II PBHs and type I/II fluctuations.
-However, in our simulation with $w=1/3$, we find that the horizon configuration is of type I for the $\mu=1.6$ case, but the initial fluctuation is of type II.
-Therefore, there is a gap between the definitions of type I/II PBHs and type I/II fluctuations.
+[PhysRevD.83.124025] did not distinguish between the definitions of type I/II and type A/B.
+However, in our simulation with $w=1/3$, we find the case of type II-A, the horizon configuration is of type A from type II fluctuation for the $\mu=1.4$ case.
 This fact is our motivation for the new definition of the PBH.
 The critical difference between the dust and the radiation cases is that trapping horizons in the FLRW regions are timelike and null for the former and the latter, respectively.
-See Figs.
-15, 16 and 18 and Ref.
-[43].
+See, e.g., Ref.
+[Harada_2018].
 Then, we confirm that the outermost boundary surface of the future-trapped region is the future-outer trapping horizon, which describes a BH apparent horizon.
-In addition, the outermost boundary surface between the past-trapped and non-trapped regions describes a cosmological apparent horizon, which asymptotically approaches the past degenerate trapping horizon in the background FLRW universe.
+Therefore, the outermost boundary surface between the past-trapped and untrapped regions, which describes a cosmological apparent horizon, asymptotically approaches the past degenerate trapping horizon in the background FLRW universe.
 It is worth noting that the schematic diagram given by Fig.
-9 implies that all light rays emitted from the center reach not the future null infinity but the future singularity in the type II PBH.
+10 implies that any light ray emitted from the center reaches not the future null infinity but the future singularity of the type B PBH.
 Although we cannot make any general statement for the moment, the null trajectories indicated by the orange arrows in Fig.
 8 seem consistent with this observation.
-Precisely, one could also consider the critical case between type I and type II PBHs when the two bifurcating trapping horizons degenerate, then we additionally classify this case as a type IId PBH, which has just one bifurcating trapping horizon.
 4 PBH mass
 4.1 PBH mass with the Gaussian initial fluctuation
+Let us introduce the mass of the PBH in two different ways at each characteristic time.
+The first is to calculate the Misner-Sharp mass at the apparent horizon at its first appearance with the future-trapped region (see Fig.
+6).
+We refer to the mass at this time as the initial BH mass.
+In a type A PBH, the BH apparent horizon appears as a future-outer trapping horizon, while it appears as a bifurcating outer trapping horizon in type B.
+We also calculate the final PBH mass by estimating the effect of the accretion.
+The accretion process is well described by the mass accretion model, often called Novikov-Zeldovich or Bondi accretion [Zeldovich:1967lct, 1952MNRAS.112..195B].
+In this accretion model, the time evolution of the mass is given by
 […]
-[…]
-where $\rho_{b}=\rho_{b,0}(t_{0}/t)^{2}$ is the background energy density, and $F$ is a constant that corresponds to the efficiency of accretion, which is commonly numerically found to be of order $\mathcal{O}(1)$.
+where $\rho_{b}=\rho_{b,0}(t_{0}/t)^{2}$ is the background energy density, and $F$ is a nondimensional constant that corresponds to the efficiency of accretion, which is commonly numerically found to be of order $\mathcal{O}(1)$ and will depend on the details of the accretion process.
 It is important to note that Eq.
-(4.1) is not valid at the time of formation of the trapping horizon since it neglects the cosmological expansion as Carr and Hawking (1974) [4] already pointed out.
-However, it will correctly describe the accretion into the PBHs for sufficiently large times once we have a quasi-stationary flow into the PBH.
+(4.1) is not valid at the time of formation of the trapping horizon since it neglects the cosmological expansion as Carr and Hawking (1974) [10.1093/mnras/168.2.399] already pointed out.
+However, once we have a quasi-stationary flow into the PBH, it will correctly describe the accretion into the PBHs for sufficiently long times.
 The integrated solution is,
@@
 where $M_{f}$ is the final BH mass.
-We fix the values of $M_{f}$ and $F$ by fitting Eq.
+We find the values of $M_{f}$ and $F$ by fitting Eq.
 (4.2) to the numerical data in a sufficiently late finite time interval denoted in Fig.
@@
 This behavior is consistent with the results in Ref.
-[45].
-Figure 10: The time evolution of the BH mass for each value of $\mu$.
+[escriva_2021].
+Figure 11: The time evolution of the BH mass for $\zeta(r)$ given by Eq.
+(2.5) for $\mu=1.2,1.4,1.6,1.8,2.0$ and $2.2$.
 The black dashed curves are fitting curves with Eq.
@@
 The fitting region is taken for the interval between $t=450t_{H}$ and $t=500t_{H}$.
-$\mu$ $M_{f}[M_{\text{H}}]$ $F$
+The values of fitting parameters are shown in Table 2.
+$\mu$ $M_{f}/M_{\text{H}}$ $F$
 1.2 11.9 5.0
@@
 2.2 21.4 4.4
-Table 2: The values of $F$ and $M_{f}$ for each amplitude of $\mu$ result from the fitting in Fig. 10.
-Let us check the $\mu$ dependence of the PBH mass. In Fig. 11, we plot the initial and final PBH mass values for each value of $\mu$.
-Figure 11: PBH mass in terms of the horizon mass at horizon entry $M_{H}$ to the amplitude of fluctuation $\mu$.
+Table 2: The values of $F$ and $M_{f}$ for $\zeta(r)$ given by Eq. (2.5) for $\mu=1.2,1.4,1.6,1.8,2.0$ and $2.2$ obtained from the fitting in Fig. 11.
+Let us check the $\mu$ dependence of the PBH mass.
+In Fig.
+12, we plot the initial and final PBH mass values for each value of $\mu$.
+This behavior can be seen for an amplitude much larger than the critical regime, where $M$ obeys a power law against $(\mu-\mu_{c})$ [PhysRevLett.70.9, Evans:1994pj, Koike:1995jm, Niemeyer:1999ak, Musco:2012au].
+Figure 12: PBH mass in terms of the horizon mass at horizon entry $M_{H}$ to the amplitude of fluctuation $\mu$ for $\zeta(r)$ given by Eq.
+(2.5).
 In this case, in terms of fluctuation types, $\mu=1.2$ is classified as type I, while $\mu=1.4\>\text{to}\>2.2$ as type II, as seen in Fig.
 2.
-In terms of formation types, on the other hand, $\mu=1.2$ to $1.6$ are classified as type I, while $\mu=1.8$ to $2.2$ as type II, as seen in Figs.
-6 and 9.
-The initial and final PBH masses get larger with increasing amplitude $\mu$. We will check the profile dependence in the next subsection.
+In terms of formation types, on the other hand, $\mu=1.2$ to $1.6$ are classified as type A, while $\mu=1.8$ to $2.2$ as type B, as seen in Figs.
+6 and 10.
+We can see that the initial and final PBH masses get larger with increasing amplitude $\mu$.
+The initial and final PBH masses get larger with increasing amplitude $\mu$. We will check whether this behavior appears for a different profile or not in the next subsection.
 4.2 PBH mass with the initial fluctuation motivated by the three-zone model
-In order to construct a specific initial profile that can be contrasted with the former Gaussian profile, let us try to utilize the three-zone profile as a reference, which is an extension of the model used in Ref.
-[1].
-To realize a similar situation to the schematic conformal diagram, let us consider the closed FLRW region for the initial overdense region.
+[…]
+To construct a specific initial profile that can be compared with the former Gaussian profile as Eq.
+(2.5) and Fig.
+1, let us try to utilize the three-zone profile as a reference, an extension of the model used in Ref.
+[PhysRevD.83.124025].
+Let us consider the closed FLRW region for the initial overdense region to realize a situation similar to the schematic conformal diagram.
 Then, as is shown in the Appendix B, under the boundary condition $\zeta=0$ at the surface of the overdense region, we obtain the following form of $\zeta(r)$:
@@
 where $k^{-1}$ is the comoving radius of $S^{3}$ associated with the closed FLRW and $0<\chi_{a}\leq\pi/k$ describes the radial coordinate of the edge of the closed FLRW region.
-Although this expression is relevant for the region $r\leq r_{a}$, we simply extend this expression in the region $r>r_{a}$, and to realize the condition $\zeta(L)=0$, we use the following specific functional form of $\zeta(r)$:
+Although this expression is relevant for the region $r\leq r_{a}$, this functional form cannot be directly used as the initial profile for our numerical simulation.
+This is because, in our simulation, we suppose that the initial profile is continuously connected to the flat FLRW background, namely, $\zeta(r)$ should vanish for $r>L$.
+To realize this condition, first, we simply extend the expression (4.3) to the region $r>r_{a}$, and use the following specific functional form of $\zeta(r)$ for the initial condition of our simulation:
 $\zeta(r)$ $=2\tilde{\mu}\ln\quantity(\Psi_{3}(r)-\Psi_{3}(L))W(r),$ (4.4)
@@
 The value of $l$ is fixed as $k^{-1}=10H_{\rm b}^{-1}$ as before, and the value of $\chi_{a}$ is set as $\pi\tilde{\mu}/k$ for a given value of $\tilde{\mu}$.
+This profile looks complicated and differs from the original three-zone model profile but is useful for showing a specific example of the non-monotonic behavior of the PBH mass.
 The initial profiles of $\zeta$, $R$ and the compaction function $\mathcal{C}_{\rm SS}$ are shown in Figs.
-12 and 13.
-The tail profile is different from the original three-zone model, but it is similar to the three-zone profile around the peak of $\zeta$.
-This profile also exhibits the type II feature with the neck structure for a large amplitude $(\tilde{\mu}\gtrsim 0.85)$.
-Figure 12: Initial three-zone profile of $\zeta(r)$ (left) and areal radius $R(r)$ (right) with respect to the amplitude parameter $\tilde{\mu}$.
-Taking the amplitude parameter $\tilde{\mu}$ larger, a type II fluctuation appeared for $\tilde{\mu}$ larger than $0.80$ with the neck $\partial_{r}R=0$.
-The original profile near the peak $\zeta_{\rm cFLRW}$ [1] is plotted in Fig.
-19 in Appendix B.
-Figure 13: Compaction function $\mathcal{C}_{\text{SS}}(r)$ for thnon-linearprofile.
-The horizontal axis is the coordinate radius with scale-up $z/L$ (left)/ the areal radius $R$ (right).
-By taking the amplitude $\tilde{\mu}$ larger, the number of peaks becomes two for $\tilde{\mu}$ = 0.89, corresponding to type II fluctuation.
+13 and 14.
+The tail profile is different from the original three-zone model, but it is similar to the three-zone profile near the peak of $\zeta$.
+This profile can also exhibit the type II feature with the neck structure for a large amplitude $(\tilde{\mu}\gtrsim 0.85)$.
+Figure 13: Same as Fig.
+1 but for $\zeta(r)$ given by Eq.
+(4.4).
+The functional form (4.4) is inspired by the three-zone model but modified for numerical simulations and is different from the Gaussian-shaped profile (2.5).
+We plot $\zeta(r)$ for the set of the amplitude parameter values $\tilde{\mu}=0.20,0.40,0.60,0.80,0.90$ and $0.92$.
+We can see that $R(r)$ is a monotonically increasing function for $\tilde{\mu}=0.20,0.40,0.60$ and $0.80$, while it increases, decreases and increases again as $r$ increases for $\tilde{\mu}=0.90$ and $0.92$.
+In the latter cases, $R(r)$ has a local maximum and a minimum with $\partial_{r}R=0$, where the minimum corresponds to the neck.
+Figure 14: Same as Fig.
+2 but for $\zeta(r)$ given by Eq.
+(4.4).
+We can see that $\mathcal{C}_{\text{SS}}(r)$ takes the only one maximum, which is smaller than $1/2$, for $\tilde{\mu}=0.20,0.40,0.60$, and $0.80$, while it takes two distinct maxima, which are equal to $1/2$, and a single minimum between the two maxima for $\tilde{\mu}=0.90$ and $0.92$.
+The latter cases correspond to type II fluctuations as shown in Appendix C.
+We can see that $\mathcal{C}_{\text{SS}}(r)$ takes the only maximum or the two maxima at the smaller areal radius $R$ for the larger values of $\tilde{\mu}$.
+This implies that the mass of the formed black hole decreases as $\tilde{\mu}$ increases as seen in Fig.
+15.
+This is not the case for the Gaussian-shaped functional forms for $\zeta(r)$ shown in Fig.
+2, for which the mass of the formed black hole monotonically increases as $\tilde{\mu}$ increases as shown in Fig.
+12 at least for the parameters computed here.
 The behavior of the horizon configuration is similar to the case of the Gaussian profile, and we do not explicitly show it.
-In contrast to the Gaussian profile, the initial PBH mass formed by this profile decreases as the amplitude of fluctuation $\tilde{\mu}$ increases for $\tilde{\mu}>0.65$ as shown in Fig.
-14, even though they are classified as type I fluctuations and formations for $\tilde{\mu}\leq 0.88$.
-This result, as well as the behavior of the Gaussian profile, indicates that whether the initial PBH mass increases or decreases with the amplitude parameter is not only up to its type but also strongly dependent on its profile, together with the ambiguity in quantifying the amplitude of non-linear perturbation.
-We also find the gap between the parameter region of type I/II fluctuations and type I/II PBHs, as indicated in Fig.
+However, in contrast to the Gaussian profile, the initial PBH mass formed by this profile decreases as the amplitude of fluctuation $\tilde{\mu}$ increases for $\tilde{\mu}>0.65$ as shown in Fig.
+15, even though they are classified as type I fluctuations and type A formations for $\tilde{\mu}\leq 0.88$.
+Again, this is well beyond the near-critical regime.
+This result, as well as the behavior of the Gaussian profile, indicates that whether the initial PBH mass increases or decreases as a function of the amplitude parameter is not only up to its type but also strongly dependent on its profile, together with the ambiguity in quantifying the amplitude of non-linear perturbation.
+We also find the gap between the parameter region of type I/II and type A/B PBHs, as indicated in Fig.
+15.
+Figure 15: Initial PBH mass with respect to the amplitude of fluctuation for $\zeta(r)$ given by Eq.
+(4.4) We can see that the initial PBH mass decreases as the amplitude parameter $\tilde{\mu}$ is increased for $\tilde{\mu}\gtrsim 0.65$, after it increases for $\tilde{\mu}_{c}<\tilde{\mu}\lesssim 0.65$.
+The decreasing behavior of the initial PBH mass with the amplitude can be easily predicted by focusing on the profile of the compaction function as a function of the areal radius.
+The initial PBH mass would be roughly estimated by the areal radius at the peak of the compaction function at the horizon entry.
+Therefore, if the areal radius at the peak of the compaction function increases and decreases with increasing the amplitude of the perturbation profile, the initial PBH mass is expected to increase and decrease as well, respectively.
+In the case of the Gaussian profile (2.5), one can find that the areal radius at the peak of the compaction function increases with increasing the value of the amplitude $\mu$ as seen in the right panel of Fig.
+2, while it decreases with increasing $\tilde{\mu}$ in the case of the specific profile (4.4) as seen in the right one of Fig.
 14.
-Figure 14: Initial PBH mass with respect to the amplitude of fluctuation for the three-zone profile.
-The decreasing behavior of the initial PBH mass with the amplitude can be easily predicted by focusing on the profile of the compaction function as a function of the areal radius.
-The initial PBH mass would be roughly estimated by the areal radius of the peak of the compaction function at the horizon entry.
-Therefore, if the areal radius around the peak of the compaction function increases and decreases with the amplitude of the perturbation profile, the initial PBH mass is expected to increase and decrease, respectively.
-In the case of the Gaussian profile, one can find that the areal radius around the peak of the compaction function increases with the value of the amplitude $\mu$ as the right panel of Fig.
-2, while it decreases with $\tilde{\mu}$ in the case of the specific profile (4.4) as the right one of Fig.
-13.
 5 Conclusion
 We have numerically investigated the PBH formation from a type II initial fluctuation $\zeta(r)$, which is defined by the non-monotonicity of the areal radius $R(r)\propto re^{\zeta}$ with respect to the coordinate radius $r$ on the spatial geometry, i.e., the existence of $\partial_{r}R(r)=0$ (called neck or throat structure), in the radiation-dominated situation.
-As in the dust case reported in KHW, the PBH formation process can be classified into types I and II according to the configuration of the trapped regions $R(t,r)\leq 2M(t,r)$ (or trapping horizons $R(t,r)=2M(t,r)$).
-Specifically, if there is a bifurcating trapping horizon, the PBH is classified into type II, and type I otherwise.
-In the dust case, KHW reported that a type I fluctuation results in a type I PBH, and a type II fluctuation does in a type II PBH.
+As in the dust case reported in KHW, the initial fluctuation can be classified into types I and II according to the configuration of the trapped regions $R(t,r)\leq 2M(t,r)$ (or trapping horizons $R(t,r)=2M(t,r)$).
+Specifically, if there is a bifurcating trapping horizon, the PBH is classified as type B, and otherwise, it is classified as type A.
+In the dust case, KHW reported that a type I fluctuation results in a type A PBH and a type II fluctuation in a type B PBH.
 However, in the case of radiation fluids, this is not true.
-We found a gap between the classifications regarding the shape of the initial spatial geometry (fluctuation type) and the structure of trapping horizons (formation type).
-That is, we found that a type II fluctuation may result in a type I PBH.
-Therefore, we propose classifying PBHs based on the existence of the bifurcating trapping horizon independently of the type of the initial fluctuation.
+We found a gap between the classifications regarding the shape of the initial spatial geometry (fluctuation type) and the structure of the trapping horizons (formation type).
+We found that a type II fluctuation may result in a type A PBH, which we will call type II-A PBH.
+Therefore, we propose classifying type A/B PBHs based on the existence of the bifurcating trapping horizon independently of the type of the initial fluctuation.
 In the leading order of the long-wavelength limit, the overdense region can be roughly approximated by the closed FLRW universe, and the amplitude of the density perturbation can be translated into the value of the radial coordinate at the edge of the closed FLRW region.
-Then, the type II fluctuation indicates that the edge is located in the other hemisphere of $S^{3}$ initial hypersurface beyond the great circle.
-Therefore, in the case of type II PBHs, one may expect that the PBH mass would decrease with increasing the amplitude of initial fluctuation since the initial time slice of $S^{3}$ in the overdensity region closes in the limit of large-density perturbation, which indicates the area of the edge is vanishing.
-However, this is not the case for the Gaussian-type initial fluctuation profile, while the other profile exhibits decreasing mass with increasing initial amplitude.
+Then, the type II fluctuation indicates that the edge is located in the other hemisphere of $S^{3}$ beyond the great circle on the initial hypersurface.
+Therefore, in the case of type II PBHs, one may expect that the PBH mass would decrease with increasing the amplitude of initial fluctuation since the edge of the overdense region in $S^{3}$ closes in the limit of large-density perturbation, which indicates that the area of the edge is vanishing.
+However, this is not the case for the Gaussian-type initial fluctuation profile, whereas the other profile exhibits decreasing mass with increasing initial amplitude.
 Therefore, we conclude that the behavior of the initial PBH mass depends on not only the type of fluctuation but also the fluctuation profile.
-Thus, systematic investigation of the profile dependence would be needed, e.g., by using the shape parameter defined as the curvature peak of the compaction function $\mathcal{C}_{\text{SS}}$ [19, 45].
-In addition to the profile dependence, the dependence on the equation of state $p=w\rho$ is also an interesting issue, which fills the gap between KHW and this work [20].
+Thus, a more systematic investigation would be needed to reveal the details of the profile dependence, e.g., by using the shape parameter defined as the curvature peak of the compaction function $\mathcal{C}_{\text{SS}}$ [Escriva:2019phb, escriva_2021].
+In addition to profile dependence, dependence on the equation of state $p=w\rho$ is also an interesting issue, filling the gap between KHW and this work [escriva_2021analy].
 Acknowledgments
-The authors are grateful to K.-I.
+The authors are grateful to I.
+Musco, K.-I.
 Nakao, B.J.
@@
 acknowledges support from the JSPS Postdoctoral Fellowships for Research in Japan (Graduate School of Sciences, Nagoya University).
-T.H.
-is supported in part by JSPS KAKENHI Grant Nos.
-19K03876 and 20H05853.
-D.S.
-would like to take this opportunity to thank the “Nagoya University Interdisciplinary Frontier Fellowship” supported by Nagoya University and JST, the establishment of university fellowships towards the creation of science technology innovation, Grant Number JPMJFS2120.
-C.Y.
-is supported in part by JSPS KAKENHI Grant Nos.
-20H05850 and 20H05853.
-C.Y.
-and T.H.
-acknowledge the hospitality at APCTP during the focus research program ”Black Hole and Gravitational Waves: from modified theories of gravity to data analysis”, where part of this work was discussed.
+The authors are supported in part by the JSPS KAKENHI Grant Nos.
+19K03876(TH), 20H05850(CY), 20H05853(TH, CY), and 24KJ1223(DS).
+KU would like to take this opportunity to thank the “THERS Make New Standards Program for the Next Generation Researchers” supported by JST SPRING, Grant Number JPMJSPS2125.
+CY and TH acknowledge the hospitality at APCTP during the focus research program ”Black Hole and Gravitational Waves: from modified theories of gravity to data analysis”, where part of this work was discussed.
 Appendix A Conformal diagrams of the three-zone model for the dust case
-In the three-zone model [7], the spacetime of PBH is described by the patchwork of the three different regions: the innermost overdense region is given by a spatially closed FLRW universe, the intermediate underdense region, and the outermost flat FLRW universe as the background.
-We also note that the intermediate underdense region must be introduced to compensate for the mass excess of the innermost region and connect the spacetime to the background spacetime without a mass gap.
-Without the intermediate underdense region, this model cannot be the exact solution, even for the dust case.
-For the underdense region with a perfect fluid with $w=1/3$, it would not be tractable to obtain any analytical and exact solutions of such dynamical inhomogeneous spacetimes with non-linear sound waves.
+In the three-zone model [PhysRevD.88.084051], the spacetime of PBH is described by the patchwork of the three different regions: the innermost overdense region is given by a spatially closed FLRW universe, the intermediate underdense region, and the outermost flat FLRW universe as the background.
+We also note that the intermediate underdense region must be introduced to compensate for the innermost region’s mass excess and connect the spacetime to the background spacetime without a mass gap.
+Without the intermediate underdense region, this model cannot be an exact solution, even for the dust case.
+For the underdense region with a perfect fluid with $w=1/3$, obtaining any analytical and exact solutions of such dynamical inhomogeneous spacetimes with non-linear sound waves would not be tractable.
 Therefore, we stick to the three-zone model with dust.
 Furthermore, for simplicity, we focus on the model where the underdense region is described by the Schwarzschild solution, where the matching is possible without any singular hypersurface for the dust case.
-Despite this great simplification, the model provides a very powerful tool for capturing the overall dynamics of PBH and the structure of trapping horizons.
+Despite this significant simplification, the model provides a powerful tool for capturing the overall dynamics of PBH and the structure of trapping horizons.
 Note that one can also approach the same problem by using the Lemaître-Tolman-Bondi exact dust solution (c.f.
-[1]).
+[PhysRevD.83.124025]).
 The conformal diagrams for the three spacetimes of which the three-zone model consists are shown in Fig.
-15, where future $(\theta_{-};\mathcal{L}_{-}\theta_{+})|_{\theta_{+}=0}=(-;*)$, past (+;)(+;*)( + ; ∗ ), and bifurcating $(0;*)$ trapping horizons are denoted by red lines, blue lines, and the intersection point between them in the Schwarzschild region, respectively.
-For the dust case $w=0$, the BH apparent horizon and cosmological apparent horizon become future-outer (;)(-;-)( - ; - ) and past-inner (+;+)(+;+)( + ; + ), denoted by red -solid and blue-dashed lines, respectively.
+16, where future, past, and bifurcating trapping horizons are denoted by red lines, blue lines, and the intersection point between them in the Schwarzschild region, respectively.
+For the dust case $w=0$, the apparent horizon and cosmological apparent horizon become future-outer and past-inner, denoted by red-solid and blue-dashed lines, respectively.
+See Table 1 for the classification of trapping horiozons.
 First, let us construct the spacetime diagram of the case of the type I fluctuation.
-The extracted regions for the patchwork in the type I case are described by the yellow shaded regions in Fig.
-15, and the resultant spacetime diagram is shown in the left panel of Fig.
-17.
-These conformal diagrams define the right and left-directed future null vectors as $l_{\rm out}$ and $l_{\rm in}$, respectively.
-For the closed FLRW region, on the boundary of the extracted region, we can find $\theta_{\rm out}>0$ and $\theta_{\rm in}<0$ at the maximum expansion.
-Therefore, in the outside Schwarzschild region, we also require the same sign of $\theta_{\rm out}$ and $\theta_{\rm in}$.
+The extracted regions for the patchwork in the type I (type A) case are described by the yellow shaded regions in Fig.
+16, and the resultant spacetime diagram is shown in the left panel of Fig.
+18.
+For the closed FLRW region, on the boundary of the extracted region, we can find $\theta<0$ and $\theta^{\prime}>0$ at the maximum expansion.
+Therefore, in the outside Schwarzschild region, we also require the same sign of $\theta$ and $\theta^{\prime}$.
 Similarly, one can understand the matching of the Schwarzsdhild region and the background spatially flat FLRW spacetime.
 The extracted regions are yellow-shaded for the type II fluctuation in Fig.
-16.
+17.
 Since the type II fluctuation should have a maximal value of the areal radius, the boundary of the closed FLRW region must be outside the central point at which the areal radius takes the maximum value corresponding to the area of the big sphere at the maximum expansion.
-Then we find $\theta_{\rm out}<0$ and $\theta_{\rm in}>0$ on the boundary, which can be matched with the Schwarzschild region only in the left part of the Schwarzschild conformal diagram.
-Thus, the resultant conformal diagram for type II is given by the right panel of Fig.
-17.
+Then we find $\theta>0$ and $\theta^{\prime}<0$ on the boundary, which can be matched with the Schwarzschild region only in the left part of the Schwarzschild conformal diagram.
+Thus, the resultant conformal diagram for type II (type B) is given by the right panel of Fig.
+18.
+For reference, Fig. 19 gives the conformal diagrams of the closed and flat FLRW spacetimes with $w=1/3$ in the left and right panels, respectively.
 (a) closed FLRW for dust
@@
 (c) flat FLRW for dust
-Figure 15: Conformal diagrams to use for a patchwork for the spacetime of the type I PBH schematically in Fig.
+Figure 16: Conformal diagrams to use for a patchwork for the spacetime of the type I (type A) PBH schematically in Fig.
 18 for dust case $w=0$.
+The left, middle, and right panels show the conformal diagrams of the closed FLRW, the Schwarzschild, and the flat FLRW spacetimes, respectively.
 Dashed gray curves represent $R$-levels.
 Red (blue)-shaded regions represent future (past) trapped regions.
-Future-outer $(\theta_{-};\mathcal{L}_{-}\theta_{+})|_{\theta_{+}=0}=(-;-)$, past-outer (+;)(+;-)( + ; - ), future-inner (;+)(-;+)( - ; + ), and past-inner (+;+)(+;+)( + ; + ) trapping horizons are denoted by red-solid, blue -solid, red -dashed, and blue -dashed lines, respectively.
+Future-outer, past-outer, future-inner, and past-inner trapping horizons are denoted by red-solid, blue-solid, red-dashed, and blue-dashed lines, respectively.
+See Table 1 for the classification of trapping horizons.
 (a) closed FLRW for dust
@@
 (c) flat FLRW for dust
-Figure 16: Same as Fig. 15 but for the type II PBH. We use the yellow-shaded regions for a patchwork.
-(a) Type I PBH
-(b) Type II PBH
+Figure 17: Same as Fig. 16 but for the type II (type B) PBH. We use the yellow-shaded regions for a patchwork.
+(a) Type A PBH
+(b) Type B PBH
 Figure 18: Conformal diagrams of PBH for dust case $w=0$ as a patchwork of the three diagrams and trapping horizons.
-The left figure defines a type I (type A) PBH as a typical case PBH causal structure.
-The right one is constructed by changing the cutting diagram range and defining a type II (type B) PBH.
+The left panel defines a type I (type A) PBH as a typical case PBH causal structure.
+The right one is that for a type II (type B) PBH.
 (a) closed FLRW for w=1/3
 (b) flat FLRW for w=1/3
-Figure 18: Conformal diagrams of the closed and flat FLRW solutions for the radiation fluid case $w=1/3$.
+Figure 19: Conformal diagrams of the closed and flat FLRW solutions in the left and right panels, respectively, for the radiation fluid case $w=1/3$.
 In this case, all trapping horizons are null.
-Future-degenerate $(\theta_{-};\mathcal{L}_{-}\theta_{+})|_{\theta_{+}=0}=(-;0)$ and past-degenerate $(+;0)$ trapping horizons are denoted by red-dotted and blue-dotted lines, respectively.
+Future-degenerate and past-degenerate trapping horizons are denoted by red-dotted and blue-dotted lines, respectively.
+See Table 1 for the classification of trapping horizons.
 Appendix B Derivation of $\zeta_{\text{cFLRW}}$ in radial coordinate $r$
@@
 where $\chi$ means the comoving radial coordinate. By comparing the correspondence of the metric, we obtain
-$d\chi$ =±$=\pm dre^{\zeta},$ (B.3)
+$d\chi$ $=\pm dre^{\zeta},$ (B.3)
 $\frac{1}{k}\sin k\chi$ $=re^{\zeta},$ (B.4)
@@
 $\zeta=2\ln\frac{\cos k\chi_{a}/2}{\sqrt{\cos^{4}k\chi_{a}/2+\frac{1}{4}k^{2}r^{2}}}.$ (B.12)
-This functional form is shown as:
+This functional form is shown as in Fig. 20.
 Figure 20: Functional form of $\zeta_{\text{cFLRW}}$ for overdense region $\chi\leq\chi_{a}$.
-This profile is obtained by imposing a boundary condition that $\zeta=0$ at the edge of the overdense region $\chi=\chi_{a}$ [1].
+This profile is obtained by imposing a boundary condition that $\zeta=0$ at the edge of the overdense region $\chi=\chi_{a}$ [PhysRevD.83.124025].
 Appendix C Double-peaked compaction function and averaged density perturbation in the type II fluctuation
@@
 we find $r\zeta^{\prime}>-1$ for $0<r<r_{m1}$, $r_{m2}<r$, $r\zeta^{\prime}=-1$ for $r=r_{m1},r_{m2}$, and $r\zeta^{\prime}<-1$ for $r_{m1}<r<r_{m2}$.
-From the above assumption, we can deduce that $r\zeta^{\prime}$ must have at least one minimum below $-1$ with $(r\zeta^{\prime})^{\prime}=0$ between $r_{m1}$ and $r_{m2}$.
+From the above assumption, we can deduce that $r\zeta^{\prime}$ must have at least one minimum below 1-1 with $(r\zeta^{\prime})^{\prime}=0$ between $r_{m1}$ and $r_{m2}$.
 For simplicity, we additionally assume within this section that $r\zeta^{\prime}$ takes only one minimum at $r=r_{m3}$ between $r_{m1}$ and $r_{m2}$ and $r\zeta^{\prime}$ monotonically decreases for $0<r<r_{m3}$ and monotonically increases for $r_{m3}<r$.
-Since we have the compaction function 777 One can rewrite $C_{\rm SS}(r)$ as $C_{\rm SS}(r)=-r\zeta^{\prime}\quantity(1+\frac{1}{2}r\zeta^{\prime})=\frac{3}{4}\delta_{\rm l}\quantity(1-\frac{3}{8}\delta_{\rm l})=-\frac{9}{32}\left(\delta_{l}-\frac{4}{3}\right)^{2}+\frac{1}{2},$ where $\delta_{\rm l}\coloneqq-\frac{4}{3}r\zeta^{\prime}.$ is the ”linear-order density perturbation”.
+Since we have the compaction function  [^7] One can rewrite $C_{\rm SS}(r)$ as $C_{\rm SS}(r)=-r\zeta^{\prime}\quantity(1+\frac{1}{2}r\zeta^{\prime})=\frac{3}{4}\delta_{\rm l}\quantity(1-\frac{3}{8}\delta_{\rm l})=-\frac{9}{32}\left(\delta_{l}-\frac{4}{3}\right)^{2}+\frac{1}{2},$ where $\delta_{\rm l}\coloneqq-\frac{4}{3}r\zeta^{\prime}.$ is the ”linear-order density perturbation”.
 Therefore, $C_{\rm SS}$ as a function of $\delta_{\rm l}$ takes two zeroes at $\delta_{\rm l}=0$ and $\delta_{\rm l}=8/3$ and a maximum $1/2$ at $\delta_{\rm l}=4/3$.
-The type II fluctuation is characterized by the existence of the interval in $r$ with $\delta_{l}(r)>4/3$, where $C_{\rm SS}(r)$ as a function of $r$ takes two maxima $1/2$ at its boundaries with $\delta_{l}(r)=4/3$.
-$\mathcal{C}_{\rm SS}(r)=\frac{1}{2}\quantity(1-\quantity(1+r\zeta^{\prime})^{2})$ (C.3)
+The type II fluctuation is characterized by the existence of the interval in $r$ with $\delta_{l}(r)>4/3$, where $C_{\rm SS}(r)$ as a function of $r$ takes two maxima of $1/2$ at its boundaries with $\delta_{l}(r)=4/3$.
+$\mathcal{C}_{\rm SS}(r)=\frac{1}{2}\left[1-\quantity(1+r\zeta^{\prime})^{2}\right]$ (C.3)
 and its derivative
@@
 (2.5) as seen in Fig.
-20.
-Figure 20: The situation of type II fluctuation $r\zeta^{\prime}<-1$ from non-monotonicity of the areal radius $R$, $R^{\prime}<0$, and corresponding to the two peaks of the compaction function $\mathcal{C}_{\rm SS}$ (or averaged density perturbation $\bar{\delta}$ at the horizon entry).
-The density perturbation $\delta\coloneqq\rho/\rho_{\rm b}-1$ in a sub-leading order ($\order{\epsilon^{2}}$ ) of the long-wavelength solutions for uniform Hubble (constant mean curvature) slice is written as [21],
+21.
+Figure 21: The functions $r\zeta^{\prime}(r)$ and the compaction function ${\mathcal{C}}_{\rm SS}(r)$ are plotted in the top and bottom panels, respectively, for $\mu=1.0,1.2,1.4,1.6$ and $1.8$.
+We can see that for $\mu=1.4,1.6$ and $1.8$, $r\zeta^{\prime}(r)$ gets smaller than 1-1, corresponding to the two peaks of the value $1/2$ of the compaction function $\mathcal{C}_{\rm SS}$ (or averaged density perturbation $\bar{\delta}$ at the horizon entry), while this is not the case for $\mu=1.0$ and $1.2$.
+[…]
 $\delta\simeq-\frac{4}{3}e^{-5\zeta/2}\bar{\Delta}\quantity(e^{\zeta/2})\quantity(\frac{1}{aH_{\rm b}})^{2},$ (C.5)
-where $\bar{\Delta}$ is the Laplacian of the flat background geometry. The compaction function can be identified with the volume averaged density perturbation at the horizon entry $R=1/H_{\rm b}$,
+where $\bar{\Delta}$ is the Laplacian of the flat geometry. The compaction function can be identified with the volume averaged density perturbation at the horizon entry $R=1/H_{\rm b}$,
 $\left.\bar{\delta}\right|_{H_{\rm b}=1/R}\simeq 2\mathcal{C}_{\rm SS}.$ (C.6)
@@
 For the type II fluctuations ($\mu\gtrsim 1.4$), the value of $\bar{\delta}$ takes a minimal value at $r=\sqrt{2}/k\simeq 0.14$ with $k=10$, and the minimal value decreases with increasing $\mu$.
-It is worth noting that the separate universe (pinch-off) limit will not occur for this Gaussian-shaped profile of $\zeta$, while Kopp et al.
-considered the limit based on the other profile via the closed FLRW (or Three-zone) geometry of Eq.
-(4.3) (or Eq.(4.4)).
+It is worth noting that the separate universe (pinch-off) limit will not occur for this Gaussian-shaped profile of $\zeta$, while KHW considered the limit based on the three-zone model via the closed FLRW geometry of Eq.
+(4.3).